### v1

# Simulating Longitudinal Data from Marginal Structural Models

Abstract
Simulating longitudinal data from specified marginal structural models is a crucial but challenging task for evaluating causal inference methods and designing clinical trials. While data generation typically proceeds in a fully conditional manner using structural equations according to a temporal ordering, marginal structural models require capturing causal effects that are marginal over time-dependent confounders, making it difficult to align conditional distributions with target marginal quantities. To address this, we propose a flexible and efficient algorithm for simulating longitudinal data that adheres exactly to a specified marginal structural model. Recognizing the importance of time-to-event outcomes in clinical trials, we extend the method to accommodate survival models. Compared to existing approaches, our method offers several key advantages: it enables exact simulation from a known causal model rather than relying on approximations; avoids imposing restrictive assumptions on the data-generating process; and it is efficient as we need only evaluate analytic functions. This last benefit contrasts with methods that use computationally intensive techniques such as Monte Carlo approximations or numerical integration. Through simulation studies replicating realistic scenarios, we validate the method’s accuracy and utility. Our method will allow researchers to effectively simulate data with target causal structures for their specific scenarios.
1 Introduction
Longitudinal data are very common in studies analyzing the causal effect of a treatment (or exposure) on an outcome, especially in the context of chronic diseases such as diabetes or HIV. However, treatments are rarely fixed for the entire period of observation, and their status often varies over time. Even in clinical trials with strict adherence protocols, intercurrent events—such as discontinuation of the prescribed treatment and use of additional medication—are inevitable. For example, participants in the treatment group might discontinue treatment due to side effects, while those in the control group could begin treatment or alternative therapies if their health declines. In observational studies, treatment initiation or cessation can occur at any time, contingent on changing circumstances. The ICH E9111International Council on Harmonization of Technical Requirements for Registration of Pharmaceuticals for Human Use [ICH] Statistical Principles for Clinical Trials [E9] emphasizes that intercurrent events must be explicitly considered when defining the clinical question of interest (European Medicines Agency, 2020). While the intention-to-treat (ITT) principle simplifies this to some extent by reducing time-varying treatment to a point-treatment problem, it may not always yield the most meaningful estimand from a clinical or regulatory perspective. Longitudinal data, however, allow us to more accurately capture the treatment dynamics, enabling estimation of causal quantities like the “on-treatment” estimand, which represents the treatment effect assuming treatment is administered as intended.
A particular challenge in analyzing longitudinal data is time-varying confounding. This occurs when a time-dependent covariate predicts both the outcome and future treatment exposure, while past treatment influences the covariate. For example, in the case study by Robins et al. (2000) on the effect of zidovudine (AZT) treatment on HIV infection, the biomarker CD4 lymphocyte count is both affected by prior AZT treatment and is predictive of future therapy initiation and HIV progression. The presence of such time-varying confounding renders conventional adjustment methods, like regression and stratification, generally biased.
Marginal structural models, first introduced by Robins (1997), have emerged as a powerful tool for estimating the causal effect of time-dependent exposures in the presence of time-varying confounding. Model parameters can be consistently estimated using inverse-probability-of-treatment-weighted estimators. Among alternative methods, including structural nested models estimated through g-computation (Robins, 1989; Robins and Tsiatis, 1991), a systematic review of the literature by Clare et al. (2019) concluded that marginal structural models fitted with inverse probability of treatment weights are the most commonly used approach for longitudinal data with time-varying confounding due to their simplicity in both implementation and interpretation.
As novel causal inference methods for longitudinal data are continually proposed to address complex real-world scenarios, it is essential to evaluate their effectiveness for practical use. Simulation studies play a key role in this, by assessing important properties such as bias, confidence interval coverage, relative efficiency, and robustness to assumption violations. These evaluations guide statisticians in developing new methods and help analysts choose appropriate methods for specific research questions.
Sample size calculation also benefits from simulations. Prior to conducting a trial or an observational study, it is crucial to determine the number of participants, or number of observational units required to achieve adequate statistical power for the hypothesis test (for example, superiority, inferiority or equivalence). An improperly sized sample is clearly problematic: if too few units are recruited it can lead to undetected non-zero effects; if too many are recruited this may be a waste of resources and potentially unethical if it means that more individuals are given an inferior treatment. While standard formulas can be used for time-fixed treatments, more complex scenarios with longitudinal data and time-varying treatments make deriving analytical formulas impractical or impossible. In such cases, simulation is invaluable, allowing researchers to estimate sample sizes based on realistic data-generating processes and causal structures.
For both method evaluation and sample size calculation, simulating data that adheres to a known causal structure is crucial. However, simulating longitudinal data from a pre-specified marginal structural model is not a straightforward task. While generating data sequentially from structural equations—creating covariates, treatment status, and then outcome conditioned on past information—may seem intuitive, marginal structural models rely on marginal, rather than conditional models. This makes it difficult to sequentially simulate from conditional models in a way that aligns with a pre-specified marginal model; in particular, it is very difficult to sample from a null hypothesis.
The remainder of this paper is structured as follows: Section 2 reviews existing methods for simulating longitudinal data from causal models. Section 3 covers the preliminaries, including the frugal parameterization (Evans and Didelez, 2024) and copulas. In Section 4, we introduce our proposed methodology and illustrate it in Section 5 through a simulation study replicating a real-world study in epidemiology (Bodnar et al., 2004). Section 6 extends the simulation approach to time-to-event data, enabling simulations from survival models such as Cox marginal structural models, including an illustrative example. The paper concludes with a discussion in Section 7.
2 Related Literature
As noted in the previous section, the challenge lies in ensuring that simulated data adhere to a specified marginal model, rather than simply following sequential conditional distributions in a structural equation model. There are two possible strategies to address this issue:
  1. (i)
    Simulate data from the conditionals and then derive the corresponding marginal models. The g-formula, simplifying assumptions, and Monte Carlo approximations, can help bridge the gap between the simulated conditionals and the target marginal model.
  2. (ii)
    Alternatively, bypass conditional simulations and develop methods to simulate directly from the target marginal model.
Both strategies are explored in existing approaches.
2.1 Simulate from the conditionals and derive the equivalent marginal models
Xiao et al. (2010) conducted a simulation study to evaluate the accuracy of two methods of fitting a Cox marginal structural model. They generated their dataset through the natural approach—sequentially through conditional models—and argued that under the assumption of a rare disease, i.e. a very low failure rate, the impact of non-collapsibility is negligible. In this scenario, the parameters of the underlying Cox marginal structural model may be approximated from the conditional model parameters. Notably, for this approximation to hold, there must be no confounders for the relationship between the time-varying confounder and the outcome, whether observed or unobserved. Essentially, their study did not directly simulate data from an marginal structural model but demonstrated a case where conditional simulation can serve as a reasonable approximation of such a model.
Young and Tchetgen Tchetgen (2014) explored bridging from conditionals to marginals using the g-formula in simulating from a discrete time Cox marginal structural model. In addition to standard identifying assumptions—consistency, positivity, and sequential exchangeability—they further assume that the g-formula uses standard parametric models, and hence derived the functional form of the marginal model. If the time-varying covariates consist only of discrete variables, then deriving the equivalent marginal relative hazard is theoretically possible, but it can be cumbersome. However, the authors noted that if the time-varying covariate includes continuous components, the summation in the above equation is replaced by integrals which may be challenging to evaluate due to the lack of closed-form solutions. The authors suggest potential simplifications such as the Markov property or approximation based on the rare disease assumption, as used by Xiao et al. (2010). A significant limitation arises when attempting to simulate data from a specific marginal structural model. Other than in very specific cases, such as for discrete or linear models, it is essentially impossible to reverse-engineer the series of equivalent conditional models. For instance, if we want to simulate from $h(t,\bar{a}_{t})=0.1+0.5a_{t}+0.3a_{t-1}-a_{t}a_{t-1}$, the corresponding series of conditionals is difficult to derive. This poses a drawback in real-world applications, especially for sample size determination, where experimenting with different effect sizes is necessary to generate power curves.
Sofrygin et al. (2017) developed the simcausal R package as a flexible, practical tool for simulating complex longitudinal data structures from non-parametric structural equation models. The package supports simulations that replicate realistic data scenarios common in causal inference, such as time-varying confounding and selection bias. Users can specify and simulate counterfactual data under various intervention regimes. A typical simcausal workflow involves parameterizing structural equations based on a causal directed acyclic graph and simulating data sequentially from the specified conditional distributions. Since deriving analytic expressions for marginal causal quantities, such as the average treatment effect, is often intractable with this type of conditional parameterization, users instead estimate these parameters through Monte Carlo approximations by simulating counterfactual data. While this approach is suitable for marginal causal quantities such as the average treatment effect, it may lack consistency for causal structures such as marginal structural models or conditional average treatment effect functions due to potential model misspecification. The authors acknowledged this limitation, as they allow users to specify a “working” model and cautioned that this approximation serves as the best parametric fit within the model class, without guaranteeing consistency.
Keogh et al. (2021) described how to simulate from an additive-hazard marginal structural model (Aalen, 1989). They utilized the result that if the conditional hazard model follows an additive form, the equivalent marginal hazard model will also be additive. Unlike Cox hazard ratios, the parameters of the additive hazard model are collapsible. Their strategy involved simulating data conditionally and then deriving the corresponding marginal hazard. However, while this approach ensures the correct form of the marginal model, the true values of the coefficients remain generally intractable. To address this, as with the workflow outlined in Sofrygin et al. (2017), the authors find a Monte Carlo approximation of the true parameters through simulating counterfactual data. A notable limitation of the additive hazard model is that it does not constrain the hazard to be non-negative. This becomes problematic when the hazard depends on continuous covariates, as ensuring positivity may be difficult in simulated data. Additionally, as with the approach by Young and Tchetgen Tchetgen (2014), this method does not allow direct specification of the model parameters, limiting its applicability in real-world scenarios.
2.2 Simulate directly from the target marginal model
Havercroft and Didelez (2012) approached this problem from a novel angle by simulating directly from the causal margin (in their case, a discrete Cox marginal structural model) and then introducing time-varying confounding to establish its relationships with the outcome and the treatment. Specifically, they used an inverse cumulative distribution function to simulate from the causal margin, with time-varying confounding introduced indirectly via a shared latent variable. This approach omits any direct edge from the time-varying confounder directly to the outcome. The primary limitation of this method is that time-varying confounding is not explicitly modeled but introduced via an artificial latent variable. As a result, the assumption that time-varying covariates have no direct effect on outcomes may not always hold in clinical contexts. For instance, in the HIV study example, this would assume that CD4 cell count has no direct effect on survival, which is unlikely. Nonetheless, their approach presents a valuable contribution.
Evans and Didelez (2024) recently proposed a frugal parameterization that prioritizes the causal margin and parameterizes the surrounding variables accordingly. This innovation enables direct simulation from the marginal structural model, addressing limitations identified in previous work. Specifically, it allows for explicit simulation of time-varying covariates and an exact marginal model. More details on the frugal parameterization are given in Section 3.1.
Building on this foundation, Seaman and Keogh (2024) applied the frugal parameterization to simulate longitudinal data with survival outcomes. Similar to the approach of Havercroft and Didelez (2012), they simulated the outcome using the inverse cumulative distribution function of the target causal margin. To handle high-dimensional time-varying covariates, they summarized their effects into a one-dimensional “risk score”, simplifying the dependence between covariates and the outcome to a bivariate copula. However, this approach has notable limitations. Since the cumulative distribution functions of the risk score are generally unknown, they relied on Monte Carlo approximations, which involve sampling a large number of individuals with the same treatment history and baseline variables. It is a computationally expensive task and must be performed at each time step. For example, with binary treatment, there could be up to $10242^{10}=$ 1024 combinations in terms of treatment history distribution functions to approximate at step 10 – potentially even more if baseline variables are included. The method becomes infeasible when treatment or baseline variables are continuous. The method also becomes impractical when treatment history or baseline variables have many possible combinations or are continuous. Moreover, the use of a risk score limits flexibility, as it prevents explicit modeling of specific covariate-outcome dependencies, such as the effect of age on survival. A more flexible approach would directly model these associations, making fuller use of the frugal parameterization.
The literature review highlights gaps in current methodologies and identifies key features desired in an improved simulation approach. These include the ability to bypass the conditional distribution of the outcome and simulate directly from the marginal structural model, robustness that avoids assumptions about the data-generating process, flexibility to accommodate diverse scenarios, and practicality for real-world applications.
3 Preliminaries
3.1 The frugal parameterization
As discussed in the previous section, the fundamental challenge in simulating data from longitudinal marginal structural models arises from the fact that, while the joint data distribution can be easily factorized into conditionals for simulation, the causal quantity of interest is a marginal distribution. Earlier methods, as detailed in Section 2, tackle this by deriving the marginal distribution implied by the conditional factorization. However, Evans and Didelez (2024) significantly advanced this area with the introduction of the frugal parameterization, which explicitly centers the causal marginal and constructs the rest of the joint distribution around it. This breakthrough has made the simulation method presented in this paper possible.
The frugal parameterization decomposes the joint distribution $p(\bm{l},a,y)$ into three separate pieces:
  1. (i)
    the past: $p_{\bm{L}A}(\bm{l},a)$—the joint distribution of covariates and treatment prior to observing outcomes;
  2. (ii)
    the causal margin: $p^{*}_{Y\mid A}(y\,|\,do(a))$—the causal margin of interest, set according to the marginal structural models relevant to the causal question;
  3. (iii)
    a dependency measure: $\phi^{*}_{Y\bm{L}\mid A}$—captures the dependency between the outcome $Y$ and the covariates $\bm{L}$ under intervention.
The first component, termed “the past”, refers to the distribution of covariates and treatment, which precede the outcome. The second component, the causal margin, aligns with the marginal structural model we aim to simulate, directly addressing the causal question at hand. The third component captures the dependency between the outcome $Y$ and the set of covariates $\bm{L}$ jointly. A common choice of this dependency, $\phi$, is a copula, which we will elaborate on in the following subsection.
3.2 Copulas
The term copula is derived from the Latin verb copulare, meaning “to join together.” We can think of continuous multivariate distributions as the joining together of the univariate marginals (or margins) and the dependence structure. This structure can be represented by a copula, which is a multivariate distribution function $C:[0,1]^{d}\rightarrow[0,1]$ with margins uniformly distributed on $(0,1)$. According to Sklar’s theorem (Sklar, 1959, 1973), any continuous multivariate distribution $F$ with marginals $F_{1},\dots F_{d}$ can be expressed as
$F(x_{1},\dots,x_{d})=C\left\{F_{1}(x_{1}),\dots,F_{d}(x_{d})\right\}=C(u_{1},\dots,u_{d}),$
where $\bm{X}=(X_{1},\dots,X_{d})$ is a vector of random variables, and $u_{i}=F_{i}(x_{i})$, $i=1,\dots,d$. The copula then takes the form
$C(u_{1},\dots,u_{n})=F\{F_{1}^{-1}(u_{1}),\dots,F_{d}^{-1}(u_{d})\},\qquad(u_{1},\dots,u_{d})\in[0,1]^{d},$
where $x_{i}=F_{i}^{-1}(u_{i})$. The mathematical relationship between copula density function $c(u_{1},\dots,u_{d})$ and joint density function $f(x_{1},\dots,x_{d})$, applying the chain rule, is given by
$f(x_{1},\dots,x_{d})=c(u_{1},\dots u_{d})\prod_{i=1}^{d}f_{i}(x_{i}).$ (1)
The choice of copula can be considered independently of the univariate margins.
3.2.1 Pair-copula construction
Building higher-dimensional copulas is generally recognized as a difficult problem. While there exist a wider variety of parametric bivariate copulas, the options for higher-dimensional copulas are limited, with Gaussian and Student t-copulas being common choices. Extending symmetric bivariate Archimedean copulas to multivariate settings is also possible but imposes strong restrictions, such as requiring the same level of dependence among all possible pairs. First introduced by Joe (1996), pair-copula constructions provide a flexible and intuitive approach to building multivariate copulas by modeling multivariate data as a series of bivariate copulas. This hierarchical structure incorporates additional variables at each level, using pair-copulas as the fundamental building blocks.
Following the general formula given in Joe (1997), the conditional distribution of one variable $y$ given $d$ other variables can be decomposed to $d$ bivariate copula density functions, each of which can be specified flexibly:
$f(y\,|\,\bm{x})=c_{YX_{j}|\bm{X}_{-j}}\left\{F(y\,|\,\bm{x}_{-j}),F(x_{j}\,|\,\bm{x}_{-j})\right\}\cdot f(y\,|\,\bm{x}_{-j}),$ (2)
where $\bm{x}$ is a $d$-dimensional vector; $x_{j}$ is an arbitrarily chosen component of $\bm{x}$, and $x_{-j}:=(x_{1},\ldots,x_{j-1},x_{j+1},\ldots,x_{d})$ denotes $\bm{x}$ with its $j$th entry removed. Applying (2) recursively to the conditional densities and, for convenience, choosing an order of $1,2,\dots d$ we have:
$f(y\,|\,\bm{x})$ $=c_{YX_{d}|\bm{X}_{1:{d-1}}}\{F(y\,|\,\bm{x}_{1:{d-1}}),F(x_{d}\,|\,\bm{x}_{1:{d-1}})\}$
$\quad\cdot c_{YX_{d-1}|\bm{X}_{1:{d-2}}}\{F(y\,|\,\bm{x}_{1:{d-2}}),F(x_{d-1}\,|\,\bm{x}_{1:{d-2}})\}$
[…]
$\quad\cdot c_{YX_{2}|X_{1}}\{F(y\,|\,x_{1}),F(x_{2}\,|\,x_{1})\}$
$\quad\cdot c_{YX_{1}}\{F(y),F(x_{1})\}\cdot f(y)$ (3)
$=f(y)\cdot\prod_{i=1}^{d}c_{YX_{d+1-i}|\bm{X}_{1:{d-i}}}\{F(y\,|\,\bm{x}_{1:{d-i}}),F(x_{d+1-i}\,|\,\bm{x}_{1:{d-i}})\},$ (4)
where each copula in the product in (4) is bivariate, i.e. a pair-copula. In principle, any ordering of the conditioned variables $x_{1},x_{2},\dots,x_{d}$ can be selected, although some orderings may be easier to use than others. In Section 4, we choose a particular ordering that is convenient and efficient for simulation.
In conclusion, this shows that under appropriate regularity conditions, a multivariate density can be expressed as a product of pair-copulas, acting on different conditional distributions.
3.2.2 Simulating from pair-copula-constructions
The pair-copula construction includes conditional distributions of the form $F(y\,|\,\bm{x})$. Joe (1996) demonstrated the following result:
$F(y\,|\,\bm{x})=\frac{\partial C_{YX_{j}\mid\bm{X}_{-j}}\left\{F(y\,|\,\bm{x}_{-j}),F(x_{j}\,|\,\bm{x}_{-j})\right\}}{\partial F(x_{j}\,|\,\bm{x}_{-j})}.$ (5)
Proof.
[…]
$=\frac{\partial}{\partial F(x_{j}\,|\,\bm{x}_{-j})}C_{YX_{j}\mid\bm{x}_{-j}}\{F(y\,|\,\bm{x}_{-j}),F(x_{j}\,|\,\bm{x}_{-j})\}.$
For convenience of notation, we introduce the h-function $h(x_{1},x_{2};\Theta)$ from Aas et al. (2009). Let $x_{1}$ and $x_{2}$ be copula random variables, i.e. with margins following the uniform distribution $x_{i}\sim\operatorname{U}(0,1)$ and $F_{i}(x_{i})=x_{i}$, for $i=1,2$. The $h$-function is defined as:
$h(x_{1},x_{2};\Theta)$ $:=F(x_{1}\,|\,x_{2})=\frac{\partial F_{12}(x_{1},x_{2})}{\partial x_{2}f_{2}(x_{2})}$
$=\frac{\partial}{\partial F_{2}(x_{2})}C_{12}(F_{1}(x_{1}),F_{2}(x_{2});\Theta)$ (6)
$=\frac{\partial}{\partial x_{2}}C_{12}(x_{1},x_{2};\Theta),$ (7)
where the second argument in $h(\cdot;\Theta)$ is the conditional variable, and $\Theta$ represents the set of copula parameters. From (7) we can see that the $h$-function is essentially the conditional copula function, $C(u_{1}\,|\,u_{2};\Theta)$.
Example 3.1.
The derivation of the analytical form of $h(\cdot)$ for common bivariate copulas is straightforward following (7) for parametric copula families. Some examples include:
  • Gaussian copula: $h(x_{1},x_{2};\rho)=\Phi\left(\frac{\displaystyle\Phi^{-1}(x_{1})-\rho\Phi^{-1}(x_{2})}{\displaystyle\sqrt{1-\rho^{2}}}\right)$;
  • Student’s $t$ copula: $h(x_{1},x_{2};\rho,\nu)=t_{\nu+1}\left(\frac{\displaystyle\sqrt{\nu+1}(t_{\nu}^{-1}(x_{1})-\rho t_{\nu}^{-1}(x_{2}))}{\displaystyle\sqrt{\left(\nu+\left(t_{\nu}^{-1}(x_{2})\right)^{2}\right)\left(1-\rho^{2}\right)}}\right)$;
  • Clayton copula: $h(x_{1},x_{2};\theta)=x_{2}^{-\theta-1}\left(x_{1}^{-\theta}+x_{2}^{-\theta}-1\right)^{-1-1/\theta}$;
  • Gumbel copula: $h(x_{1},x_{2};\theta)=C(x_{1},x_{2};\theta)\frac{1}{x_{2}}(-\log x_{2})^{\theta-1}\left((-\log x_{1})^{\theta}+(-\log x_{2})^{\theta}\right)^{1/\theta-1}$.
The $h$-functions are foundational for simulating from $F(y\,|\,\bm{x})$. Similar to (4), by imposing an order $1,\dots,d$ on the components of $\bm{x}$, we can recursively apply $h$-functions to obtain:
$F(y\,|\,\bm{x})$ $=h\{F(y\,|\,\bm{x}_{1:{d-1}}),F(x_{d}\,|\,\bm{x}_{1:{d-1}}),\Theta_{1}\}$
$=h[h\{F(y\,|\,\bm{x}_{1:{d-2}}),F(x_{d-1}\,|\,\bm{x}_{1:{d-2}};\Theta_{2}\},F(x_{d}\,|\,\bm{x}_{1:{d-1}});\Theta_{1}]$
$=h\left[h\{h\left(F(y\,|\,\bm{x}_{1:{d-3}}),F(x_{d-2}\,|\,\bm{x}_{1:{d-3}};\Theta_{3})\right),F(x_{d-1}\,|\,\bm{x}_{1:{d-2})};\Theta_{2}\},F(x_{d}\,|\,\bm{x}_{1:{d-1}});\Theta_{1}\right]$
$\dots$
$=h[\dots h\{h(F(y),F(x_{1});\Theta_{d}),F(x_{2}\,|\,x_{1});\Theta_{d-1}\}\dots;\Theta_{1}].$ (8)
Thus, $Y$ given $\bm{X}$ can be simulated by first drawing a uniform $U\sim{\rm U}(0,1)$, and then evaluating:
$Y=h^{-1}[\dots h^{-1}\{h^{-1}(U,\,F(X_{1});\Theta_{d}),F(X_{2}\,|\,X_{1});\Theta_{d}\}\dots;\Theta_{1}].$ (9)
This means that to simulate $Y$ given a $d$-dimensional vector $\bm{X}$, we recursively calculate the $d$ inverse $h$-functions, forming a PCC. This forms the basis of our proposed simulation method.
4 Simulation Algorithm
[…]
Figure 1: An example directed acyclic graph showing temporal relationships in a longitudinal discrete setting, with baseline covariates $Z$; treatments $A_{0},A_{1},\ldots,A_{K}$; time-varying covariates $L_{1},\ldots,L_{K}$ and outcomes $Y$ observed at the end of the K follow up intervals. Given the topological ordering $Z,L_{0},A_{0},L_{1},A_{1}\ldots,L_{K},A_{K},Y$, any variable may depend upon any subset of the previous variables.
We consider a longitudinal study with $n$ individuals observed over discrete time points, $\mathcal{K}=\{0,1,\dots,K\}$. At each follow-up $k$, treatment status $A_{k}$, and a set of time-varying covariates $L_{k}$ are recorded. Additionally, $Z$ denotes a set of static covariates measured at baseline. There may also be latent variables, represented by $\{U_{k}\}_{k\in\mathcal{K}}$. To indicate the history of a variable up to time $k$, we use a bar notation. For instance, $\bar{A}_{k}=\{A_{0},A_{1}\dots,A_{k}\}$ and $\bar{L}_{k}=\{L_{0},L_{1},\dots,L_{k}\}$. For convenience, we let $\bar{A}_{-1}=\emptyset$ and $\bar{L}_{-1}=\emptyset$. The outcome $Y$ is observed at the final time point $K$. Figure 1 shows a DAG representing this setup, which is common in medical and epidemiological research, such as examining the effects of iron supplement use during pregnancy on the odds of anemia at delivery (Bodnar et al., 2004), anti-diabetic drugs on weight loss over a specified period (e.g. Jastreboff et al., 2022; Ryan et al., 2024) and exposure to community violence on youth’s mental health (Kennedy et al., 2023). In Section 5, we will illustrate our method with a simulation study that replicates such setups.
Following the notation of Evans and Didelez (2024), we use an asterisk to represent causal or interventional distributions where treatment is set to a specific value $A_{k}=a$ via intervention. In graphical terms, this corresponds to removing all arrows directed into $A_{k}$.
4.1 Fully conditional approach
For simplicity, we initially assume the time-varying covariate $L_{k}$ is one-dimensional and exclude the latent process $\{U_{k}\}_{k\in\mathcal{K}}$; the extension to multi-dimensional time-varying covariates and latent variables is discussed later in this section.
The usual strategy for simulation of longitudinal samples $(Z,\bar{L}_{K},\bar{A}_{K},Y)$ is by time step in blocks using the Rosenblatt transformation (Rosenblatt, 1952). Specifically:
  1. 1.
    Baseline simulation: Simulate static baseline variables $z=F_{Z}^{-1}(w_{z})$, where $w_{z}$ is a uniform random variable.
  2. 2.
    Time-step simulation: For each time point $k=0,1,\dots,K$, simulate:
    $l_{k}$ $=F_{L_{k}}^{-1}(w_{k,1},\,|\,z,\bar{l}_{k-1},\bar{a}_{k-1})$ (10)
    $a_{k}$ $=F_{A_{k}}^{-1}(w_{k,2},\,|\,z,\bar{l}_{k},\bar{a}_{k-1}),$
    where $w_{k,1},w_{k,2}$ are random uniform variables.
  3. 3.
    Outcome simulation: After simulating all time-varying covariates and treatment, simulate the outcome $Y$ as
    $y$ $=F_{Y}^{-1}(w_{y}\,|\,z,\bar{l}_{K},\bar{a}_{K}),$ (11)
    with $w_{y}$ uniformly distributed.
4.2 Frugal Parameterization for Outcome Simulation
Our method diverges from traditional conditional approaches by using the frugal parameterization for $Y$ to avoid intractable analytical calculations for the marginal structural model. Instead of directly assuming a conditional distribution $F_{Y}^{-1}(w_{y}\,|\,z,\bar{l}_{K},\bar{a}_{K})$, we apply interventional distributions to decompose this into interventional margins and dependency structures.
Evans and Didelez (2024) showed that the conditional distribution $F_{Y}^{-1}(w_{y}\,|\,z,\bar{l}_{k},\bar{a}_{K})$ can be frugally parameterized as:
$p(y\,|\,z,\bar{l}_{k},\bar{a}_{K})=p^{*}_{Y\mid Z\bar{A}_{K}}(y\,|\,z,do(\bar{a}_{K}))\cdot c^{*}_{Y\bar{L}_{K}\mid Z\bar{A}_{K}}(u^{*}_{y},u^{*}_{l_{0}},\dots,u^{*}_{l_{K}}\,|\,z,do(\bar{a}_{K})),$ (12)
where $u^{*}_{y}=F(y\,|\,z,do(\bar{a}_{K}))$ and $u^{*}_{l_{k}}=F(l_{k}\,|\,z,do(\bar{a}_{K}))$ for $k=0,1,\ldots,K$ are the marginal quantiles under interventions that set $\bar{A}_{K}=\bar{a}_{K}$. We mark these quantiles with an asterisk to emphasize their interventional nature. This decomposition implies that we can simulate $Y$ through the interventional copula and the interventional margin. Applying the pair-copula-construction from (4) to the expression in (12), we obtain:
$p(y\,|\,z,\bar{l}_{K},\bar{a}_{K})$ $=c^{*}_{YL_{K}\mid\bar{L}_{K-1},Z,\bar{A}_{K}}\{F(y\,|\,\bar{l}_{K-1},z,do(\bar{a}_{K})),F(l_{K}\,|\,\bar{l}_{K-1},z,do(\bar{a}_{K}));\Theta_{K}\}$
$\quad\cdot c^{*}_{YL_{K-1}\mid\bar{L}_{K-2},Z,\bar{A}_{K}}\{F(y\,|\,\bar{l}_{K-2},z,do(\bar{a}_{K})),F(l_{K-1}\,|\,\bar{l}_{K-2},z,do(\bar{a}_{K}));\Theta_{K-1}\}$
[…]
$\quad\cdot c^{*}_{YL_{0}\mid Z,\bar{A}_{K}}\{F(y\,|\,z,do(\bar{a}_{K})),F(l_{0}\,|\,z,do(\bar{a}_{K}));\Theta_{0}\}\cdot p^{*}_{Y\mid Z\bar{A}_{K}}(y\,|\,z,do(\bar{a}_{K}))$ (13)
$=\prod_{k=0}^{K}c^{*}_{Y,L_{k}\mid\bar{L}_{k-1},Z,\bar{A}_{K}}\{F(y\,|\,\bar{l}_{k-1},z,do(\bar{a}_{K})),F(l_{k}\,|\,\bar{l}_{k-1},z,do(\bar{a}_{K}));\Theta_{k}\}.$ (14)
This decomposes the conditional density into the product of $K+1$ pair-copulas and the target marginal causal density. Under intervention, graphically speaking, all arrows into nodes $\bar{A}_{k}$ are removed in the directed acyclic graph, rendering $L_{k}$ independent of subsequent interventions. Consequently, $F(l_{k}\,|\,\bar{l}_{k-1},z,do(\bar{a}_{K}))=F(l_{k}\,|\,\bar{l}_{k-1},z,\bar{a}_{k-1})$ for $k=1,2,\ldots,K$ and $F(l_{0}\,|\,z,do(\bar{a}_{K}))=F(l_{0}\,|\,z)$. Given this, following the simulation strategy from pair-copula constructions as described in Section 3.2.2, the conditional cumulative distribution function in (11) becomes
$F(y\,|\,z,\bar{l}_{K},\bar{a}_{K})$ $=h[\dots h\{h(F^{*}(y\,|\,z,\bar{a}_{K}),F(l_{0}\,|\,z);\Theta_{0}),F(l_{1}\,|\,l_{0},z,a_{0});\Theta_{1}\}\dots;\Theta_{K}].$
Through inversion the outcome can be simulated in two steps:
  1. 1.
    Calculate $u^{*}_{y}$ as
    $u^{*}_{y}$ $=h^{-1}\big{[}\dots h^{-1}\big{\{}h^{-1}(w_{y},F(l_{K}\,|\,\bar{l}_{K-1},z,\bar{a}_{K});\Theta_{K}),$
    $\qquad\qquad\qquad F(l_{K-1}\,|\,\bar{l}_{K-2},z,\bar{a}_{K-1});\Theta_{K-1}\big{\}}\dots,F(l_{0}\,|\,z);\Theta_{0}\big{]}$ (15)
    $=h^{-1}\big{[}\dots h^{-1}\big{\{}h^{-1}(w_{y},w_{K,1};\Theta_{K}),w_{K-1,1};\Theta_{K-1}\big{\}},\dots,w_{0,1};\Theta_{0}\big{]}.$ (16)
  2. 2.
    Obtain $y$ as:
    $y=F^{*-1}(u^{*}_{y}\,|\,z,do(\bar{a}_{K})).$
The quantiles in the inverse $h$ functions, such as $F(l_{k}\,|\,\bar{l}_{k-1},\bar{a}_{k},z)$, are precisely the random variables $w_{k,1}$ used in simulate them in (2). This means that we can directly substitute the $l_{k}$ quantiles into (15) without the need for computation.
4.3 Simulation Algorithm
To simulate data $(Z,\bar{L}_{K},\bar{A}_{K},Y)$ in this example, the user specifies:
  • the population distribution of the static baseline covariate $Z$: $F_{Z}$;
  • the development of the time-varying confounder $L_{k}$ as influenced by $Z$ and treatment history $\bar{A}_{k-1}$: $F_{L_{k}\mid Z,\bar{L}_{k-1},\bar{A}_{k-1}}$;
  • treatment assignment based on the time-varying confounder $L_{k}$, past treatment and baseline covariate $Z$: $F_{A_{k}\mid Z,\bar{L}_{k},\bar{A}_{k-1}}$;
  • the target causal marginal structural model: $F_{Y_{k+1}\mid\bar{A}_{k},Z}^{*}(\cdot\,|\,do(\bar{a}_{K}),z)$;
  • $K+1$ pair-copulas describing dependencies between outcome $Y$ and $L_{0},L_{1},\ldots,L_{K}$:
    $C_{YL_{k}\mid Z,do(\bar{A}_{k}),\bar{L}_{k-1}}\quad\text{for}\quad k=0,1,\ldots,K.$
Section 5 provides examples of specifying and interpreting these inputs through a simulation study that capture real-world complexities.
Algorithm 1 below details the simulation procedure:
Algorithm 1 Simulation algorithm for an individual $i$ over $K$ follow-ups
Sample $w_{z}\sim U(0,1)$
$z\leftarrow F_{Z}^{-1}(w_{z})$
Let $k=0$
while $k\leq K$ do
     Sample $w_{k,1},w_{k,2}$ independent uniform on $[0,1]$.
     $l_{k}\leftarrow F^{-1}_{L_{k}\mid Z,\bar{L}_{k-1},\bar{A}_{k-1}}(w_{k,1}\,|\,z,\bar{l}_{k-1},\bar{a}_{k-1})$
     $a_{k}\leftarrow F^{-1}_{A_{k}\mid Z,\bar{L}_{k},\bar{A}_{k-1}}(w_{k,2}\,|\,z,\bar{l}_{k},\bar{a}_{k-1})$
     $k\leftarrow k+1$
end while
Sample $w_{y}\sim U(0,1)$
Let $\nu_{0}=w_{y}$
for $j=0$ to $K$ do
     $\nu_{j+1}\leftarrow h^{-1}(\nu_{j},w_{(K-j),1};\Theta_{K-j})$
end for
$u^{*}_{y}\leftarrow\nu_{K+1}$
$y\leftarrow F_{Y_{k+1}\mid\bar{A}_{k},Z}^{*-1}(u^{*}_{y}\,|\,do(\bar{a}_{K}),z)$
return $(z,\bar{l}_{K},\bar{a}_{K},y)$
4.4 Extension to Include Time-Varying Latent Variables
We now extend the simulation algorithm discussed above to allow for a latent process that confounds the covariates and the final outcome $Y$. As an example, the steps to simulate longitudinal data from the model depicted in Figure 2 are:
[…]
Figure 2: Extension to the model in Figure 1 in which we also allow an unobserved process $U_{t}$ affecting only $Z$, $L_{k}$ and $Y$.
  1. 1.
    Baseline simulation: The DAG in Figure 1 imposes that the latent variable $U_{0}$ precedes $Z$ in topological order, which may indicate an initial selection bias. Thus, we first simulate the baseline value of $U$ as $u_{0}=F_{U_{0}}^{-1}(w_{0,0})$, followed by the static baseline covariate $Z$ given $U_{0}$: $z=F_{Z\mid U_{0}}^{-1}(w_{z}\,|\,u_{0})$, where $w_{0,0}$ and $w_{z}$ are uniform random numbers.
    $u_{0}$ $=F_{U_{0}}^{-1}(w_{0,0})$
    $z$ $=F_{Z\mid U_{0}}^{-1}(w_{z}\,|\,u_{0})$
    $l_{0}$ $=F_{L_{0}\mid ZU_{0}}^{-1}(w_{0,1}\,|\,z,u_{0})$
    $a_{0}$ $=F_{A_{0}\mid ZU_{0}L_{0}}^{-1}(w_{0,2}\,|\,z,u_{0},l_{0})$
  2. 2.
    Time-step simulation: For each time point $k=1,\dots,K$, simulate:
    $u_{k}$ $=F_{U_{k}\mid\bar{U}_{k-1}}^{-1}(w_{k,1}\,|\,\bar{u}_{k-1})$
    $l_{k}$ $=F_{L_{k}\mid Z\bar{U}_{k}\bar{L}_{k-1}\bar{A}_{k-1}}^{-1}(w_{k,2},\,|\,z,\bar{u}_{k},\bar{l}_{k-1},\bar{a}_{k-1})$
    $a_{k}$ $=F_{A_{k}\mid Z\bar{U}_{k}\bar{L}_{k}\bar{A}_{k-1}}^{-1}(w_{k,3},\,|\,z,\bar{u}_{k},\bar{l}_{k},\bar{a}_{k-1})$
    where $w_{k,1},w_{k,2},w_{k,3}$ are random uniform variables.
  3. 3.
    Outcome simulation: After simulating all time-varying covariates and treatment, and applying the same derivation following the frugal parameterization as in Section 4.2, simulate the outcome $Y$ in two steps:
    1. (a)
      Calculate
      $u^{*}_{y}$ $=h^{-1}\big{[}\dots h^{-1}[h^{-1}\left\{h^{-1}\left(w_{y},F(l_{K}\,|\,\bar{l}_{K-1},\bar{u}_{K},z,\bar{a}_{K});\Theta_{K,1}\right),F(u_{K}\,|\,\bar{l}_{K-1},\bar{u}_{K-1},z,\bar{a}_{K});\Theta_{K,2}\right\}$
      $F(l_{K-1}\,|\,\bar{l}_{K-2},\bar{u}_{K-1},z,\bar{a}_{K-1});\Theta_{K-1,1}]\dots;\Theta_{0,2}\big{]}.$
    2. (b)
      Obtain: $y=F^{*^{-1}}(u^{*}_{y}\,|\,z,do(\bar{a}_{K}))$.
    Here the $w_{y}$ are independent and uniformly distributed on $(0,1)$.
In addition to the information required in the case without latent confounders, the user also specifies:
  • the development of the latent confounder $U_{k}$ as influenced by $Z$ and treatment history $\bar{A}_{k}$;
  • $K+1$ pair-copula densities describing dependencies between outcome $Y$ and each of $U_{0},U_{1},\ldots,U_{K}$:
    $c_{YU_{k}\mid Z,\bar{U}_{k-1},\bar{L}_{k-1},do(\bar{A}_{k-1})},$
    for $k=0,1,\ldots,K$.
Extending to include additional time-varying covariates follows a similar approach. We specify the evolution of each time-varying covariate and use pair copulas to parameterize their association with the outcome, conditioned on all preceding variables. With common choices of parametric bivariate copulas, evaluating the $h$-functions is fast, making the simulation efficient. Increasing the dimensions of time-varying confounders, and the number of time steps does not impose significant computational challenges.
5 Demonstration
We first demonstrate our simulation method with an example replicating the setting in Bodnar et al. (2004), which examines the causal effect of iron supplement use during pregnancy on the odds of anemia at delivery.
Participants are initially randomized into treatment and control groups and are followed up in four visits leading up to delivery. Like the Iron Supplementation Study analyzed in Bodnar et al. (2004), we assume that the participants adhere to their randomly assigned treatment up to $T=3$, after which the standard clinical protocol applies.
[…]
Baseline:
[…]
For subsequent follow-up visits $k>0$:
[…]
$A_{2}$ $=A_{1}=A_{0}$ $A_{k}\sim\operatorname{Bernoulli}(\operatorname{expit}(10-L_{k}+0.1\,A_{k-1}))\text{ for }k=3,4.$
We choose a logistic model for a discrete-time hazard marginal structural model:
$\mathbb{P}[Y=1\,|\,do(\bar{A}_{4}),B,C]=\operatorname{expit}(-2+0.1\,B+0.02\,C-\beta\,\operatorname{cum}(\bar{A}_{4})),$
where $\operatorname{cum}(\bar{A}_{4})=\Sigma_{i=1}^{4}A_{i}$. This assumes $\exp(\beta)$ as the causal odds ratio of anemia for an incremental dose. If all women receive iron supplements at all four visits, the causal odds ratio would be $\exp(4\beta)$, representing the true value for the “always treated” efficacy estimand.
We use a t-copula $T_{\rho,\operatorname{df}}$ (with $\rho=-0.5$ and $\operatorname{df}=5$) across all pair-copula construction components to model the dependency between $Y$ and each of $L_{0},L_{1},L_{2},L_{3}$, and $L_{4}$. Compared to a Gaussian copula, the t-copula allows for heavier-tailed dependency, capturing stronger associations in the tails. In our example, t-copulas reflect the likelihood that extreme Hb levels (e.g., severely low) are more strongly associated with extreme outcomes (anemia at delivery). This dependency structure may better reflect the underlying association than a Gaussian copula. Our algorithm offers full flexibility, allowing users to choose any copula family and specify different parameters for each bivariate copula in the pair-copula construction.
We simulated scenarios with varying causal coefficients $\beta=\{-0.5,0.3,0\}$, representing strong, moderate and no effect, respectively and considered different samples size $n=\{500,1000,2000\}$. For each combination of $\beta$ and $n$ we generated 400 datasets. We fitted the model $Y=\operatorname{expit}(\alpha+\gamma\,B+\theta\,C-\beta\,\operatorname{cum}(\bar{A}_{4}))$ using two approaches: with stablized inverse probability of treatment weights, and an unweighted approach. For each fit, we computed 95% confidence intervals for $\beta$ using two methods: (1) the robust sandwich estimator of the variance, with rationale provided by Hernán et al. (2000) that the use of weights induces within-subject correlation, and (2) non-parametric bootstrap, constructing 95% confidence intervals from the 2.5% and 97.5% percentiles from 200 bootstrap samples.
Table 1 presents the bias with corresponding Monte Carlo standard errors. The inverse-probability-weighted estimator shows minimal bias, with Monte Carlo standard errors decreasing as sample size increases.In contrast, the unweighted estimator remains significantly biased across sample sizes, showing no convergence. Table 2 reports coverage rates. Both the sandwich and bootstrap CIs approach 95% coverage, though with minor under-coverage. The unweighted estimator, however, exhibits poor coverage due to significant bias in $\beta$. Table 2 shows the statistical power of the test of treatment effect (null hypothesis: $\beta=0$) for each estimator, with power of the weighted estimator increasing as the causal effect size and sample size grow. In summary, these results confirm that data generated using our simulation algorithm accurately reflect the intended causal effects.
$\beta$ $n$ IPTW Unweighted
--0.5 500 0.027 (0.186) 0.270 (0.115)
1000 0.012 (0.140) 0.269 (0.079)
2000 0.009 (0.096) 0.274 (0.055)
--0.3 500 0.015 (0.179) 0.236 (0.103)
1000 0.004 (0.119) 0.234 (0.073)
2000 0.009 (0.088) 0.235 (0.051)
0 500 0.017 (0.138) 0.189 (0.083)
1000 0.015 (0.087) 0.188 (0.059)
2000 0.014 (0.072) 0.193 (0.043)
Table 1: Bias and Monte Carlo standard errors (in parentheses, defined as $sd(\hat{\beta})$) for inverse-probability-of-treatment-weighted (‘IPTW’) and unweighted estimators (‘unweighted) across different $\beta$ and $n$ values.
$\beta$ $n$ IPTW Unweighted
Sandwich Bootstrap
$-0.5$ 500 91.8% 93.0% 31.8%
1000 92.0% 91.8% 7.5%
2000 92.3% 92.8% 0.0%
$-0.3$ 500 91.0% 91.5% 35.3%
1000 94.3% 92.5% 8.0%
2000 94.3% 93.8% 0.3%
0.0 500 92.8% 93.0% 41.8%
1000 95.5% 94.3% 12.0%
2000 92.8% 90.0% 0.5%
Table 2: Coverage of 95% confidence intervals for IPTW-sandwich, IPTW-bootsrap, and unweighted estimators across scenarios with varying $\beta$ and $n$ values.
$\beta$ $n$ IPTW Unweighted
Sandwich Bootstrap
$-0.5$ 500 83.3% 79.5% 55.3%
1000 97.5% 97.8% 82.5%
2000 100.0% 100.0% 98.3%
$-0.3$ 500 45.8% 43.0% 12.0%
1000 80.0% 77.8% 15.3%
2000 95.8% 96.5% 26.8%
0.0 500 7.3% 7.0% 58.3%
1000 4.5% 5.8% 88.0%
2000 7.3% 10.0% 99.5%
Table 3: Statistical power of the test for the null hypothesis $\beta=0$, i.e., no causal effect.
6 Simulating Survival Outcomes
[…]
Figure 3: Example directed acyclic graph showing temporal relationships in a longitudinal survival setting. The sequence of outcomes observed over time is denoted as $\{Y_{k}\}_{k\in\mathcal{K}^{+}}$, where $\mathcal{K}^{+}=\{1,\dots,K+1\}$.
We have described and demonstrated our method to simulate longitudinal data where the outcome is measured at the final time point. Since time-to-event and survival outcomes are very common in medical and epidemiological contexts, we shall now demonstrate how our method can be extended to simulate longitudinal data from survival models.
As shown in Figure 3, we consider the setup described in Section 4 to remain the same except that the outcome now becomes a sequence of outcomes observed over time as $\{Y_{k}\}_{k\in\mathcal{K}^{+}}$, where $\mathcal{K}^{+}=\{1,\dots,K+1\}$ and $Y_{k}=1$ represents failure. An example study that follows this setup would be one evaluating the effect of an anti-diabetic drug on cardiovascular risk. The primary endpoint is time to the first occurrence of a major adverse cardiovascular event (MACE) [^2] A composite endpoint of death from cardiovascular causes, non-fatal myocardial infarction, or non-fatal stroke.. The baseline variables include sex and medical history, and an example time-varying confounder might be HbA1c; the treatment influences HbA1c which in turn predicts intercurrent events—for example, use of additional medication—and MACE. We will refer to this example as we explain the algorithm and assumptions in this section.
Similarly, we describe the algorithm for the case where the time-varying confounder is univariate in the absence of any latent procedures. It is fairly straightforward to extend to multiple confounders and latent variables in a manner analogous to ordinary longitudinal outcomes (see Section 4.4).
Depending on the chosen marginal structural model, there are various ways to model the outcomes $Y_{k}$’s. Examples include:
  • Discrete-time Cox marginal structural model, where we model the binary survival outcome $Y_{k}\in\{0,1\}$ directly:
    $\frac{P(Y_{k+1}=1\,|\,do(\bar{a}_{k}),Z,Y_{k}=0)}{P(Y_{k+1}=1\,|\,do(\bar{0}_{k}),Z,Y_{k}=0)}=f(\bar{a}_{k},Z;\alpha).$ (17)
  • Continuous-time hazard MSM, where we model survival time $T$ and observe the binary survival outcome $Y_{k}=I(T>k)$. Specifically
    • Cox proportional hazard: $\lambda(t\,|\,do(\bar{a}_{k}),Z)=\lambda(t\,|\,do(\bar{0}_{k}),Z=z_{0})\exp\{g(\bar{a}_{t},Z;\beta)\}$,
    • Additive hazard: $\lambda(t\,|\,do(\bar{a}_{k}),Z)=\lambda(t\,|\,do(\bar{0}_{k}),Z=z_{0})+g(\bar{a}_{t},Z;\gamma)$.
In Section 6.2 we illustrate the simulation of time-to-event outcomes from a Cox marginal structural model.
6.1 Simulation algorithm
The general idea is very similar to the non-survival case and the difference is that in the time-step simulation, the simulation of step $k$ is conditional on survival up to the previous time step.
Unlike the derivation of outcome simulation under the frugal parameterization for the non-survival case discussed in Section 4.2, we now need to condition all quantities involved in the copula on survival up to the previous time interval. For simulating survival outcome at the end of the $k$th interval—that is, $y_{k+1}$—(14) becomes:
$p(y_{k+1}$ $\,|\,z,\bar{l}_{k},\bar{a}_{k},y_{k}=0)$
$=\prod_{j=0}^{k}c^{*}_{Y_{k+1},L_{j}\mid\bar{L}_{j-1},Z,\bar{A}_{k},Y_{k}=0}\big{\{}F(y_{k+1}\,|\,\bar{l}_{j-1},z,do(\bar{a}_{k}),Y_{k}=0),$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},z,do(\bar{a}_{k}),Y_{k}=0);\Theta_{k,j}\}.$ (18)
It is important to note the distinction between the subscripts $j$ and $k$ in the product. To reiterate, our goal is to simulate the outcome at the end of the $k$th interval, that is the $k$th panel in Figure 3. Each copula in (18) associates the outcome $Y_{k+1}$ with all the prior values of the time-varying confounder $L_{j}$, for $j=0,1,\dots,k$.
In the non-survival case discussed in Section 4.2, the quantiles, $F(l_{j}\,|\,\bar{l}_{j-1},z,\bar{a}_{j-1})$, for $j=0,1,\dots,k$, are directly available at each simulation at step $j$ and can be simply plugged into the iterative application of inverse $h$-functions without further computation. However, this convenience does not extend when the quantiles are conditioned on survival at later time points. For instance, suppose that at $k=0$, we simulate baseline health, $P(L_{0})\sim N(0,1)$. Conditioning on survival after 12 months, however, alters the distribution. Survivors’ baseline health likely represents a healthier subgroup than the initial population, meaning the distribution is no longer standard normal. This means that the required conditional quantiles of $L_{k}$ are not directly available and require additional computation. Fortunately, employing copulas to define the dependence structure offers an efficient method to derive these conditional distributions, eliminating the need for resource-intensive techniques such as Monte Carlo approximation or numerical integration.
The proposed simulation algorithm follows an iterative structure and is formally presented and validated using an inductive approach.
Proposed simulation method: At time step $k$, we specify the following inputs:
$F(L_{k}\,|\,\bar{L}_{k-1},Y_{k}=0)$ $=w_{k,1}$
$F(A_{k}\,|\,\bar{L}_{k},Y_{k}=0)$ $=w_{k,2}$
$F(Y_{k+1}\,|\,\bar{L}_{k},do(\bar{A}_{k}),Y_{k}=0)$ $=w_{k,3},$ (19)
where $w_{k,1},w_{k,2},w_{k,3}\sim U(0,1)$. We also have parameters $\Theta_{k,j}$ for each copula $C_{Y_{k+1}L_{j}\mid\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0}$ and from the previous step we assume that—by induction—we are given quantiles for each $L_{j}\mid\overline{L}_{j-1},\overline{A}_{j-1},Y_{k-1}=0$. These inputs include the distributions and quantiles of ‘the past’ (i.e. $L_{k}$ and $A_{k}$) given survival, the quantile of the the outcome, $Y_{k+1}$, given the past, and $k+1$ bivariate copulas that ‘weave’ the dependency between the outcome $Y_{k+1}$ and the history of $L$. Crucially, we do not need to compute the full conditional distribution of $Y_{k+1}L_{k}\mid\bar{L}_{k-1},do(\bar{A}_{k}),Y_{k}=0$, but rather just a quantile corresponding to our simulated value; this is central to the efficiency of our approach. We then compute the following for each $j=0,\ldots,k-1$:
$F(L_{j}\,|\,\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0)$ $=h\big{\{}F(L_{j}\,|\,\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k-1}=0),$
$\qquad\quad F(Y_{k}=0\,|\,\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k-1}=0);\Theta_{k-1,j}\big{\}}$ (20)
$F(Y_{k+1}\,|\,\bar{L}_{j},do(\bar{A}_{k}),Y_{k}=0)$ $=h^{-1}\big{\{}F(Y_{k+1}\,|\,\bar{L}_{j+1},do(\bar{A}_{k}),Y_{k}=0),$
$\qquad\qquad F(L_{j+1}\,|\,\bar{L}_{j},do(\bar{A}_{k}),Y_{k}=0);\Theta_{k,j+1}\big{\}},$ (21)
before
$Y_{k+1}=F^{-1}\big{\{}F(Y_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0)\big{\}}.$ (22)
Equation (20) involves $k$ calculations of the distributions of each $L_{j}\,|\,\bar{L}_{j-1}$ among those who survived to time $k$, in which all the quantities, including quantiles and copulas, are available from the previous step $k-1$. In (21) we iteratively apply $k+1$ $h$-functions to obtain the marginal quantile of $F(Y_{k+1}\,|\,Y_{k}=0)$. Once this quantile is computed, we simulate $Y_{k}$ from the target marginal structural model via inversion, as in (22). In total we evaluate $42k+$ 4 functions (including the simulation of $L_{k}$ and $A_{k}$) for this step.
We proceed to show that the proposed method works using mathematical induction.
Proposition 6.1.
Given the quantiles $w_{k,1},w_{k,2},w_{k,3}$ for each $k=0,\ldots,K-1$ from (19) and associated univariate distributions, together with copula parameters $\Theta_{k,j}$ for each $k=0,\ldots,K$ and $j=0,\ldots,k$, we can use the procedure outlined above to simulate a survival outcome according to the specified distributions.
Proof.
Base case: For $k=0$, initialize the algorithm by specifying:
$F(L_{0})=w_{0,1}$
$F(A_{0})=w_{0,2}$
$F(Y_{1}\,|\,do(A_{0}),L_{0})=w_{0,3}$
$C_{Y_{1}L_{0},do(A_{0})}:\Theta_{0,0}.$
Using these inputs, we calculate
$F(Y_{1}\,|\,do(A_{0}))$ $=h^{-1}\big{\{}F(Y_{1}\,|\,L_{0}),F(L_{0});\Theta_{0,0}\big{\}}$ (23)
$Y_{1}$ $=F^{-1}\big{\{}F(Y_{1}\,|\,do(A_{0})\big{\}}.$
For the next time step ($k=1$), the algorithm requires:
$F(L_{0}\,|\,do(\bar{A}_{1}),Y_{1}=0)$ $=h\big{\{}F(L_{0}),F(Y_{1}=0\,|\,do(A_{0}));\Theta_{0,0}\big{\}}$ (24)
$F(Y_{2}\,|\,L_{0},do(\bar{A}_{1}),Y_{1}=0)$ $=h^{-1}\big{\{}F(Y_{2}\,|\,\bar{L}_{1},do(\bar{A}_{1}),Y_{1}=0),F(L_{1}\,|\,L_{0},do(\bar{A}_{1}),Y_{1}=0);\Theta_{1,1}\big{\}}$
$F(Y_{2}\,|\,Y_{1}=0)$ $=h^{-1}\big{\{}F(Y_{2}\,|\,L_{0},do(\bar{A}_{1}),Y_{1}=0),F(L_{0}\,|\,do(\bar{A}_{1}),Y_{1}=0);\Theta_{1,0}\big{\}}.$
In (24) we derive the distribution of $L_{0}$ among survivors at $k=1$, using $F(Y_{1}=0)$, which is already computed in (23) in step $k=0$. All other necessary quantities are specified at $k=1$. Therefore, the base case is satisfied since the algorithm simulates the first time step and provides the required inputs for the next.
Inductive step: Assumes that we have reached the $k$th iteration of the algorithm. We are given $F(L_{k}\mid\overline{L}_{k-1},do(\bar{A}_{k-1}),Y_{k}=0)$ and $F(Y_{k+1}\mid\overline{L}_{k},do(\bar{A}_{k}),Y_{k}=0)$, as well as a collection of pair copulas between $Y_{k+1}$ and $L_{j}$ with parameters $\Theta_{k+1,j}$. We need to compute the updated distributions $F(L_{j}\mid\overline{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0)$ for each $j<k$, given (from the $(k-1)$th iteration) $F(L_{j}\mid\overline{L}_{j-1},do(\bar{A}_{k-1}),Y_{k-1}=0)$ and $F(Y_{k}\mid\overline{L}_{j},do(\bar{A}_{k-1}),Y_{k-1}=0)$. This can be achieved by noting that $F(L_{j}\mid\overline{L}_{j-1},do(\bar{A}_{k-1}),Y_{k-1}=0)=F(L_{j}\mid\overline{L}_{j-1},do(\bar{A}_{k}),Y_{k-1}=0)$, and then applying (20).
We must then obtain $F(Y_{k+1}\mid\overline{L}_{j},do(\bar{A}_{k}),Y_{k}=0)$ for each $j$ to use in the next iteration. This can be done using (21), and hence, we can move to the $(k+1)$th iteration. By the principle of induction, the statement is true for all $k\geq 0$. ∎
In this basic setting, simulating $K$ time steps for each individual requires $\sum_{k=0}^{K-1}(2k+4)=K^{2}+3K$ calculations. This implies that the computational complexity grows quadratically with $K$. If an individual ceases to survive at any time step, the simulation is terminated for them.
Building upon this basic case, we analyze various extensions:
  • Multiple time-varying confounders: If the number of time-vary confounders increases, making $\bm{L}$ a vector of $p$ variables, the number of calculations for the distributions of ¯$\bm{\bar{}}L_{k-1}$, and $F(Y_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0)$ among the survivors at $k$ scales proportionally with $p$. The total number of computations becomes $\sum_{k=0}^{K-1}(2pk+3p+2)=pK^{2}+(2p+2)K$. Hence the computational complexity grows linearly with $p$. In the demonstration presented in the next subsection, we simulate data where there are two time-varying confounders.
  • Multiple outcomes: The approach to simulating additional outcomes depends on whether they are are survival-related or non-survival-related.
    • For non-survival outcomes, such as weight loss, simulated from marginal structural models – as opposed to from conditional distributions) – the number of calculations to derive $F(\bm{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y^{\triangle}_{k}=0)$ increases proportionally with the number of outcomes, $m$. Here, $\bm{Y}_{k}$ is a m-dimensional vector, and $Y_{k}^{\triangle}$ denotes the primary survival outcome. Specifically, the total number of calculations becomes $\sum_{k=0}^{K-1}(pmk+pk+pm+p+m+1)=\frac{mp+p}{2}K^{2}+(mp+m+p+1)K$.
    • For survival outcomes, such as competing risks (e.g., death from other causes), a hierarchical simulation approach is adopted. This involves defining a hypothetical sequence of events, where any order can be chosen, though some may be more intuitive than others. For instance, the simulation first evaluates death from the primary cause; if an individual survives, it then assesses death from other causes. The distribution of ¯$\bm{\bar{}}L_{k-1}$ is calculated sequentially conditional on surviving from both risks according to this order. As a result, the number of calculations scales with $m$. Additionally, similar to the non-survival case, deriving $(F(\bm{Y}_{k+1}\mid do(\bar{A}_{k}),\bm{Y}_{k}=0)$ requires $m$–fold more computations. In total, $\sum_{k=0}^{K-1}(pmk+pmk+pm+p+m+1)=mpK^{2}+(mp+m+p+1)K$ calculations are needed.
    In both scenarios, the computational complexity grows linearly with the number of outcomes, $m$.
  • Censoring: It is customary to specify and simulate censoring $M_{k}$ using a full conditional distribution $F(M_{k+1}\,|\,\bar{L}_{k},do(\bar{A}_{k}),\bm{Y}_{k}=0)$, which represents the likelihood of an individual being censored given their past history. This requires adding one additional calculation at the end of each time step to simulate $M_{k+1}$.
These calculations are all evaluations of $h$-functions and their inverses. For commonly used copulas with analytic $h$-functions, these evaluations are generally efficient and computationally fast.
6.2 Demonstration
We use an example to demonstrate the proposed algorithm simulating longitudinal time-to-event data. This setup broadly follows the simulation study in Seaman and Keogh (2024).
We consider baseline variables $B$, $C$ and two time-varying covariates $L$ and $W$. $A_{k}$ indicates treatment at follow-up $k$. Let the time to event be $T$ and $Y_{k}=I(T>k)$ be the survival status observed at each visit. Specifically we choose that at baseline:
$B\sim\operatorname{Exp}(2)$ $L_{0}\sim\operatorname{Gamma}(k=1+0.5\,B+0.5\,C,\;\theta=1)$
$C\sim\operatorname{Bernoulli}(0.7)$ $W_{0}\sim\operatorname{Bernoulli}(\operatorname{expit}\{-0.2+0.5\,B+0.5\,C\})$
We specify that throughout the treatment assignment is influenced by $B$, $C$, $L_{k}$ and $W_{k}$ as:
$A_{k}\sim\operatorname{Bernoulli}(\operatorname{expit}\{-1.5+0.5\,B+0.5\,C+0.5\,L_{k}+0.5\,W_{k}\}).$
For subsequent follow-up visits $k>0$:
$L_{k}$ $\sim\operatorname{Gamma}(k=1+0.5\,B+0.5\,C+0.1\,L_{k-1}-0.5\,A_{k-1},\;\theta=1)$
$W_{k}$ $\sim\operatorname{Bernoulli}(\operatorname{expit}\{-0.2+0.5\,B+0.5\,C+W_{k-1}-0.6\,A_{k-1}\}).$
We assume a Cox marginal structural model given by:
$\lambda(t\,|\,do(\bar{A}_{\left\lfloor t\right\rfloor}),C)=\lambda_{0}\exp\left(\beta_{C}\,C+\beta_{A}\,A_{\left\lfloor t\right\rfloor}+\beta_{CA}\,C\,A_{\left\lfloor t\right\rfloor}\right),$ (25)
where $\lambda_{0}$ is the baseline hazard and is assumed to be constant over time, $A_{\left\lfloor t\right\rfloor}$ represent the treatment status at the most recent visit prior to time $t$ (with $\left\lfloor t\right\rfloor$ denoting the largest integer less than $t$). We set $\lambda_{0}=\exp(-2)$ and $\beta_{C}=-0.5$. We vary the values of $(\beta_{A},\beta_{CA})$ to construct scenarios with different strengths of treatment: a strong effect scenario $(-0.5,0.3)$ and a weak effect scenario $(0.2,0.1)$. This parameterization assumes a constant hazard rate between follow-up visits, influenced by the treatment status at the most recent visit. This implies that given survival at follow-up time $k$, the survival time $T$ follows an exponential distribution with the rate equal to the hazard above:
$T-k\,|\,T>k\sim\operatorname{Exp}(\lambda(t\,|\,do(\bar{A}_{\left\lfloor t\right\rfloor}),C)).$
If $T-k<1$ then $Y_{k+1}=1$, and 00 otherwise. We set the total number of follow-ups $K$ to 6. For simplicity we set the copulas that weave the dependence between $L$, $W$ and $Y$ so those that describe the dependency between $L_{j}$ and $Y_{k+1}$ for $j=0,1,2,\dots,k$ are Student t with parameter $\rho_{L}$ and 5 degrees of freedom, and those for $W_{j}$ and $Y_{k+1}$ (given $L_{k}$) for $j=0,1,2,\dots,k$ are Gaussian with correlation $\rho_{W}$. We vary the values of $\rho_{L}$ and $\rho_{W}$ to construct scenarios with different strengths of confounding. Table 4 summarizes the different parameter values. In total, we consider 6 scenarios corresponding to the combination of different effect size and confounding scenarios, and three sample sizes: $n=\{1000,2000,5000\}$.
Effect size Large $(\beta_{A},\beta_{CA})=(-0.5,0.3)$ Small $(\beta_{A},\beta_{CA})=(0.2,0.1)$ Level of confounding High $(\rho_{L},\rho_{W})=(0.9,0.7)$ Medium $(\rho_{L},\rho_{W})=(-0.5,0.4)$ Low $(\rho_{L},\rho_{W})=(-0.1,-0.2)$
Table 4: Parameter values of treatment effect size and confounding scenarios.
Remark 6.2.
The proposed algorithm, as previously discussed, exclusively relies on evaluating analytic functions, making it highly efficient. Table 5 reports the runtime required to simulate data under the setup for the large causal effect with medium confounding scenario. The simulations were carried out on a local laptop using a single score without parallelization. Generating data for $n$ up to 10,000 units takes only a few seconds, while even for a very large sample size ($n=50,\!000$) and 11 time steps ($K=10$), the runtime remains under one minute. This computational efficiency is crucial for tasks such as method evaluation and sample size calculations in clinical trial design, where simulations must be repeated across various parameter values. For example, determining the sample size might require testing 10 different parameter combinations (such as effect sizes, prevalence of drop-in treatments, or adherence levels) across four sample sizes, each bootstrapped 200 times to estimate statistical power. This requires $10\times 4\times 200=8,\!000$ simulations. Using our method, suppose $n=10,000$ and $K=10$, the total runtime is approximately 13 hours on a single core, which can be reduced to about 2 hours using an 8-core laptop or under 10 minutes on a 50-core CPU through parallelization.
In contrast, simulation methods requiring computationally intensive operations, such as Monte Carlo approximations (see, for example, Seaman and Keogh, 2024) or numerical integration, are often too slow to support such large-scale simulations. This highlights a key advantage of our proposed method: its ability to handle extensive simulation demands efficiently, making it well-suited for practical applications.
$n$ $K=5$ $K=10$ 1,000 0.27 0.79 2,000 0.45 1.68 5,000 1.69 2.49 10,000 2.70 6.09 50,000 16.50 45.49
Table 5: Runtime (in seconds) to simulate a dataset of size $n$, averaged over 20 repeats.
We contrast the results of two approaches for fitting the Cox marginal structural model specified in (25). The first employs stabilized inverse probability of treatment weighting, with weights for subject $i$ at time step $k$ calculated as:
$\prod_{j=0}^{k}\frac{\hat{p}_{j}(A_{j}\mid C,\bar{A}_{j-1},Y_{j}=0)}{\hat{p}_{j}(A_{j}\mid B,C,\bar{L}_{j},\bar{W}_{j},\bar{A}_{j-1},Y_{j}=0)},$
where $\hat{p}_{j}(A_{j}\mid C,\bar{A}_{j-1},Y_{j}=0)$ and $\hat{p}_{j}(A_{j}\mid B,C,\bar{L}_{j},\bar{W}_{j},\bar{A}_{j-1},Y_{j}=0)$ are estimated using logistic regression models fitted to the survivors at time step $j$.
Remark 6.3.
The denominator in the weight formula represents the correctly specified propensity score model under the assumed data-generating process, while the numerator stabilizes the weights. To avoid unintended confounding, the baseline covariates included in the numerator must also be adjusted for in the marginal structural model (Cole and Hernán, 2008). In accordance with the model in (25), we only include $C$ in the numerator for stabilization.
The other method is to fit model (25) nominally, without weights or any adjustment for confounding. Table 6 summarizes the bias of estimates produced by inverse probability of treatment weighted and unweighted Cox regressions, from 500 replications. The biases are compared with their corresponding Monte Carlo standard errors as a measure of uncertainty. As the sample size increases, generally, the biases of estimates for $\beta_{C}$, $\beta_{A}$ and $\beta_{AC}$ obtained with inverse probability weights converge to zero, with their absolute values substantially smaller than the corresponding Monte Carlo standard deviations. In contrast, the unweighted estimates exhibit considerable bias, particularly for the estimate of the treatment effect, $\beta_{A}$, where the bias does not diminish as $n$ increases to $5000$. The magnitude of this bias far exceeds the Monte Carlo standard errors. The unweighted estimates are most biased under the high level of confounding by $L$ and $W$, since this is not adjusted for.
Level of confounding Effect size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ IPTW Unweighted IPTW Unweighted IPTW Unweighted High $1000$ $-0.04(2.27)$ $0.27(3.09)$ $3.73(1.73)$ $-1.05(1.54)$ $0.49(2.22)$ $2.31(9.34)$ $\beta_{A}=-0.5$ $2000$ $-0.07(1.68)$ $0.02(2.09)$ $3.81(1.21)$ $-1.03(1.06)$ $0.34(1.51)$ $1.63(9.52)$ $5000$ $0.02(1.25)$ $-0.09(1.58)$ $3.89(0.76)$ $-0.96(0.68)$ $0.24(0.97)$ $1.08(9.44)$ $1000$ $-0.03(2.35)$ $0.03(2.78)$ $3.64(1.46)$ $-0.93(1.59)$ $0.44(1.94)$ $1.99(9.62)$ $\beta_{A}=0.2$ $2000$ $-0.11(1.60)$ $0.17(2.02)$ $3.65(1.18)$ $-0.93(1.11)$ $0.49(1.51)$ $1.49(9.46)$ $5000$ $-0.03(1.14)$ $0.07(1.36)$ $3.65(0.71)$ $-0.95(0.71)$ $0.51(0.89)$ $1.01(9.38)$ Medium $1000$ $-0.07(2.04)$ $0.23(2.91)$ $1.90(1.73)$ $-0.43(1.51)$ $0.08(2.22)$ $2.19(9.48)$ $\beta_{A}=-0.5$ $2000$ $-0.02(1.60)$ $0.09(2.27)$ $1.95(1.22)$ $-0.35(1.01)$ $-0.01(1.55)$ $1.65(9.52)$ $5000$ $-0.09(1.10)$ $0.07(1.59)$ $1.97(0.80)$ $-0.42(0.64)$ $-0.02(1.03)$ $1.21(9.32)$ $1000$ $-0.12(2.06)$ $0.14(2.70)$ $1.75(1.54)$ $-0.36(1.43)$ $0.09(1.92)$ $2.02(9.50)$ $\beta_{A}=0.2$ $2000$ $-0.10(1.62)$ $0.09(2.09)$ $1.68(1.11)$ $-0.42(1.08)$ $0.17(1.44)$ $1.49(9.36)$ $5000$ $-0.03(1.04)$ $0.03(1.32)$ $1.78(0.68)$ $-0.39(0.66)$ $0.14(0.88)$ $0.95(9.54)$ Low $1000$ $0.01(1.82)$ $0.21(2.90)$ $-1.95(1.83)$ $0.46(1.29)$ $-0.20(2.20)$ $2.36(9.50)$ $\beta_{A}=-0.5$ $2000$ $-0.10(1.32)$ $0.07(2.13)$ $-1.83(1.25)$ $0.44(0.89)$ $-0.30(1.59)$ $1.69(9.58)$ $5000$ $-0.02(1.04)$ $0.03(1.46)$ $-1.77(0.84)$ $0.47(0.58)$ $-0.32(0.99)$ $1.07(9.42)$ $1000$ $0.04(1.98)$ $-0.03(2.65)$ $-1.88(1.46)$ $0.53(1.30)$ $-0.31(1.88)$ $1.98(9.64)$ $\beta_{A}=0.2$ $2000$ $-0.02(1.45)$ $0.01(1.87)$ $-1.85(1.10)$ $0.43(0.99)$ $-0.26(1.38)$ $1.43(9.44)$ $5000$ $-0.01(0.92)$ $-0.06(1.25)$ $-1.89(0.68)$ $0.43(0.59)$ $-0.25(0.84)$ $0.95(9.50)$
Table 6: Bias and Monte Carlo standard errors (×$\times 10$) of estimates produced by inverse probability of treatment weighted (‘IPTW’) and unweighted Cox regressions, under different strengths of confounding, effect sizes, and sample sizes.
Table 7 presents the coverage rates of the 95% confidence intervals for both the inverse probability weighted and unweighted estimators. Consistent with the illustrative example in Section 5, the confidence intervals for the inverse probability weighted estimator were computed using two methods:one based on sandwich variance estimation and the other on bootstrap resampling. The results indicate that the confidence intervals for the inverse probability weighted estimator, obtained using both methods, achieve coverage rates close to the nominal 95% level. However, slight undercoverage is observed with both approaches, and the extent of undercoverage is comparable between them. In contrast, the sandwich-based confidence intervals for the unweighted estimator perform poorly, particularly for the treatment effect estimate, $\beta_{A}$.
Level of confounding Effect Size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ W-S W-B UW W-S W-B UW W-S W-B UW High 1000 93.8% 95.2% 88.4% 93.4% 92.8% 42.2% 93.8% 93.8% 93.4% $\beta_{A}=-0.5$ 2000 94.0% 94.0% 84.4% 95.2% 95.6% 12.4% 95.6% 95.6% 94.6% 5000 93.2% 92.8% 70.2% 94.4% 93.4% 0.2% 94.2% 93.6% 93.8% 1000 92.8% 93.8% 91.8% 96.2% 95.8% 31.6% 93.4% 94.0% 95.6% $\beta_{A}=0.2$ 2000 95.2% 95.0% 87.6% 94.6% 95.0% 9.8% 94.8% 94.6% 92.2% 5000 93.6% 93.6% 72.4% 93.8% 93.0% 0.2% 94.0% 94.0% 90.0% Medium 1000 93.8% 94.2% 93.4% 94.8% 93.6% 79.6% 93.8% 94.0% 93.6% $\beta_{A}=-0.5$ 2000 93.0% 93.8% 92.4% 95.2% 94.6% 64.8% 92.2% 93.4% 96.2% 5000 95.4% 94.8% 90.0% 93.2% 92.6% 30.2% 93.0% 93.8% 93.6% 1000 94.2% 94.8% 95.0% 95.0% 93.8% 77.2% 94.2% 93.2% 95.8% $\beta_{A}=0.2$ 2000 94.6% 93.8% 91.6% 93.6% 93.2% 64.0% 92.8% 92.2% 94.2% 5000 95.4% 94.6% 91.6% 95.4% 95.2% 26.8% 94.8% 94.4% 95.4% Low 1000 95.4% 95.4% 93.0% 95.0% 95.2% 84.0% 95.2% 95.2% 94.6% $\beta_{A}=-0.5$ 2000 95.6% 96.4% 93.0% 95.8% 96.0% 71.6% 95.0% 95.6% 94.6% 5000 93.2% 93.2% 87.2% 94.2% 94.8% 42.6% 94.0% 94.4% 94.4% 1000 94.6% 95.6% 94.8% 96.4% 96.6% 77.4% 92.8% 94.8% 96.0% $\beta_{A}=0.2$ 2000 94.8% 95.6% 92.8% 94.4% 94.6% 60.6% 95.0% 95.2% 94.2% 5000 94.8% 94.2% 89.4% 95.0% 95.4% 20.0% 95.0% 94.2% 93.8%
Table 7: Coverage rates for IPTW-Sandwich (W-S), IPTW-Bootstrap (W-B), and unweighted (UW) estimators of $\beta_{C}$, $\beta_{A}$, and $\beta_{AC}$ under different strengths of confounding, effect sizes, and sample sizes.
In summary, the asymptotically unbiased nature of the inverse probability weighted estimates, compared to the biased unweighted estimates, further confirms that the simulated data were generated in accordance with the Cox model in (25), and that confounding by $L$ and $W$ was appropriately introduced. R code for implementing this study is provided in Appendix A.
7 Discussion
In this paper, we proposed and demonstrated a principled strategy for parametrically simulating longitudinal data from marginal structural models, leveraging the frugal parameterization (Evans and Didelez, 2024). Our proposed method is highly flexible and can accommodate a variety of practical scenarios, including multiple treatments (for example, primary and drop-in), dynamic treatment regimes, dose-response analyses, multiple outcomes (such as secondary outcomes and competing risks), and censoring. Additionally, it imposes no restrictions on the data-generating process or marginal structural causal models.
In Section 6, we extended this method to survival models and time-to-event outcomes, which are common in clinical trials. Unlike the non-survival case, survival data simulation requires conditioning each time step $k+1$ on surviving through $k$, meaning all parameterized quantities must account for survival. This posed a challenge: while we can specify the distribution of the time-varying variable, say $L_{j}$, conditional on all prior history up to $j-1$, its distribution conditioning on a future survival state at time $k$ ($k>j$) is generally intractable. A very similar issue occurs in the approach proposed by Seaman and Keogh (2024) where Monte Carlo approximations are used to address the problem. While effective in theory, their approach is computationally expensive and imposes restrictions on the types of data that can be generated. We resolved this challenge elegantly by leveraging pair-copulas, so that these conditional quantiles can be iteratively computed through h-functions. Thus, our solution relies solely on evaluating analytic functions, making it highly efficient. This efficiency is crucial for practical applications, such as sample size calculation, which require repeated simulations across a wide range of parameter combinations.
[…]
If enforcing such Markov conditions is absolutely necessary, an alternative pair-copula construction can be adopted. However, such an approach requires computing marginal quantiles of time-varying confounders through numerical integration, which quickly becomes high-dimensional—a well-known computational challenge.
Appendix A Example R code for Section 6.2
Here is an example code to simulate data following the setup in Section 6.2.
1library(data.table)
2library(copula)
3
4sim_func <- function(n = 500,
5 rho1 = -0.5, rho2 = 2,
6 K = 5, betaS = -0.5,
7 betaB2 = -0.5, betaSB2 = 0.3){
8 # copulas
9 copL <- tCopula(param = rho1, df = 5, dim = 2, dispstr = "un")
10 copW <- normalCopula(param = rho2, dim = 2, dispstr = "un")
11 dat <- data.table()
12 dat\$surv <- rep(1,n)
13
14 # baseline vars
15 dat[, B1 := rexp(n, rate = 2)]
16 dat[, B2 := rbinom(n, size = 1, prob = 0.7)]
17 baseline_vars <- c("B1","B2")
18
19 # a data table to store calculated quantiles to be carried over to the next time step
20 comp_dat <- copy(dat[,baseline_vars, with = F])
21
22 for (k in 0:K){
23 n_surv <- sum(dat\$surv, na.rm = T)
24 if (k>0){
25 comp_dat <- comp_dat[Y == 0,]
26 if (k > 1){
27 comp_dat[,grep("_prev\$", colnames(comp_dat), value = TRUE):= NULL]
28 }
29 # reset names
30 setnames(comp_dat,
31 old = setdiff(names(comp_dat), baseline_vars),
32 new = paste0(setdiff(names(comp_dat), baseline_vars), "_prev"))
33 }
34
35 # random quantiles
36 comp_dat[, paste0("F_L",k,"|surv"):= runif(n_surv)]
37 comp_dat[, paste0("F_W",k,"|L",k,"_surv"):= runif(n_surv)]
38 comp_dat[, paste0("F_S",k,"|L",k,"_surv"):= runif(n_surv)]
39 comp_dat[, paste0("F_Y|LW",k,"_surv"):= runif(n_surv)]
40
41 if (k == 0){
42 comp_dat[,L := qgamma(get(paste0("F_L",k,"|surv")), shape = 1 + 0.5*B1 + 0.5*B2, scale = 1)]
43 comp_dat[,W := qbinom(get(paste0("F_W",k,"|L",k,"_surv")), size = 1, prob = expit(-0.2 +0.5 * B1 + 0.5*B2))]
44 } else{
45 comp_dat[,L := qgamma(get(paste0("F_L",k,"|surv")), shape = 1 + 0.5*B1 + 0.5*B2 + 0.1 * L_prev - 0.5 * S_prev, scale = 1)]
46 comp_dat[,W := qbinom(get(paste0("F_W",k,"|L",k,"_surv")), size = 1, prob = expit(-0.2 + 0.5 * B1 + 0.5*B2 + W_prev - 0.6* S_prev))]
47 }
48
49 comp_dat[,S := qbinom(get(paste0("F_S",k,"|L",k,"_surv")),size = 1, prob = expit(-1.5 + 0.5*B1 + 0.5*B2 + 0.5* L + 0.5 * W))]
50
51 if (k >0){
52 #get the distributions of Ls and Ws in the survivors
53 for (j in (1:k)){
54 if (j<k){
55 comp_dat[,paste0("F_W",k-j,"|L",k-j,"_surv"):= cCopula(as.matrix(cbind(get(paste0("F_Y|L",k-j,"_W", k-j-1,"_surv_prev")),
56 get(paste0("F_W",k-j,"|L",k-j,"_surv_prev")))),copula = copW)[,2]]
57 comp_dat[,paste0("F_L",k-j,"|surv"):= cCopula(as.matrix(cbind(get(paste0("F_Y|LW",k-j-1,"_surv_prev")),
58 get(paste0("F_L",k-j,"|surv_prev")))),copula = copL)[,2]]
59 } else{
60 comp_dat[,paste0("F_W",k-j,"|L",k-j,"_surv"):= cCopula(as.matrix(cbind(get(paste0("F_Y|L",k-j,"_surv_prev")),
61 get(paste0("F_W",k-j,"|L",k-j,"_surv_prev")))),copula = copW)[,2]]
62 comp_dat[,paste0("F_L",k-j,"|surv"):= cCopula(as.matrix(cbind(F_Y_surv_prev,
63 get(paste0("F_L",k-j,"|surv_prev")))),copula = copL)[,2]]
64 }
65 }
66 }
67
68 for (j in (0:k)){
69 if (j<k){
70 comp_dat[,paste0("F_Y|L",k-j,"_W", k-j-1,"_surv"):= cCopula(as.matrix(cbind(get(paste0("F_W",k-j,"|L",k-j,"_surv")),
71 get(paste0("F_Y|LW",k-j,"_surv")))),copula = copW,inverse = T)[,2]]
72 comp_dat[,paste0("F_Y|LW",k-j-1,"_surv"):= cCopula(as.matrix(cbind(get(paste0("F_L",k-j,"|surv")),
73 get(paste0("F_Y|L",k-j,"_W", k-j-1,"_surv")))),copula = copL,inverse = T)[,2]]
74 }else{
75 comp_dat[,paste0("F_Y|L",k-j,"_surv"):= cCopula(as.matrix(cbind(get(paste0("F_W",k-j,"|L",k-j,"_surv")),
76 get(paste0("F_Y|LW",k-j,"_surv")))),copula = copW,inverse = T)[,2]]
77 comp_dat[,F_Y_surv:= cCopula(as.matrix(cbind(get(paste0("F_L",k-j,"|surv")),
78 get(paste0("F_Y|L",k-j,"_surv")))),copula = copL,inverse = T)[,2]]
79 }
80 }
81
82 # hazard
83 comp_dat[, lambda := exp(-2) *exp(betaS * S + betaB2 * B2 + betaSB2*B2*S)]
84 comp_dat[, T := qexp(F_Y_surv, rate = lambda)]
85 comp_dat[, Y := 1*(T<1)]
86
87 dat[surv == 1, paste0("L","_",k) := comp_dat\$L]
88 dat[surv == 1, paste0("W","_",k) := comp_dat\$W]
89 dat[surv == 1, paste0("S","_",k) := comp_dat\$S]
90 dat[surv == 1, paste0("Y","_",k) := comp_dat\$Y]
91 dat[surv == 1, T := (comp_dat\$T + k)]
92 dat[,surv := 1*(get(paste0("Y_",k)) == 0)]
93 return(dat)
94 }

### v2

@@
-Simulating Longitudinal Data from Marginal Structural Models
+Exact Simulation of Longitudinal Data from Marginal Structural Models
 Abstract
-Simulating longitudinal data from specified marginal structural models is a crucial but challenging task for evaluating causal inference methods and designing clinical trials.
-While data generation typically proceeds in a fully conditional manner using structural equations according to a temporal ordering, marginal structural models require capturing causal effects that are marginal over time-dependent confounders, making it difficult to align conditional distributions with target marginal quantities.
+Simulating longitudinal data from specified marginal structural models is a crucial but challenging task for evaluating causal inference methods and clinical trial design.
+While data generation typically proceeds in a fully conditional manner using structural equations according to a temporal ordering, it is difficult to ensure alignment between conditional distributions and the target marginal causal effects.
+This misalignment presents a fundamental challenge in simulating data that adheres to marginal structural model specifications.
 To address this, we propose a flexible and efficient algorithm for simulating longitudinal data that adheres exactly to a specified marginal structural model.
-Recognizing the importance of time-to-event outcomes in clinical trials, we extend the method to accommodate survival models.
-Compared to existing approaches, our method offers several key advantages: it enables exact simulation from a known causal model rather than relying on approximations; avoids imposing restrictive assumptions on the data-generating process; and it is efficient as we need only evaluate analytic functions.
+Recognizing the importance of time-to-event outcomes in clinical research, we extend our approach to accommodate survival models.
+Compared to existing approaches, our method offers several advantages: it enables exact simulation from a known causal model rather than relying on approximations; avoids restrictive assumptions about the data-generating process; and remains computationally efficient by requiring only the evaluation of analytic functions.
 This last benefit contrasts with methods that use computationally intensive techniques such as Monte Carlo approximations or numerical integration.
 Through simulation studies replicating realistic scenarios, we validate the method’s accuracy and utility.
-Our method will allow researchers to effectively simulate data with target causal structures for their specific scenarios.
+Our method will facilitate researchers in effectively simulating data with target causal structures for their specific scenarios.
 1 Introduction
-Longitudinal data are very common in studies analyzing the causal effect of a treatment (or exposure) on an outcome, especially in the context of chronic diseases such as diabetes or HIV.
-However, treatments are rarely fixed for the entire period of observation, and their status often varies over time.
-Even in clinical trials with strict adherence protocols, intercurrent events—such as discontinuation of the prescribed treatment and use of additional medication—are inevitable.
-For example, participants in the treatment group might discontinue treatment due to side effects, while those in the control group could begin treatment or alternative therapies if their health declines.
-In observational studies, treatment initiation or cessation can occur at any time, contingent on changing circumstances.
-The ICH E9111International Council on Harmonization of Technical Requirements for Registration of Pharmaceuticals for Human Use [ICH] Statistical Principles for Clinical Trials [E9] emphasizes that intercurrent events must be explicitly considered when defining the clinical question of interest (European Medicines Agency, 2020).
-While the intention-to-treat (ITT) principle simplifies this to some extent by reducing time-varying treatment to a point-treatment problem, it may not always yield the most meaningful estimand from a clinical or regulatory perspective.
-Longitudinal data, however, allow us to more accurately capture the treatment dynamics, enabling estimation of causal quantities like the “on-treatment” estimand, which represents the treatment effect assuming treatment is administered as intended.
-A particular challenge in analyzing longitudinal data is time-varying confounding.
-This occurs when a time-dependent covariate predicts both the outcome and future treatment exposure, while past treatment influences the covariate.
-For example, in the case study by Robins et al.
+Longitudinal data are fundamental for studies analyzing the causal effect of a treatment or exposure on an outcome, especially for research into chronic diseases such as diabetes or HIV.
+Unlike cross-sectional studies, longitudinal designs capture the dynamic nature of treatment regimens, which often vary over time due to clinical decisions, patient adherence, or external factors.
+Even in clinical trials with strict adherence protocols, intercurrent events—such as treatment discontinuation or supplementary medication use—are common.
+By accounting for these treatment dynamics, longitudinal data enable the estimation of meaningful causal quantities, such as the “on-treatment” estimand, which quantifies the treatment effect assuming treatment is administered as intended.
+A central challenge in analyzing longitudinal data is addressing time-varying confounding, where a time-dependent covariate predicts both the outcome and future treatments, while itself being influenced by past treatments.
+For example, in the case study in Robins et al.
 (2000) on the effect of zidovudine (AZT) treatment on HIV infection, the biomarker CD4 lymphocyte count is both affected by prior AZT treatment and is predictive of future therapy initiation and HIV progression.
 The presence of such time-varying confounding renders conventional adjustment methods, like regression and stratification, generally biased.
-Marginal structural models, first introduced by Robins (1997), have emerged as a powerful tool for estimating the causal effect of time-dependent exposures in the presence of time-varying confounding.
-Model parameters can be consistently estimated using inverse-probability-of-treatment-weighted estimators.
-Among alternative methods, including structural nested models estimated through g-computation (Robins, 1989; Robins and Tsiatis, 1991), a systematic review of the literature by Clare et al.
-(2019) concluded that marginal structural models fitted with inverse probability of treatment weights are the most commonly used approach for longitudinal data with time-varying confounding due to their simplicity in both implementation and interpretation.
-As novel causal inference methods for longitudinal data are continually proposed to address complex real-world scenarios, it is essential to evaluate their effectiveness for practical use.
-Simulation studies play a key role in this, by assessing important properties such as bias, confidence interval coverage, relative efficiency, and robustness to assumption violations.
-These evaluations guide statisticians in developing new methods and help analysts choose appropriate methods for specific research questions.
-Sample size calculation also benefits from simulations.
-Prior to conducting a trial or an observational study, it is crucial to determine the number of participants, or number of observational units required to achieve adequate statistical power for the hypothesis test (for example, superiority, inferiority or equivalence).
-An improperly sized sample is clearly problematic: if too few units are recruited it can lead to undetected non-zero effects; if too many are recruited this may be a waste of resources and potentially unethical if it means that more individuals are given an inferior treatment.
-While standard formulas can be used for time-fixed treatments, more complex scenarios with longitudinal data and time-varying treatments make deriving analytical formulas impractical or impossible.
+Marginal structural models, introduced by Robins (1997), have emerged as a powerful tool for estimating causal effects under time-varying confounding.
+Model parameters can be consistently estimated using inverse probability of treatment weighting (Robins, 1999, 2000).
+Among alternative methods, including structural nested models estimated through g-computation (Robins, 1989; Robins and Tsiatis, 1991), marginal structural models have gained prominence due to their interpretability and ease of implementation.
+A systematic review by Clare et al.
+(2019) highlights their widespread use in applied research.
+As causal inference methods for longitudinal data are continually proposed to address complex real-world settings, simulation studies are essential to evaluate their effectiveness for practical use.
+These studies assess important properties such as bias, confidence interval coverage, relative efficiency, and robustness to violations of the assumptions used, guiding both methodological development and application.
+Sample size calculations also benefit from simulations.
+Prior to conducting a trial or an observational study, it is crucial to determine the number of participants or number of observational units required to achieve adequate statistical power.
+While standard formulas can be used for time-fixed treatments, more complex scenarios with time-varying treatments make deriving analytical formulas impractical or impossible.
 In such cases, simulation is invaluable, allowing researchers to estimate sample sizes based on realistic data-generating processes and causal structures.
-For both method evaluation and sample size calculation, simulating data that adheres to a known causal structure is crucial.
+A fundamental requirement for both method evaluation and sample size determination is the ability to simulate data that conform to a specified causal model.
 However, simulating longitudinal data from a pre-specified marginal structural model is not a straightforward task.
-While generating data sequentially from structural equations—creating covariates, treatment status, and then outcome conditioned on past information—may seem intuitive, marginal structural models rely on marginal, rather than conditional models.
-This makes it difficult to sequentially simulate from conditional models in a way that aligns with a pre-specified marginal model; in particular, it is very difficult to sample from a null hypothesis.
+While generating data sequentially from structural equations—creating covariates, treatment status, and then the outcome conditioned on past information—may seem intuitive, these models rely on conditional rather than marginal specifications.
+This makes it difficult to sequentially simulate with conditional distributions in a way that aligns with a pre-specified marginal model; in particular, it is extremely challenging to sample from a null hypothesis —the so-called g-null paradox (Robins and Wasserman, 1997).
 The remainder of this paper is structured as follows: Section 2 reviews existing methods for simulating longitudinal data from causal models.
-Section 3 covers the preliminaries, including the frugal parameterization (Evans and Didelez, 2024) and copulas.
-In Section 4, we introduce our proposed methodology and illustrate it in Section 5 through a simulation study replicating a real-world study in epidemiology (Bodnar et al., 2004).
+Section 3 introduces key preliminaries, including the frugal parameterization (Evans and Didelez, 2024) and copulas.
+In Section 4, we introduce our proposed methodology and illustrate it in Section 4.4 through a simulation study replicating a real-world study in epidemiology (Bodnar et al., 2004).
 Section 5 extends the simulation approach to time-to-event data, enabling simulations from survival models such as Cox marginal structural models, including an illustrative example.
 The paper concludes with a discussion in Section 6.
+Example code for the simulations in Sections 4.4 and 5.3 is available at https://github.com/XiLinStats/LongSiMSM.
 2 Related Literature
-As noted in the previous section, the challenge lies in ensuring that simulated data adhere to a specified marginal model, rather than simply following sequential conditional distributions in a structural equation model.
-There are two possible strategies to address this issue:
+As noted in Section 1, the challenge lies in ensuring that the simulated data adhere to a specified marginal model, rather than simply following sequential conditional distributions in a structural equation model.
+Existing approaches address this issue through two strategies:
   1. (i)
-    Simulate data from the conditionals and then derive the corresponding marginal models.
-The g-formula, simplifying assumptions, and Monte Carlo approximations, can help bridge the gap between the simulated conditionals and the target marginal model.
+    Simulating from conditional models and deriving or approximating the corresponding marginal models post-hoc.
   2. (ii)
-    Alternatively, bypass conditional simulations and develop methods to simulate directly from the target marginal model.
-Both strategies are explored in existing approaches.
-2.1 Simulate from the conditionals and derive the equivalent marginal models
+    Bypassing conditional simulations and simulating directly from the target marginal model.
+Several studies have explored the first approach.
 Xiao et al.
-(2010) conducted a simulation study to evaluate the accuracy of two methods of fitting a Cox marginal structural model.
-They generated their dataset through the natural approach—sequentially through conditional models—and argued that under the assumption of a rare disease, i.e.
-a very low failure rate, the impact of non-collapsibility is negligible.
-In this scenario, the parameters of the underlying Cox marginal structural model may be approximated from the conditional model parameters.
-Notably, for this approximation to hold, there must be no confounders for the relationship between the time-varying confounder and the outcome, whether observed or unobserved.
-Essentially, their study did not directly simulate data from an marginal structural model but demonstrated a case where conditional simulation can serve as a reasonable approximation of such a model.
-Young and Tchetgen Tchetgen (2014) explored bridging from conditionals to marginals using the g-formula in simulating from a discrete time Cox marginal structural model.
-In addition to standard identifying assumptions—consistency, positivity, and sequential exchangeability—they further assume that the g-formula uses standard parametric models, and hence derived the functional form of the marginal model.
-If the time-varying covariates consist only of discrete variables, then deriving the equivalent marginal relative hazard is theoretically possible, but it can be cumbersome.
-However, the authors noted that if the time-varying covariate includes continuous components, the summation in the above equation is replaced by integrals which may be challenging to evaluate due to the lack of closed-form solutions.
-The authors suggest potential simplifications such as the Markov property or approximation based on the rare disease assumption, as used by Xiao et al.
-(2010).
-A significant limitation arises when attempting to simulate data from a specific marginal structural model.
-Other than in very specific cases, such as for discrete or linear models, it is essentially impossible to reverse-engineer the series of equivalent conditional models.
-For instance, if we want to simulate from $h(t,\bar{a}_{t})=0.1+0.5a_{t}+0.3a_{t-1}-a_{t}a_{t-1}$, the corresponding series of conditionals is difficult to derive.
-This poses a drawback in real-world applications, especially for sample size determination, where experimenting with different effect sizes is necessary to generate power curves.
+(2010) used sequential conditional models to approximate Cox marginal structural models under the rare disease assumption, i.e.
+very low failure rates, which mitigates the impact of noncollapsibility but assumes the absence of unmeasured confounding between time-varying confounders and outcomes.
+Young and Tchetgen Tchetgen (2014) employed the g-formula to bridge conditional and marginal models in discrete-time Cox marginal structural models.
+While theoretically feasible for discrete time-varying covariates, deriving marginal models becomes intractable for continuous covariates due to the need for numerical integration.
+Simplifications, such as the Markov property or rare disease assumptions, are often necessary but limit generalizability.
 Sofrygin et al.
-(2017) developed the simcausal R package as a flexible, practical tool for simulating complex longitudinal data structures from non-parametric structural equation models.
-The package supports simulations that replicate realistic data scenarios common in causal inference, such as time-varying confounding and selection bias.
-Users can specify and simulate counterfactual data under various intervention regimes.
-A typical simcausal workflow involves parameterizing structural equations based on a causal directed acyclic graph and simulating data sequentially from the specified conditional distributions.
-Since deriving analytic expressions for marginal causal quantities, such as the average treatment effect, is often intractable with this type of conditional parameterization, users instead estimate these parameters through Monte Carlo approximations by simulating counterfactual data.
-While this approach is suitable for marginal causal quantities such as the average treatment effect, it may lack consistency for causal structures such as marginal structural models or conditional average treatment effect functions due to potential model misspecification.
-The authors acknowledged this limitation, as they allow users to specify a “working” model and cautioned that this approximation serves as the best parametric fit within the model class, without guaranteeing consistency.
+(2017) developed the simcausal R package for flexible simulation from non-parametric structural equation models.
+The workflow involves specifying structural equations and sequentially simulating data from conditional distributions.
+However, deriving analytic expressions for marginal causal quantities under this conditional parameterization is often intractable and prone to model misspecification.
+To address this, users typically specify low-dimensional working marginal structural models (Neugebauer and van der Laan, 2007) and estimate their parameters via Monte Carlo simulations of counterfactual data.
+This introduces two limitations: first, the true marginal model remains unknown; second, reverse-engineering simulations to match a target causal function is challenging.
 Keogh et al.
-(2021) described how to simulate from an additive-hazard marginal structural model (Aalen, 1989).
+(2021) described simulation from additive-hazard marginal structural models (Aalen, 1989).
 They utilized the result that if the conditional hazard model follows an additive form, the equivalent marginal hazard model will also be additive.
-Unlike Cox hazard ratios, the parameters of the additive hazard model are collapsible.
-Their strategy involved simulating data conditionally and then deriving the corresponding marginal hazard.
-However, while this approach ensures the correct form of the marginal model, the true values of the coefficients remain generally intractable.
-To address this, as with the workflow outlined in Sofrygin et al.
-(2017), the authors find a Monte Carlo approximation of the true parameters through simulating counterfactual data.
-A notable limitation of the additive hazard model is that it does not constrain the hazard to be non-negative.
-This becomes problematic when the hazard depends on continuous covariates, as ensuring positivity may be difficult in simulated data.
-Additionally, as with the approach by Young and Tchetgen Tchetgen (2014), this method does not allow direct specification of the model parameters, limiting its applicability in real-world scenarios.
-2.2 Simulate directly from the target marginal model
-Havercroft and Didelez (2012) approached this problem from a novel angle by simulating directly from the causal margin (in their case, a discrete Cox marginal structural model) and then introducing time-varying confounding to establish its relationships with the outcome and the treatment.
-Specifically, they used an inverse cumulative distribution function to simulate from the causal margin, with time-varying confounding introduced indirectly via a shared latent variable.
-This approach omits any direct edge from the time-varying confounder directly to the outcome.
-The primary limitation of this method is that time-varying confounding is not explicitly modeled but introduced via an artificial latent variable.
-As a result, the assumption that time-varying covariates have no direct effect on outcomes may not always hold in clinical contexts.
-For instance, in the HIV study example, this would assume that CD4 cell count has no direct effect on survival, which is unlikely.
-Nonetheless, their approach presents a valuable contribution.
-Evans and Didelez (2024) recently proposed a frugal parameterization that prioritizes the causal margin and parameterizes the surrounding variables accordingly.
-This innovation enables direct simulation from the marginal structural model, addressing limitations identified in previous work.
-Specifically, it allows for explicit simulation of time-varying covariates and an exact marginal model.
-More details on the frugal parameterization are given in Section 3.1.
-Building on this foundation, Seaman and Keogh (2024) applied the frugal parameterization to simulate longitudinal data with survival outcomes.
-Similar to the approach of Havercroft and Didelez (2012), they simulated the outcome using the inverse cumulative distribution function of the target causal margin.
-To handle high-dimensional time-varying covariates, they summarized their effects into a one-dimensional “risk score”, simplifying the dependence between covariates and the outcome to a bivariate copula.
-However, this approach has notable limitations.
-Since the cumulative distribution functions of the risk score are generally unknown, they relied on Monte Carlo approximations, which involve sampling a large number of individuals with the same treatment history and baseline variables.
-It is a computationally expensive task and must be performed at each time step.
-For example, with binary treatment, there could be up to $10242^{10}=$ 1024 combinations in terms of treatment history distribution functions to approximate at step 10 – potentially even more if baseline variables are included.
-The method becomes infeasible when treatment or baseline variables are continuous.
-The method also becomes impractical when treatment history or baseline variables have many possible combinations or are continuous.
-Moreover, the use of a risk score limits flexibility, as it prevents explicit modeling of specific covariate-outcome dependencies, such as the effect of age on survival.
-A more flexible approach would directly model these associations, making fuller use of the frugal parameterization.
-The literature review highlights gaps in current methodologies and identifies key features desired in an improved simulation approach.
-These include the ability to bypass the conditional distribution of the outcome and simulate directly from the marginal structural model, robustness that avoids assumptions about the data-generating process, flexibility to accommodate diverse scenarios, and practicality for real-world applications.
+While the marginal model retains the correct form, true parameter values are generally intractable and require Monte Carlo approximations.
+A further limitation is the difficulty in ensuring non-negative hazards for continuous covariates.
+Alternative methods focus on simulating directly from the target marginal model.
+Havercroft and Didelez (2012) simulated directly from a discrete Cox marginal structural model, introducing time-varying confounding indirectly via a shared latent variable.
+However, this approach assumes no direct effect of time-varying confounders on outcomes, which may not hold in practice, e.g.
+the effect of CD4 count on HIV survival (Robins et al., 2000).
+Evans and Didelez (2024) proposed a frugal parameterization that prioritizes the causal margin and parameterizes the surrounding variables accordingly, enabling direct simulation from the exact marginal model while explicitly modeling time-varying covariates.
+This addresses key limitations of earlier methods.
+Building on this, Seaman and Keogh (2024) applied the frugal parameterization to simulate survival outcomes, using a one-dimensional “risk score” to summarize multi-dimensional covariates.
+While simplifying dependence structures, this method relies on computationally expensive Monte Carlo approximations to estimate the cumulative distribution functions of the risk score, making it infeasible for continuous treatments or baseline variables.
+The aims for a new simulation approach therefore include the ability to: precisely specify the marginal structural model; include multiple time-varying covariates, treatments, outcomes, and censoring; have good control over their distributions; avoid restrictive assumptions about the data-generating process.
+In addition, we would like it to scale well in the number of time-points and variables, and be applicable to real-world problems.
 3 Preliminaries
@@
 As discussed in the previous section, the fundamental challenge in simulating data from longitudinal marginal structural models arises from the fact that, while the joint data distribution can be easily factorized into conditionals for simulation, the causal quantity of interest is a marginal distribution.
-Earlier methods, as detailed in Section 2, tackle this by deriving the marginal distribution implied by the conditional factorization.
-However, Evans and Didelez (2024) significantly advanced this area with the introduction of the frugal parameterization, which explicitly centers the causal marginal and constructs the rest of the joint distribution around it.
-This breakthrough has made the simulation method presented in this paper possible.
+Earlier methods, as detailed in Section 2, tackled this by deriving the marginal distribution implied by the conditional factorization.
+However, Evans and Didelez (2024) advanced this area with the introduction of the frugal parameterization, which explicitly parameterizes the causal marginal and constructs the rest of the joint distribution around it.
+This development has made the simulation method presented in this paper possible.
 The frugal parameterization decomposes the joint distribution $p(\bm{l},a,y)$ into three separate pieces:
@@
     a dependency measure: $\phi^{*}_{Y\bm{L}\mid A}$—captures the dependency between the outcome $Y$ and the covariates $\bm{L}$ under intervention.
-The first component, termed “the past”, refers to the distribution of covariates and treatment, which precede the outcome.
-The second component, the causal margin, aligns with the marginal structural model we aim to simulate, directly addressing the causal question at hand.
-The third component captures the dependency between the outcome $Y$ and the set of covariates $\bm{L}$ jointly.
-A common choice of this dependency, $\phi$, is a copula, which we will elaborate on in the following subsection.
+A common choice of this dependency measure is a copula, which we elaborate on in the following subsection.
 3.2 Copulas
-The term copula is derived from the Latin verb copulare, meaning “to join together.” We can think of continuous multivariate distributions as the joining together of the univariate marginals (or margins) and the dependence structure.
+We can think of continuous multivariate distributions as the joining together of the univariate marginals (or margins) and the dependence structure.
 This structure can be represented by a copula, which is a multivariate distribution function $C:[0,1]^{d}\rightarrow[0,1]$ with margins uniformly distributed on $(0,1)$.
-According to Sklar’s theorem (Sklar, 1959, 1973), any continuous multivariate distribution $F$ with marginals $F_{1},\dots F_{d}$ can be expressed as
-$F(x_{1},\dots,x_{d})=C\left\{F_{1}(x_{1}),\dots,F_{d}(x_{d})\right\}=C(u_{1},\dots,u_{d}),$
-where $\bm{X}=(X_{1},\dots,X_{d})$ is a vector of random variables, and $u_{i}=F_{i}(x_{i})$, $i=1,\dots,d$. The copula then takes the form
-$C(u_{1},\dots,u_{n})=F\{F_{1}^{-1}(u_{1}),\dots,F_{d}^{-1}(u_{d})\},\qquad(u_{1},\dots,u_{d})\in[0,1]^{d},$
-where $x_{i}=F_{i}^{-1}(u_{i})$.
-The mathematical relationship between copula density function $c(u_{1},\dots,u_{d})$ and joint density function $f(x_{1},\dots,x_{d})$, applying the chain rule, is given by
-$f(x_{1},\dots,x_{d})=c(u_{1},\dots u_{d})\prod_{i=1}^{d}f_{i}(x_{i}).$ (1)
+According to Sklar’s theorem (Sklar, 1959, 1973), any continuous multivariate distribution $F$ with marginals $F_{1},\ldots,F_{d}$ can be expressed as
+$F(x_{1},\ldots,x_{d})=C\left\{F_{1}(x_{1}),\ldots,F_{d}(x_{d})\right\}=C(u_{1},\ldots,u_{d}),$
+where $\bm{X}=(X_{1},\ldots,X_{d})$ is a vector of random variables, and $u_{i}=F_{i}(x_{i})$, $i=1,\ldots,d$. The copula then takes the form
+$C(u_{1},\ldots,u_{d})=F\{F_{1}^{-1}(u_{1}),\ldots,F_{d}^{-1}(u_{d})\},\qquad(u_{1},\ldots,u_{d})\in[0,1]^{d}.$
 The choice of copula can be considered independently of the univariate margins.
 3.2.1 Pair-copula construction
-Building higher-dimensional copulas is generally recognized as a difficult problem.
-While there exist a wider variety of parametric bivariate copulas, the options for higher-dimensional copulas are limited, with Gaussian and Student t-copulas being common choices.
+While there exist a wide variety of parametric bivariate copulas, the options for higher-dimensional copulas are limited, with Gaussian and Student t-copulas being common choices.
 Extending symmetric bivariate Archimedean copulas to multivariate settings is also possible but imposes strong restrictions, such as requiring the same level of dependence among all possible pairs.
-First introduced by Joe (1996), pair-copula constructions provide a flexible and intuitive approach to building multivariate copulas by modeling multivariate data as a series of bivariate copulas.
+Building on the work of Joe (1996), Bedford and Cooke (2001, 2002) introduced pair-copula constructions, a flexible and intuitive approach to constructing multivariate copulas by modeling multivariate data as a series of bivariate copulas.
 This hierarchical structure incorporates additional variables at each level, using pair-copulas as the fundamental building blocks.
@@
 where $\bm{x}$ is a $d$-dimensional vector; $x_{j}$ is an arbitrarily chosen component of $\bm{x}$, and $x_{-j}:=(x_{1},\ldots,x_{j-1},x_{j+1},\ldots,x_{d})$ denotes $\bm{x}$ with its $j$th entry removed.
-Applying (2) recursively to the conditional densities and, for convenience, choosing an order of $1,2,\dots d$ we have:
-$f(y\,|\,\bm{x})$ $=c_{YX_{d}|\bm{X}_{1:{d-1}}}\{F(y\,|\,\bm{x}_{1:{d-1}}),F(x_{d}\,|\,\bm{x}_{1:{d-1}})\}$
-$\quad\cdot c_{YX_{d-1}|\bm{X}_{1:{d-2}}}\{F(y\,|\,\bm{x}_{1:{d-2}}),F(x_{d-1}\,|\,\bm{x}_{1:{d-2}})\}$
-[…]
-$\quad\cdot c_{YX_{2}|X_{1}}\{F(y\,|\,x_{1}),F(x_{2}\,|\,x_{1})\}$
-$\quad\cdot c_{YX_{1}}\{F(y),F(x_{1})\}\cdot f(y)$ (3)
-$=f(y)\cdot\prod_{i=1}^{d}c_{YX_{d+1-i}|\bm{X}_{1:{d-i}}}\{F(y\,|\,\bm{x}_{1:{d-i}}),F(x_{d+1-i}\,|\,\bm{x}_{1:{d-i}})\},$ (4)
-where each copula in the product in (4) is bivariate, i.e.
-a pair-copula.
-In principle, any ordering of the conditioned variables $x_{1},x_{2},\dots,x_{d}$ can be selected, although some orderings may be easier to use than others.
-In Section 4, we choose a particular ordering that is convenient and efficient for simulation.
-In conclusion, this shows that under appropriate regularity conditions, a multivariate density can be expressed as a product of pair-copulas, acting on different conditional distributions.
-3.2.2 Simulating from pair-copula-constructions
+Applying (1) recursively to the conditional densities, we have:
+$f(y\,|\,\bm{x})$ $=f(y)\cdot\prod_{i=1}^{d}c_{YX_{d+1-i}|\bm{X}_{1:{d-i}}}\{F(y\,|\,\bm{x}_{1:{d-i}}),F(x_{d+1-i}\,|\,\bm{x}_{1:{d-i}})\},$ (2)
+where each copula in the product in (2) is bivariate; this defines a pair-copula.
+In principle, any ordering of the conditioned variables $x_{1},x_{2},\ldots,x_{d}$ can be selected, although some orderings may be easier to use than others.
+In our proposed method, we choose an ordering that is convenient and efficient for simulation.
+In conclusion, under appropriate regularity conditions, a multivariate density can be expressed as a product of pair-copulas, acting on different conditional distributions.
+3.2.2 Simulating from pair-copula constructions
 The pair-copula construction includes conditional distributions of the form $F(y\,|\,\bm{x})$. Joe (1996) demonstrated the following result:
 $F(y\,|\,\bm{x})=\frac{\partial C_{YX_{j}\mid\bm{X}_{-j}}\left\{F(y\,|\,\bm{x}_{-j}),F(x_{j}\,|\,\bm{x}_{-j})\right\}}{\partial F(x_{j}\,|\,\bm{x}_{-j})}.$ (3)
-Proof.
-[…]
-$=\frac{\partial}{\partial F(x_{j}\,|\,\bm{x}_{-j})}C_{YX_{j}\mid\bm{x}_{-j}}\{F(y\,|\,\bm{x}_{-j}),F(x_{j}\,|\,\bm{x}_{-j})\}.$
-For convenience of notation, we introduce the h-function $h(x_{1},x_{2};\Theta)$ from Aas et al.
+For notational convenience, we introduce the h-function $h(u_{1},u_{2};\Theta)$ from Aas et al.
 (2009).
-Let $x_{1}$ and $x_{2}$ be copula random variables, i.e.
-with margins following the uniform distribution $x_{i}\sim\operatorname{U}(0,1)$ and $F_{i}(x_{i})=x_{i}$, for $i=1,2$.
-The $h$-function is defined as:
-$h(x_{1},x_{2};\Theta)$ $:=F(x_{1}\,|\,x_{2})=\frac{\partial F_{12}(x_{1},x_{2})}{\partial x_{2}f_{2}(x_{2})}$
-$=\frac{\partial}{\partial F_{2}(x_{2})}C_{12}(F_{1}(x_{1}),F_{2}(x_{2});\Theta)$ (6)
-$=\frac{\partial}{\partial x_{2}}C_{12}(x_{1},x_{2};\Theta),$ (7)
+Let $u_{1}$ and $u_{2}$ be random variables with a uniform distribution on $(0,1)$.
+The $h$-function is defined as
+$h(u_{1},u_{2};\Theta):=F(u_{1}\,|\,u_{2})=\frac{\partial}{\partial u_{2}}C_{12}(u_{1},u_{2};\Theta)$
 where the second argument in $h(\cdot;\Theta)$ is the conditional variable, and $\Theta$ represents the set of copula parameters.
-From (7) we can see that the $h$-function is essentially the conditional copula function, $C(u_{1}\,|\,u_{2};\Theta)$.
-Example 3.1.
-The derivation of the analytical form of $h(\cdot)$ for common bivariate copulas is straightforward following (7) for parametric copula families. Some examples include:
-  • Gaussian copula: $h(x_{1},x_{2};\rho)=\Phi\left(\frac{\displaystyle\Phi^{-1}(x_{1})-\rho\Phi^{-1}(x_{2})}{\displaystyle\sqrt{1-\rho^{2}}}\right)$;
-  • Student’s $t$ copula: $h(x_{1},x_{2};\rho,\nu)=t_{\nu+1}\left(\frac{\displaystyle\sqrt{\nu+1}(t_{\nu}^{-1}(x_{1})-\rho t_{\nu}^{-1}(x_{2}))}{\displaystyle\sqrt{\left(\nu+\left(t_{\nu}^{-1}(x_{2})\right)^{2}\right)\left(1-\rho^{2}\right)}}\right)$;
-  • Clayton copula: $h(x_{1},x_{2};\theta)=x_{2}^{-\theta-1}\left(x_{1}^{-\theta}+x_{2}^{-\theta}-1\right)^{-1-1/\theta}$;
-  • Gumbel copula: $h(x_{1},x_{2};\theta)=C(x_{1},x_{2};\theta)\frac{1}{x_{2}}(-\log x_{2})^{\theta-1}\left((-\log x_{1})^{\theta}+(-\log x_{2})^{\theta}\right)^{1/\theta-1}$.
+The $h$-function is effectively the conditional copula $C(u_{1}\,|\,u_{2};\Theta)$.
+Remark 3.1.
+The derivation of the analytical form of $h(\cdot)$ for common bivariate copulas is straightforward by definition for parametric families.
+Examples include Gaussian, Student t, Clayton, Frank, and Gumbel copulas.
 The $h$-functions are foundational for simulating from $F(y\,|\,\bm{x})$.
-Similar to (4), by imposing an order $1,\dots,d$ on the components of $\bm{x}$, we can recursively apply $h$-functions to obtain:
-$F(y\,|\,\bm{x})$ $=h\{F(y\,|\,\bm{x}_{1:{d-1}}),F(x_{d}\,|\,\bm{x}_{1:{d-1}}),\Theta_{1}\}$
-$=h[h\{F(y\,|\,\bm{x}_{1:{d-2}}),F(x_{d-1}\,|\,\bm{x}_{1:{d-2}};\Theta_{2}\},F(x_{d}\,|\,\bm{x}_{1:{d-1}});\Theta_{1}]$
-$=h\left[h\{h\left(F(y\,|\,\bm{x}_{1:{d-3}}),F(x_{d-2}\,|\,\bm{x}_{1:{d-3}};\Theta_{3})\right),F(x_{d-1}\,|\,\bm{x}_{1:{d-2})};\Theta_{2}\},F(x_{d}\,|\,\bm{x}_{1:{d-1}});\Theta_{1}\right]$
-$\dots$
-$=h[\dots h\{h(F(y),F(x_{1});\Theta_{d}),F(x_{2}\,|\,x_{1});\Theta_{d-1}\}\dots;\Theta_{1}].$ (8)
-Thus, $Y$ given $\bm{X}$ can be simulated by first drawing a uniform $U\sim{\rm U}(0,1)$, and then evaluating:
-$Y=h^{-1}[\dots h^{-1}\{h^{-1}(U,\,F(X_{1});\Theta_{d}),F(X_{2}\,|\,X_{1});\Theta_{d}\}\dots;\Theta_{1}].$ (9)
-This means that to simulate $Y$ given a $d$-dimensional vector $\bm{X}$, we recursively calculate the $d$ inverse $h$-functions, forming a PCC. This forms the basis of our proposed simulation method.
+Similar to (2), by imposing an ordering $1,\ldots,d$ on the components of $\bm{x}$, we can recursively apply $h$-functions to obtain:
+$F(y\,|\,\bm{x})$ $=h\{F(y\,|\,\bm{x}_{1:{d-1}}),F(x_{d}\,|\,\bm{x}_{1:{d-1}}),\Theta_{d}\}$
+$=h[h\{F(y\,|\,\bm{x}_{1:{d-2}}),F(x_{d-1}\,|\,\bm{x}_{1:{d-2}};\Theta_{d-1}\},F(x_{d}\,|\,\bm{x}_{1:{d-1}});\Theta_{d}]$
+$\ldots$
+$=h[\ldots h\{h(F(y),F(x_{1});\Theta_{1}),F(x_{2}\,|\,x_{1});\Theta_{2}\},\ldots,F(x_{d}\,|\,\bm{x}_{1:{d-1}});\Theta_{d}].$ (4)
+Intuitively, the bivariate copulas “weave” the dependence between the outcome and each of the covariates.
+Thus, given a value for $\bm{X}$ we can simulate $Y$ by first drawing a uniform $U\sim{\rm U}(0,1)$, and then evaluating:
+$F(Y)=h^{-1}[\ldots h^{-1}\{h^{-1}(U,\,F(X_{1});\Theta_{1}),F(X_{2}\,|\,X_{1});\Theta_{2}\},\ldots,F(X_{d}\,|\,X_{1},\ldots,X_{d-1});\Theta_{d}].$ (5)
+This means that to simulate $Y$ given a $d$-dimensional vector $\bm{X}$, we recursively calculate the $d$ inverse $h$-functions, forming a pair-copula construction.
+This underpins our proposed simulation method.
 4 Simulation Algorithm
 […]
-Figure 1: An example directed acyclic graph showing temporal relationships in a longitudinal discrete setting, with baseline covariates $Z$; treatments $A_{0},A_{1},\ldots,A_{K}$; time-varying covariates $L_{1},\ldots,L_{K}$ and outcomes $Y$ observed at the end of the K follow up intervals.
+Figure 1: An example directed acyclic graph showing temporal relationships in a longitudinal discrete setting, with baseline covariates $Z$; treatments $A_{0},\ldots,A_{K}$; time-varying covariates $L_{0},\ldots,L_{K}$ and an outcome $Y$ observed at the end of the K follow up intervals.
 Given the topological ordering $Z,L_{0},A_{0},L_{1},A_{1}\ldots,L_{K},A_{K},Y$, any variable may depend upon any subset of the previous variables.
-We consider a longitudinal study with $n$ individuals observed over discrete time points, $\mathcal{K}=\{0,1,\dots,K\}$.
+We first consider a longitudinal study with $n$ individuals observed at discrete time points, $\mathcal{K}=\{0,\ldots,K\}$.
 At each follow-up $k$, treatment status $A_{k}$, and a set of time-varying covariates $L_{k}$ are recorded.
-Additionally, $Z$ denotes a set of static covariates measured at baseline.
-There may also be latent variables, represented by $\{U_{k}\}_{k\in\mathcal{K}}$.
-To indicate the history of a variable up to time $k$, we use a bar notation.
-For instance, $\bar{A}_{k}=\{A_{0},A_{1}\dots,A_{k}\}$ and $\bar{L}_{k}=\{L_{0},L_{1},\dots,L_{k}\}$.
-For convenience, we let $\bar{A}_{-1}=\emptyset$ and $\bar{L}_{-1}=\emptyset$.
+Additionally, $Z$ represents a set of static baseline covariatesatent variables, if present, are written as $\{H_{k}\}_{k\in\mathcal{K}}$.
+To indicate the history of a variable up to time $k$, we use bar notation.
+For instance, $\bar{A}_{k}=\{A_{0},\ldots,A_{k}\}$ and $\bar{L}_{k}=\{L_{0},\ldots,L_{k}\}$, with the convention $\bar{A}_{-1}=\emptyset$ and $\bar{L}_{-1}=\emptyset$.
 The outcome $Y$ is observed at the final time point $K$.
-Figure 1 shows a DAG representing this setup, which is common in medical and epidemiological research, such as examining the effects of iron supplement use during pregnancy on the odds of anemia at delivery (Bodnar et al., 2004), anti-diabetic drugs on weight loss over a specified period (e.g.
-Jastreboff et al., 2022; Ryan et al., 2024) and exposure to community violence on youth’s mental health (Kennedy et al., 2023).
-In Section 5, we will illustrate our method with a simulation study that replicates such setups.
+Figure 1 is a graphical representation of this setup, which is common in medical and epidemiological research.
+Examples include examining the effects of an HIV treatment on infection status assessed at the end of the study period (Robins et al., 2000), and assessing the impact of iron supplement use during pregnancy on the odds of anemia at delivery (Bodnar et al., 2004).
+In Section 4.4, we illustrate our method through a simulation study that replicates such settings.
 Following the notation of Evans and Didelez (2024), we use an asterisk to represent causal or interventional distributions where treatment is set to a specific value $A_{k}=a$ via intervention.
 In graphical terms, this corresponds to removing all arrows directed into $A_{k}$.
-4.1 Fully conditional approach
-For simplicity, we initially assume the time-varying covariate $L_{k}$ is one-dimensional and exclude the latent process $\{U_{k}\}_{k\in\mathcal{K}}$; the extension to multi-dimensional time-varying covariates and latent variables is discussed later in this section.
-The usual strategy for simulation of longitudinal samples $(Z,\bar{L}_{K},\bar{A}_{K},Y)$ is by time step in blocks using the Rosenblatt transformation (Rosenblatt, 1952). Specifically:
+4.1 The fully-conditional approach
+To outline our approach, we initially assume the time-varying covariate $L_{k}$ is one-dimensional and omit any latent process $\{H_{k}\}_{k\in\mathcal{K}}$; the extension to multi-dimensional time-varying covariates and latent variables is discussed in Appendix A.
+The standard method for simulation of longitudinal data $(Z,\bar{L}_{K},\bar{A}_{K},Y)$ is by time step in blocks using the Rosenblatt transformation (Rosenblatt, 1952). Specifically:
   1. 1.
-    Baseline simulation: Simulate static baseline variables $z=F_{Z}^{-1}(w_{z})$, where $w_{z}$ is a uniform random variable.
+    Baseline simulation: Simulate static baseline variables, $z$ following a specified factorization, e.g.  $F_{Z_{1}},F_{Z_{2}\mid Z_{1}},\ldots$.
   2. 2.
-    Time-step simulation: For each time point $k=0,1,\dots,K$, simulate:
-    $l_{k}$ $=F_{L_{k}}^{-1}(w_{k,1},\,|\,z,\bar{l}_{k-1},\bar{a}_{k-1})$ (10)
-    $a_{k}$ $=F_{A_{k}}^{-1}(w_{k,2},\,|\,z,\bar{l}_{k},\bar{a}_{k-1}),$
-    where $w_{k,1},w_{k,2}$ are random uniform variables.
+    Time-step simulation: At each time point $k=0,\dots,K$, simulate $l_{k}=F_{L_{k}\mid Z,\bar{L}_{k-1},\bar{A}_{k-1}}^{-1}(w_{k,1},\,|\,z,\bar{l}_{k-1},\bar{a}_{k-1})$ and $a_{k}=F_{A_{k}\mid Z,\bar{L}_{k},\bar{A}_{k-1}}^{-1}(w_{k,2},\,|\,z,\bar{l}_{k},\bar{a}_{k-1})$, where $w_{k,1},w_{k,2}\sim\operatorname{U}(0,1)$.
   3. 3.
-    Outcome simulation: After simulating all time-varying covariates and treatment, simulate the outcome $Y$ as
-    $y$ $=F_{Y}^{-1}(w_{y}\,|\,z,\bar{l}_{K},\bar{a}_{K}),$ (11)
-    with $w_{y}$ uniformly distributed.
+    Outcome simulation: Given the simulated covariates and treatments, generate the outcome $y=F_{Y\mid Z,\bar{L}_{K},\bar{A}_{K}}^{-1}(w_{y}\,|\,z,\bar{l}_{K},\bar{a}_{K})$, with $w_{Y}\sim\operatorname{U}(0,1)$.
+Our proposed method diverges from this fully-conditional approach at Step 3.
+Instead of specifying the conditional distribution $F_{Y\mid Z\bar{L}_{K}\bar{A}_{K}}$—which inherently introduces misalignment between conditional and the desired marginal models—we introduce a margin-preserving parameterization for outcome simulation.
 4.2 Frugal parameterization for outcome simulation
-Our method diverges from traditional conditional approaches by using the frugal parameterization for $Y$ to avoid intractable analytical calculations for the marginal structural model.
-Instead of directly assuming a conditional distribution $F_{Y}^{-1}(w_{y}\,|\,z,\bar{l}_{K},\bar{a}_{K})$, we apply interventional distributions to decompose this into interventional margins and dependency structures.
-Evans and Didelez (2024) showed that the conditional distribution $F_{Y}^{-1}(w_{y}\,|\,z,\bar{l}_{k},\bar{a}_{K})$ can be frugally parameterized as:
-$p(y\,|\,z,\bar{l}_{k},\bar{a}_{K})=p^{*}_{Y\mid Z\bar{A}_{K}}(y\,|\,z,do(\bar{a}_{K}))\cdot c^{*}_{Y\bar{L}_{K}\mid Z\bar{A}_{K}}(u^{*}_{y},u^{*}_{l_{0}},\dots,u^{*}_{l_{K}}\,|\,z,do(\bar{a}_{K})),$ (12)
-where $u^{*}_{y}=F(y\,|\,z,do(\bar{a}_{K}))$ and $u^{*}_{l_{k}}=F(l_{k}\,|\,z,do(\bar{a}_{K}))$ for $k=0,1,\ldots,K$ are the marginal quantiles under interventions that set $\bar{A}_{K}=\bar{a}_{K}$.
-We mark these quantiles with an asterisk to emphasize their interventional nature.
-This decomposition implies that we can simulate $Y$ through the interventional copula and the interventional margin.
-Applying the pair-copula-construction from (4) to the expression in (12), we obtain:
-$p(y\,|\,z,\bar{l}_{K},\bar{a}_{K})$ $=c^{*}_{YL_{K}\mid\bar{L}_{K-1},Z,\bar{A}_{K}}\{F(y\,|\,\bar{l}_{K-1},z,do(\bar{a}_{K})),F(l_{K}\,|\,\bar{l}_{K-1},z,do(\bar{a}_{K}));\Theta_{K}\}$
-$\quad\cdot c^{*}_{YL_{K-1}\mid\bar{L}_{K-2},Z,\bar{A}_{K}}\{F(y\,|\,\bar{l}_{K-2},z,do(\bar{a}_{K})),F(l_{K-1}\,|\,\bar{l}_{K-2},z,do(\bar{a}_{K}));\Theta_{K-1}\}$
-[…]
-$\quad\cdot c^{*}_{YL_{0}\mid Z,\bar{A}_{K}}\{F(y\,|\,z,do(\bar{a}_{K})),F(l_{0}\,|\,z,do(\bar{a}_{K}));\Theta_{0}\}\cdot p^{*}_{Y\mid Z\bar{A}_{K}}(y\,|\,z,do(\bar{a}_{K}))$ (13)
-$=\prod_{k=0}^{K}c^{*}_{Y,L_{k}\mid\bar{L}_{k-1},Z,\bar{A}_{K}}\{F(y\,|\,\bar{l}_{k-1},z,do(\bar{a}_{K})),F(l_{k}\,|\,\bar{l}_{k-1},z,do(\bar{a}_{K}));\Theta_{k}\}.$ (14)
+Our method uses the frugal parameterization to simulate outcomes, avoiding potentially intractable analytical calculations to obtain the marginal structural model.
+Evans and Didelez (2024) showed that the fully-conditional distribution of the outcome can be frugally parameterized as:
+$p_{Y\mid\bar{L}_{K},\bar{A}_{K}}(y\,|\,\bar{l}_{k},\bar{a}_{K})=p^{*}_{Y\mid\bar{A}_{K}}(y\,|\,do(\bar{a}_{K}))\cdot c^{*}_{Y\bar{L}_{K}\mid\bar{A}_{K}}(u^{*}_{Y},u_{L_{0}},\dots,u_{L_{K}}\,|\,do(\bar{a}_{K})),$ (6)
+where $u^{*}_{Y}=F_{Y\mid\bar{A}_{K}}(y\,|\,do(\bar{a}_{K}))$ and $u_{L_{k}}=F_{L_{k}\mid\bar{A}_{K}}(l_{k}\,|\,do(\bar{a}_{K}))$ for $k=0,\ldots,K$ are the marginal quantiles under interventions that set $\bar{A}_{K}=\bar{a}_{K}$.
+Remark 4.1.
+For brevity, we omit conditioning on baseline variables $Z$ in all quantities in (6), treating them as if no baseline variables exist.
+This omission will continue hereafter, though all expressions can be extended to condition on $Z$ by including it in their respective conditioning sets.
+This decomposition implies that we can simulate $Y$ through the multivariate copula and its interventional margin. Applying the pair-copula construction in (2) to (6), we obtain:
+$p(y\,|\,\bar{l}_{K},\bar{a}_{K})$ $=p(y\,|\,do(\bar{a}_{K}))\cdot\prod_{k=0}^{K}c^{*}\{F(y\,|\,\bar{l}_{k-1},do(\bar{a}_{K})),F(l_{k}\,|\,\bar{l}_{k-1},do(\bar{a}_{K}));\Theta_{k}\}.$ (7)
+Note that we omit some subscripts on densities for conciseness where doing so does not introduce ambiguity.
 This decomposes the conditional density into the product of $K+1$ pair-copulas and the target marginal causal density.
+The pair-copula linking $Y$ and $L_{k}$ is parameterized by $\Theta_{k}$, which may depend on intervened treatments $do(\bar{A}_{k})$, past time-varying confounders $\bar{L}_{k-1}$ and baseline covariates $\bm{Z}$.
 Under intervention, graphically speaking, all arrows into nodes $\bar{A}_{k}$ are removed in the directed acyclic graph, rendering $L_{k}$ independent of subsequent interventions.
-Consequently, $F(l_{k}\,|\,\bar{l}_{k-1},z,do(\bar{a}_{K}))=F(l_{k}\,|\,\bar{l}_{k-1},z,\bar{a}_{k-1})$ for $k=1,2,\ldots,K$ and $F(l_{0}\,|\,z,do(\bar{a}_{K}))=F(l_{0}\,|\,z)$.
-Given this, following the simulation strategy from pair-copula constructions as described in Section 3.2.2, the conditional cumulative distribution function in (11) becomes
-$F(y\,|\,z,\bar{l}_{K},\bar{a}_{K})$ $=h[\dots h\{h(F^{*}(y\,|\,z,\bar{a}_{K}),F(l_{0}\,|\,z);\Theta_{0}),F(l_{1}\,|\,l_{0},z,a_{0});\Theta_{1}\}\dots;\Theta_{K}].$
+Consequently, $F(l_{k}\,|\,\bar{l}_{k-1},do(\bar{a}_{K}))=F(l_{k}\,|\,\bar{l}_{k-1},\bar{a}_{k-1})$ for $k=1,\ldots,K$ and $F(l_{0}\,|\,do(\bar{a}_{K}))=F(l_{0})$.
+Following the simulation strategy from pair-copula constructions outlined in Section 3.2.2, the fully-conditional distribution of the outcome is alternatively parameterized as:
+$F(y\,|\,\bar{l}_{K},\bar{a}_{K})$ $=h\big{[}\dots h\{h(F^{*}(y\,|\,do(\bar{a}_{K})),F(l_{0});\Theta_{0}),F(l_{1}\,|\,l_{0},a_{0});\Theta_{1}\}\dots;\Theta_{K}\big{]},$
+which is centered around the causal margin $F^{*}(y\,|\,\bar{a}_{K})$.
 Through inversion the outcome can be simulated in two steps:
   1. 1.
-    Calculate $u^{*}_{y}$ as
-    $u^{*}_{y}$ $=h^{-1}\big{[}\dots h^{-1}\big{\{}h^{-1}(w_{y},F(l_{K}\,|\,\bar{l}_{K-1},z,\bar{a}_{K});\Theta_{K}),$
-    $\qquad\qquad\qquad F(l_{K-1}\,|\,\bar{l}_{K-2},z,\bar{a}_{K-1});\Theta_{K-1}\big{\}}\dots,F(l_{0}\,|\,z);\Theta_{0}\big{]}$ (15)
-    $=h^{-1}\big{[}\dots h^{-1}\big{\{}h^{-1}(w_{y},w_{K,1};\Theta_{K}),w_{K-1,1};\Theta_{K-1}\big{\}},\dots,w_{0,1};\Theta_{0}\big{]}.$ (16)
+    Calculate the marginal quantile of the outcome, $u^{*}_{Y}$, as:
+    $u^{*}_{Y}$ $=h^{-1}\big{[}\dots h^{-1}\big{\{}h^{-1}(w_{Y},F(l_{K}\,|\,\bar{l}_{K-1},\bar{a}_{K-1});\Theta_{K}),$
+    $\qquad\qquad\qquad F(l_{K-1}\,|\,\bar{l}_{K-2},\bar{a}_{K-2});\Theta_{K-1}\big{\}}\dots,F(l_{0});\Theta_{0}\big{]}$ (8)
+    $=h^{-1}\big{[}\dots h^{-1}\big{\{}h^{-1}(w_{y},w_{K,1};\Theta_{K}),w_{K-1,1};\Theta_{K-1}\big{\}},\ldots,w_{0,1};\Theta_{0}\big{]},$
+    where $w_{Y}\sim\operatorname{U}(0,1)$.
   2. 2.
-    Obtain $y$ as:
-    $y=F^{*-1}(u^{*}_{y}\,|\,z,do(\bar{a}_{K})).$
-The quantiles in the inverse $h$ functions, such as $F(l_{k}\,|\,\bar{l}_{k-1},\bar{a}_{k},z)$, are precisely the random variables $w_{k,1}$ used in simulate them in (2).
-This means that we can directly substitute the $l_{k}$ quantiles into (15) without the need for computation.
+    Obtain $y$ through the inverse cumulative distribution function of the marginal distribution: $y=F^{*-1}_{Y\mid\bar{A}_{K}}(u^{*}_{Y}\,|\,do(\bar{a}_{K}))$.
+An intuitive way to understand the calculation in Step 1 is that the inverse $h$-functions “disentangle” the dependencies between $Y$ and each of $L_{K},L_{K-1},\dots L_{0}$ sequentially, and remove the conditioning of $Y$ on $\bar{L}_{K}$ until the marginal quantile of $Y$ is obtained.
+The quantiles in the inverse $h$-functions, such as $F(l_{k}\,|\,\bar{l}_{k-1},\bar{a}_{k-1})$, correspond exactly to the random standard uniform variables $w_{k,1}$ used to simulate $l_{k}$ at time step $k$.
+This enables direct substitution of the $l_{k}$ quantiles into (8) without additional computation.
 4.3 Simulation Algorithm
-To simulate data $(Z,\bar{L}_{K},\bar{A}_{K},Y)$ in this example, the user specifies:
-  • the population distribution of the static baseline covariate $Z$: $F_{Z}$;
-  • the development of the time-varying confounder $L_{k}$ as influenced by $Z$ and treatment history $\bar{A}_{k-1}$: $F_{L_{k}\mid Z,\bar{L}_{k-1},\bar{A}_{k-1}}$;
-  • treatment assignment based on the time-varying confounder $L_{k}$, past treatment and baseline covariate $Z$: $F_{A_{k}\mid Z,\bar{L}_{k},\bar{A}_{k-1}}$;
-  • the target causal marginal structural model: $F_{Y_{k+1}\mid\bar{A}_{k},Z}^{*}(\cdot\,|\,do(\bar{a}_{K}),z)$;
-  • $K+1$ pair-copulas describing dependencies between outcome $Y$ and $L_{0},L_{1},\ldots,L_{K}$:
-    $C_{YL_{k}\mid Z,do(\bar{A}_{k}),\bar{L}_{k-1}}\quad\text{for}\quad k=0,1,\ldots,K.$
-Section 5 provides examples of specifying and interpreting these inputs through a simulation study that capture real-world complexities.
+In summary, to simulate the data $(Z,\bar{L}_{K},\bar{A}_{K},Y)$, the user specifies:
+  • the population distribution of the static baseline covariates $Z$: $F_{Z}$, either as a joint distribution, or through a conditional factorization;
+  • the development of the time-varying confounder $L_{k}$ as influenced by $\bm{Z}$ and $\bar{A}_{k-1}$: $F_{L_{k}\mid Z,\bar{L}_{k-1},\bar{A}_{k-1}}$;
+  • treatment assignment based on the time-varying confounder $\bar{L}_{k}$, past treatment and baseline covariates $Z$: $F_{A_{k}\mid\bm{Z},\bar{L}_{k},\bar{A}_{k-1}}$;
+  • the target causal marginal structural model: $F_{Y\mid\bar{A}_{k},Z}^{*}(\cdot\,|\,do(\bar{a}_{K}),z)$;
+  • $K+1$ bivariate copulas describing dependencies between outcome $Y$ and $L_{0},\ldots,L_{K}$:
+    $C_{YL_{k}\mid Z,do(\bar{A}_{k}),\bar{L}_{k-1}}(\cdot;\Theta_{k})\quad\text{for}\quad k=0,\ldots,K.$
+Section 4.4 provides an example of specifying and interpreting these inputs through a simulation study that captures real-world complexities.
 Algorithm 1 below details the simulation procedure:
@@
 while $k\leq K$ do
-     Sample $w_{k,1},w_{k,2}$ independent uniform on $[0,1]$.
+     Sample $w_{k,1},w_{k,2}$ independent standard uniform variables
      $l_{k}\leftarrow F^{-1}_{L_{k}\mid Z,\bar{L}_{k-1},\bar{A}_{k-1}}(w_{k,1}\,|\,z,\bar{l}_{k-1},\bar{a}_{k-1})$
@@
 $u^{*}_{y}\leftarrow\nu_{K+1}$
-$y\leftarrow F_{Y_{k+1}\mid\bar{A}_{k},Z}^{*-1}(u^{*}_{y}\,|\,do(\bar{a}_{K}),z)$
+$y\leftarrow F_{Y\mid\bar{A}_{k},Z}^{*-1}(u^{*}_{y}\,|\,do(\bar{a}_{K}),z)$
 return $(z,\bar{l}_{K},\bar{a}_{K},y)$
-4.4 Extension to Include Time-Varying Latent Variables
-We now extend the simulation algorithm discussed above to allow for a latent process that confounds the covariates and the final outcome $Y$.
-As an example, the steps to simulate longitudinal data from the model depicted in Figure 2 are:
-[…]
-Figure 2: Extension to the model in Figure 1 in which we also allow an unobserved process $U_{t}$ affecting only $Z$, $L_{k}$ and $Y$.
-  1. 1.
-    Baseline simulation: The DAG in Figure 1 imposes that the latent variable $U_{0}$ precedes $Z$ in topological order, which may indicate an initial selection bias.
-Thus, we first simulate the baseline value of $U$ as $u_{0}=F_{U_{0}}^{-1}(w_{0,0})$, followed by the static baseline covariate $Z$ given $U_{0}$: $z=F_{Z\mid U_{0}}^{-1}(w_{z}\,|\,u_{0})$, where $w_{0,0}$ and $w_{z}$ are uniform random numbers.
-    $u_{0}$ $=F_{U_{0}}^{-1}(w_{0,0})$
-    $z$ $=F_{Z\mid U_{0}}^{-1}(w_{z}\,|\,u_{0})$
-    $l_{0}$ $=F_{L_{0}\mid ZU_{0}}^{-1}(w_{0,1}\,|\,z,u_{0})$
-    $a_{0}$ $=F_{A_{0}\mid ZU_{0}L_{0}}^{-1}(w_{0,2}\,|\,z,u_{0},l_{0})$
-  2. 2.
-    Time-step simulation: For each time point $k=1,\dots,K$, simulate:
-    $u_{k}$ $=F_{U_{k}\mid\bar{U}_{k-1}}^{-1}(w_{k,1}\,|\,\bar{u}_{k-1})$
-    $l_{k}$ $=F_{L_{k}\mid Z\bar{U}_{k}\bar{L}_{k-1}\bar{A}_{k-1}}^{-1}(w_{k,2},\,|\,z,\bar{u}_{k},\bar{l}_{k-1},\bar{a}_{k-1})$
-    $a_{k}$ $=F_{A_{k}\mid Z\bar{U}_{k}\bar{L}_{k}\bar{A}_{k-1}}^{-1}(w_{k,3},\,|\,z,\bar{u}_{k},\bar{l}_{k},\bar{a}_{k-1})$
-    where $w_{k,1},w_{k,2},w_{k,3}$ are random uniform variables.
-  3. 3.
-    Outcome simulation: After simulating all time-varying covariates and treatment, and applying the same derivation following the frugal parameterization as in Section 4.2, simulate the outcome $Y$ in two steps:
-    1. (a)
-      Calculate
-      $u^{*}_{y}$ $=h^{-1}\big{[}\dots h^{-1}[h^{-1}\left\{h^{-1}\left(w_{y},F(l_{K}\,|\,\bar{l}_{K-1},\bar{u}_{K},z,\bar{a}_{K});\Theta_{K,1}\right),F(u_{K}\,|\,\bar{l}_{K-1},\bar{u}_{K-1},z,\bar{a}_{K});\Theta_{K,2}\right\}$
-      $F(l_{K-1}\,|\,\bar{l}_{K-2},\bar{u}_{K-1},z,\bar{a}_{K-1});\Theta_{K-1,1}]\dots;\Theta_{0,2}\big{]}.$
-    2. (b)
-      Obtain: $y=F^{*^{-1}}(u^{*}_{y}\,|\,z,do(\bar{a}_{K}))$.
-    Here the $w_{y}$ are independent and uniformly distributed on $(0,1)$.
-In addition to the information required in the case without latent confounders, the user also specifies:
-  • the development of the latent confounder $U_{k}$ as influenced by $Z$ and treatment history $\bar{A}_{k}$;
-  • $K+1$ pair-copula densities describing dependencies between outcome $Y$ and each of $U_{0},U_{1},\ldots,U_{K}$:
-    $c_{YU_{k}\mid Z,\bar{U}_{k-1},\bar{L}_{k-1},do(\bar{A}_{k-1})},$
-    for $k=0,1,\ldots,K$.
-Extending to include additional time-varying covariates follows a similar approach.
-We specify the evolution of each time-varying covariate and use pair copulas to parameterize their association with the outcome, conditioned on all preceding variables.
-With common choices of parametric bivariate copulas, evaluating the $h$-functions is fast, making the simulation efficient.
-Increasing the dimensions of time-varying confounders, and the number of time steps does not impose significant computational challenges.
-5 Demonstration
-We first demonstrate our simulation method with an example replicating the setting in Bodnar et al.
+Extending the simulation algorithm to include a latent process $\{H_{k}\}_{k\in\mathcal{K}}$, or multiple time-varying confounders is straightforward.
+The key modification lies in the outcome simulation: the marginal quantile of $Y$ is obtained by sequentially removing the conditioning on all the time-dependent confounders at each step, through inverse $h$-functions.
+We provide an illustrative example including a latent variable $H_{k}$ in Appendix A.
+4.4 Illustrative example
+We demonstrate our simulation method by replicating the setting in Bodnar et al.
 (2004), which examines the causal effect of iron supplement use during pregnancy on the odds of anemia at delivery.
-Participants are initially randomized into treatment and control groups and are followed up in four visits leading up to delivery.
-Like the Iron Supplementation Study analyzed in Bodnar et al.
-(2004), we assume that the participants adhere to their randomly assigned treatment up to $T=3$, after which the standard clinical protocol applies.
-[…]
-Baseline:
+Participants are randomized into treatment and control groups and followed up over four visits leading up to delivery.
+Treatment is not time-fixed as participants are assumed to adhere to assigned treatment up to $k=2$, after which standard clinical protocols dictate further treatment based on hemoglobin levels.
 […]
 […]
-For subsequent follow-up visits $k>0$:
-[…]
-$A_{2}$ $=A_{1}=A_{0}$ $A_{k}\sim\operatorname{Bernoulli}(\operatorname{expit}(10-L_{k}+0.1\,A_{k-1}))\text{ for }k=3,4.$
-We choose a logistic model for a discrete-time hazard marginal structural model:
 $\mathbb{P}[Y=1\,|\,do(\bar{A}_{4}),B,C]=\operatorname{expit}(-2+0.1\,B+0.02\,C-\beta\,\operatorname{cum}(\bar{A}_{4})),$
 where $\operatorname{cum}(\bar{A}_{4})=\Sigma_{i=1}^{4}A_{i}$.
-This assumes $\exp(\beta)$ as the causal odds ratio of anemia for an incremental dose.
-If all women receive iron supplements at all four visits, the causal odds ratio would be $\exp(4\beta)$, representing the true value for the “always treated” efficacy estimand.
+Here, $\exp(\beta)$ represents the causal odds ratio for an incremental dose, and $\exp(4\beta)$ corresponds to the “always treated” efficacy estimand.
 We use a t-copula $T_{\rho,\operatorname{df}}$ (with $\rho=-0.5$ and $\operatorname{df}=5$) across all pair-copula construction components to model the dependency between $Y$ and each of $L_{0},L_{1},L_{2},L_{3}$, and $L_{4}$.
-Compared to a Gaussian copula, the t-copula allows for heavier-tailed dependency, capturing stronger associations in the tails.
-In our example, t-copulas reflect the likelihood that extreme Hb levels (e.g., severely low) are more strongly associated with extreme outcomes (anemia at delivery).
-This dependency structure may better reflect the underlying association than a Gaussian copula.
+Compared to a Gaussian copula, the t-copula captures heavier-tailed dependency, capturing stronger associations in the tails reflecting stronger associations in extreme scenarios (e.g., severely low hemoglobin levels and anemia).
 Our algorithm offers full flexibility, allowing users to choose any copula family and specify different parameters for each bivariate copula in the pair-copula construction.
-We simulated scenarios with varying causal coefficients $\beta=\{-0.5,0.3,0\}$, representing strong, moderate and no effect, respectively and considered different samples size $n=\{500,1000,2000\}$.
-For each combination of $\beta$ and $n$ we generated 400 datasets.
-We fitted the model $Y=\operatorname{expit}(\alpha+\gamma\,B+\theta\,C-\beta\,\operatorname{cum}(\bar{A}_{4}))$ using two approaches: with stablized inverse probability of treatment weights, and an unweighted approach.
-For each fit, we computed 95% confidence intervals for $\beta$ using two methods: (1) the robust sandwich estimator of the variance, with rationale provided by Hernán et al.
-(2000) that the use of weights induces within-subject correlation, and (2) non-parametric bootstrap, constructing 95% confidence intervals from the 2.5% and 97.5% percentiles from 200 bootstrap samples.
-Table 1 presents the bias with corresponding Monte Carlo standard errors.
-The inverse-probability-weighted estimator shows minimal bias, with Monte Carlo standard errors decreasing as sample size increases.In contrast, the unweighted estimator remains significantly biased across sample sizes, showing no convergence.
-Table 2 reports coverage rates.
-Both the sandwich and bootstrap CIs approach 95% coverage, though with minor under-coverage.
-The unweighted estimator, however, exhibits poor coverage due to significant bias in $\beta$.
-Table 2 shows the statistical power of the test of treatment effect (null hypothesis: $\beta=0$) for each estimator, with power of the weighted estimator increasing as the causal effect size and sample size grow.
-In summary, these results confirm that data generated using our simulation algorithm accurately reflect the intended causal effects.
+We simulated scenarios with varying causal coefficients $\beta=\{-0.5,-0.3,0\}$, representing strong, moderate and no effect, respectively and samples sizes $n=\{500,1000,2000\}$.
+For each combination, we generated 400 datasets.
+The model $Y=\operatorname{expit}(\alpha+\gamma\,B+\theta\,C-\beta\,\operatorname{cum}(\bar{A}_{4}))$ is fitted using stabilized inverse probability of treatment weights, and an unweighted approach.
+For each fit, we estimated 95% confidence intervals for $\beta$ using both the robust sandwich estimator and non-parametric bootstrap.
+Results are summarized in Table 1, 2 and 3.
+The inverse-probability-weighted estimator shows minimal bias, with magnitude well below Monte Carlo standard errors.
+In contrast, the unweighted estimator remains significantly biased across sample sizes, showing no convergence.
+Both the sandwich and bootstrap CIs of the inverse-probability-weighted estimator approach 95% coverage, though with minor under-coverage.
+The unweighted estimator, however, exhibits poor coverage due to bias.
+Statistical power for testing the treatment effect (null hypothesis: $\beta=0$) increases with effect size and sample size for the weighted estimator.
+In summary, these results confirm that data generated using our simulation algorithm accurately reflect the intended causal effects and time-varying confounding is introduced.
 $\beta$ $n$ IPTW Unweighted
@@
 2000 0.014 (0.072) 0.193 (0.043)
-Table 1: Bias and Monte Carlo standard errors (in parentheses, defined as $sd(\hat{\beta})$) for inverse-probability-of-treatment-weighted (‘IPTW’) and unweighted estimators (‘unweighted) across different $\beta$ and $n$ values.
+Table 1: Bias and Monte Carlo standard errors (in parentheses, defined as $sd(\hat{\beta})$) for inverse-probability-of-treatment-weighted (‘IPTW’) and unweighted estimators (‘unweighted’) across different $\beta$ and $n$ scenarios.
 $\beta$ $n$ IPTW Unweighted
@@
 2000 92.8% 90.0% 0.5%
-Table 2: Coverage of 95% confidence intervals for IPTW-sandwich, IPTW-bootsrap, and unweighted estimators across scenarios with varying $\beta$ and $n$ values.
+Table 2: Coverage of 95% confidence intervals for IPTW-sandwich, IPTW-bootstrap, and unweighted estimators across scenarios with varying $\beta$ and $n$ values.
 $\beta$ $n$ IPTW Unweighted
@@
 2000 7.3% 10.0% 99.5%
-Table 3: Statistical power of the test for the null hypothesis $\beta=0$, i.e., no causal effect.
+Table 3: Statistical power to reject the null hypothesis $\beta=0$; that is, no causal effect.
 5 Simulating Survival Outcomes
@@
 Figure 2: Example directed acyclic graph showing temporal relationships in a longitudinal survival setting.
-The sequence of outcomes observed over time is denoted as $\{Y_{k}\}_{k\in\mathcal{K}^{+}}$, where $\mathcal{K}^{+}=\{1,\dots,K+1\}$.
+The sequence of outcomes observed over time is denoted as $\{Y_{k}\}_{k\in\mathcal{K}^{+}}$, where $\mathcal{K}^{+}=\{1,\ldots,K+1\}$.
 We have described and demonstrated our method to simulate longitudinal data where the outcome is measured at the final time point.
 Since time-to-event and survival outcomes are very common in medical and epidemiological contexts, we shall now demonstrate how our method can be extended to simulate longitudinal data from survival models.
-As shown in Figure 3, we consider the setup described in Section 4 to remain the same except that the outcome now becomes a sequence of outcomes observed over time as $\{Y_{k}\}_{k\in\mathcal{K}^{+}}$, where $\mathcal{K}^{+}=\{1,\dots,K+1\}$ and $Y_{k}=1$ represents failure.
-An example study that follows this setup would be one evaluating the effect of an anti-diabetic drug on cardiovascular risk.
-The primary endpoint is time to the first occurrence of a major adverse cardiovascular event (MACE) [^2] A composite endpoint of death from cardiovascular causes, non-fatal myocardial infarction, or non-fatal stroke..
-The baseline variables include sex and medical history, and an example time-varying confounder might be HbA1c; the treatment influences HbA1c which in turn predicts intercurrent events—for example, use of additional medication—and MACE.
-We will refer to this example as we explain the algorithm and assumptions in this section.
-Similarly, we describe the algorithm for the case where the time-varying confounder is univariate in the absence of any latent procedures.
-It is fairly straightforward to extend to multiple confounders and latent variables in a manner analogous to ordinary longitudinal outcomes (see Section 4.4).
-Depending on the chosen marginal structural model, there are various ways to model the outcomes $Y_{k}$’s. Examples include:
-  • Discrete-time Cox marginal structural model, where we model the binary survival outcome $Y_{k}\in\{0,1\}$ directly:
-    $\frac{P(Y_{k+1}=1\,|\,do(\bar{a}_{k}),Z,Y_{k}=0)}{P(Y_{k+1}=1\,|\,do(\bar{0}_{k}),Z,Y_{k}=0)}=f(\bar{a}_{k},Z;\alpha).$ (17)
-  • Continuous-time hazard MSM, where we model survival time $T$ and observe the binary survival outcome $Y_{k}=I(T>k)$. Specifically
-    • Cox proportional hazard: $\lambda(t\,|\,do(\bar{a}_{k}),Z)=\lambda(t\,|\,do(\bar{0}_{k}),Z=z_{0})\exp\{g(\bar{a}_{t},Z;\beta)\}$,
-    • Additive hazard: $\lambda(t\,|\,do(\bar{a}_{k}),Z)=\lambda(t\,|\,do(\bar{0}_{k}),Z=z_{0})+g(\bar{a}_{t},Z;\gamma)$.
-In Section 6.2 we illustrate the simulation of time-to-event outcomes from a Cox marginal structural model.
+In Figure 2, each plate represents a single interval.
+For the $k$th interval, where $k=0,\ldots,K$, the observations include $L_{k},A_{k}$ at the start of the interval, and the survival status $Y_{k+1}$ at the end.
+The length of each interval is set to 1.
+In addition to the notation in Section 4, let $T$ denote an individual’s failure time.
+The outcome is represented be a sequence $\{Y_{k}\}_{k\in\mathcal{K}^{+}}$, where $\mathcal{K}^{+}=\{1,\ldots,K+1\}$ and $Y_{k+1}=I(T<k+1)$ indicates failure.
+We describe the algorithm for the case where the time-varying confounder is univariate and no latent processes are present.
+The algorithm can be extended analogously to the case of ordinary longitudinal outcomes.
+The illustrative example at the end of this section demonstrates the simulation with two time-varying confounders.
+Depending on the chosen marginal structural model, there are various ways to model the outcomes. Examples include:
+  • Discrete-time marginal structural model, where the binary survival outcome $Y_{k}$ is modeled directly:
+    $P(Y_{k+1}=1\,|\,do(\bar{A}_{k})=\bar{a}_{k},Z=z,Y_{k}=0)=f(\bar{a}_{k},z;\alpha),$ (9)
+    where $f$ can be any real-valued functions bounded between 0 and 1, for example, an expit function for a logistic marginal structural model.
+  • Continuous-time hazard MSM, where we introduce an auxiliary continuous variable, $\tilde{Y}_{k+1}=T-k\,|\,T>k$, representing the incremental failure time given survival at the start of the $k$th interval.
+We model incremental survival time $\tilde{Y}_{k+1}$ and truncated it to obtain the observed binary survival outcome $Y_{k+1}$.
+Given survival at $k$, the time to failure is $T\,|\,T>k=k+\tilde{Y}_{k+1}$.
+Examples of marginal models for $\tilde{Y}_{k+1}$ include:
+    • Cox proportional hazard: $\lambda(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),z,y_{k}=0)=\lambda(\tilde{y}_{k+1}\,|\,do(\bar{0}_{k}),z,y_{k}=0)\exp\{g(\bar{a}_{k},z;\beta)\}$,
+    • Additive hazard: $\lambda(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),z,y_{k}=0)=\lambda(\tilde{y}_{k+1}\,|\,do(\bar{0}_{k}),z,y_{k}=0)+g(\bar{a}_{k},z;\gamma)$.
+We focus on describing the simulation algorithm for time-to-event outcomes, with a remark on adapting the approach for discrete survival outcomes, which follows a similar but simpler framework.
 5.1 Simulation algorithm
-The general idea is very similar to the non-survival case and the difference is that in the time-step simulation, the simulation of step $k$ is conditional on survival up to the previous time step.
-Unlike the derivation of outcome simulation under the frugal parameterization for the non-survival case discussed in Section 4.2, we now need to condition all quantities involved in the copula on survival up to the previous time interval.
-For simulating survival outcome at the end of the $k$th interval—that is, $y_{k+1}$—(14) becomes:
-$p(y_{k+1}$ $\,|\,z,\bar{l}_{k},\bar{a}_{k},y_{k}=0)$
-$=\prod_{j=0}^{k}c^{*}_{Y_{k+1},L_{j}\mid\bar{L}_{j-1},Z,\bar{A}_{k},Y_{k}=0}\big{\{}F(y_{k+1}\,|\,\bar{l}_{j-1},z,do(\bar{a}_{k}),Y_{k}=0),$
-$\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},z,do(\bar{a}_{k}),Y_{k}=0);\Theta_{k,j}\}.$ (18)
+The general approach remains similar to the non-survival case, but with the key difference that at each time step $k$, simulation is now conditioned on survival up to the start of the current time interval.
+For brevity, we continue to omit conditioning on $Z$.
+Unlike the derivation of outcome simulation for the non-survival case in Section 4.2, all quantities involved in the copulas must now be conditioned on survival up to this point.
+Specifically, when simulating the incremental failure time, $\tilde{y}_{k+1}$, given $Y_{k}=0$, (7) becomes:
+$p(\tilde{y}_{k+1}$ $\,|\,\bar{l}_{k},\bar{a}_{k},y_{k}=0)$
+$=\prod_{j=0}^{k}c^{*}_{\tilde{Y}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},Y_{k}=0}\big{\{}F(\tilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0),$
+$\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0);\Theta_{k,j}\}.$ (10)
 It is important to note the distinction between the subscripts $j$ and $k$ in the product.
-To reiterate, our goal is to simulate the outcome at the end of the $k$th interval, that is the $k$th panel in Figure 3.
-Each copula in (18) associates the outcome $Y_{k+1}$ with all the prior values of the time-varying confounder $L_{j}$, for $j=0,1,\dots,k$.
-In the non-survival case discussed in Section 4.2, the quantiles, $F(l_{j}\,|\,\bar{l}_{j-1},z,\bar{a}_{j-1})$, for $j=0,1,\dots,k$, are directly available at each simulation at step $j$ and can be simply plugged into the iterative application of inverse $h$-functions without further computation.
+To reiterate, our goal is to simulate the outcome at the end of the $k$th interval, which corresponds to the $k$th panel in Figure 2.
+Each copula in (10) associates $\tilde{y}_{k+1}$ with all the prior values of the time-varying confounder $L_{j}$, for $j=0,\ldots,k$.
+In the non-survival case discussed in Section 4.2, the quantiles $F(l_{j}\,|\,\bar{l}_{j-1},\bar{a}_{j-1})$, for $j=0,\ldots,k$, are directly available at each simulation at step $j$ and can be simply plugged into the iterative application of inverse $h$-functions without further computation.
 However, this convenience does not extend when the quantiles are conditioned on survival at later time points.
-For instance, suppose that at $k=0$, we simulate baseline health, $P(L_{0})\sim N(0,1)$.
+For instance, suppose that at $k=0$, we simulate baseline health, $P(L_{0})\sim\operatorname{N}(0,1)$.
 Conditioning on survival after 12 months, however, alters the distribution.
-Survivors’ baseline health likely represents a healthier subgroup than the initial population, meaning the distribution is no longer standard normal.
-This means that the required conditional quantiles of $L_{k}$ are not directly available and require additional computation.
+The baseline health of the survivors will almost certainly not be representative of the initial population, meaning that the distribution is no longer standard normal.
+This means that the conditional quantiles of $L_{j}$ that we need are not directly available and require additional computation.
 Fortunately, employing copulas to define the dependence structure offers an efficient method to derive these conditional distributions, eliminating the need for resource-intensive techniques such as Monte Carlo approximation or numerical integration.
 The proposed simulation algorithm follows an iterative structure and is formally presented and validated using an inductive approach.
-Proposed simulation method: At time step $k$, we specify the following inputs:
-$F(L_{k}\,|\,\bar{L}_{k-1},Y_{k}=0)$ $=w_{k,1}$
-$F(A_{k}\,|\,\bar{L}_{k},Y_{k}=0)$ $=w_{k,2}$
-$F(Y_{k+1}\,|\,\bar{L}_{k},do(\bar{A}_{k}),Y_{k}=0)$ $=w_{k,3},$ (19)
-where $w_{k,1},w_{k,2},w_{k,3}\sim U(0,1)$.
-We also have parameters $\Theta_{k,j}$ for each copula $C_{Y_{k+1}L_{j}\mid\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0}$ and from the previous step we assume that—by induction—we are given quantiles for each $L_{j}\mid\overline{L}_{j-1},\overline{A}_{j-1},Y_{k-1}=0$.
-These inputs include the distributions and quantiles of ‘the past’ (i.e.
-$L_{k}$ and $A_{k}$) given survival, the quantile of the the outcome, $Y_{k+1}$, given the past, and $k+1$ bivariate copulas that ‘weave’ the dependency between the outcome $Y_{k+1}$ and the history of $L$.
-Crucially, we do not need to compute the full conditional distribution of $Y_{k+1}L_{k}\mid\bar{L}_{k-1},do(\bar{A}_{k}),Y_{k}=0$, but rather just a quantile corresponding to our simulated value; this is central to the efficiency of our approach.
-We then compute the following for each $j=0,\ldots,k-1$:
+Proposed simulation method: At time step $k$, we specify the cumulative distribution functions:
+$F_{L_{k}\mid\bar{L}_{k-1},\bar{A}_{k-1},Y_{k}=0}$ $F_{A_{k}\mid\bar{L}_{k},\bar{A}_{k-1},Y_{k}=0}$ $F_{\tilde{Y}_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}.$
+Additionally, we define $k+1$ bivariate copulas $C_{\tilde{Y}_{k+1}L_{j}\mid\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0}$, parameterized by $\Theta_{k,j}$ for $j=0,\ldots,k$.
+These copulas captures the dependence between $\tilde{Y}_{k+1}$ and each preceding time-varying covariate.
+Draw $w_{k,1}$, $w_{k,2}$, $w_{k,3}\sim\operatorname{U}(0,1)$ and compute:
+$l_{k}=F^{-1}_{L_{k}\mid\bar{L}_{k-1},\bar{A}_{k-1},Y_{k}=0}(w_{k,1}\,|\,\bar{l}_{k-1},\bar{a}_{k-1})$ $a_{k}=F^{-1}_{A_{k}\mid\bar{L}_{k},\bar{A}_{k-1},Y_{k}=0}(w_{k,2}\,|\,\bar{l}_{k},\bar{a}_{k-1}),$ (11)
+to obtain a sample $(l_{k},a_{k})$. We initialize by setting $F(\tilde{y}_{k+1}\,|\,\bar{l}_{k},do(\bar{a}_{k}),y_{k}=0)=w_{k,3}$. Then, for $j=k-1,\ldots,0$, we iteratively apply the transformations:
 $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$ $=h\big{\{}F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0),$
-$\qquad\quad F(Y_{k}=0\,|\,\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k-1}=0);\Theta_{k-1,j}\big{\}}$ (20)
-$F(Y_{k+1}\,|\,\bar{L}_{j},do(\bar{A}_{k}),Y_{k}=0)$ $=h^{-1}\big{\{}F(Y_{k+1}\,|\,\bar{L}_{j+1},do(\bar{A}_{k}),Y_{k}=0),$
-$\qquad\qquad F(L_{j+1}\,|\,\bar{L}_{j},do(\bar{A}_{k}),Y_{k}=0);\Theta_{k,j+1}\big{\}},$ (21)
-before
-$Y_{k+1}=F^{-1}\big{\{}F(Y_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0)\big{\}}.$ (22)
+$\qquad\quad F(\tilde{y}_{k}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0);\Theta_{{k-1},j}\big{\}}$ (12)
+$F(\tilde{y}_{k+1}\,|\,\bar{l}_{j},do(\bar{a}_{k}),y_{k}=0)$ $=h^{-1}\big{\{}F(\tilde{y}_{k+1}\,|\,\bar{l}_{j+1},do(\bar{a}_{k}),y_{k}=0),$
+$\qquad\qquad F(l_{j+1}\,|\,\bar{l}_{j},do(\bar{a}_{k}),y_{k}=0);\Theta_{k,j+1}\big{\}}.$ (13)
+By induction, the quantiles in the $h$-function in (12) are available from the previous time step $k-1$. We prove this claim as Proposition 5.1.
+When the recursion reaches $j=0$, we obtain the quantile $F(\tilde{y}_{k+1}\,|\,l_{0},do(\bar{a}_{k}),y_{k}=0)$. Applying the $h$-function one final time, we set:
+$u_{k+1}=F(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}))$ $=h^{-1}\big{\{}F(\tilde{y}_{k+1}\,|\,l_{0},do(\bar{a}_{k}),y_{k}=0),$
+$\qquad\qquad F(l_{0}\,|\,do(\bar{a}_{k}),y_{k}=0);\Theta_{k,0}\big{\}}.$
+A sample of $\tilde{y}_{k+1}$ is then obtained via inversion:
+$\tilde{y}_{k+1}=F_{Y_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}^{-1}(u_{k+1}),$ (14)
+with the corresponding binary outcome $y_{k+1}=I(\tilde{y}_{k+1}<1)$.
+This completes the $k$th time step, yielding the sample $(l_{k},a_{k},\tilde{y}_{k+1},y_{k+1})$.
+If $y_{k+1}=1$, the simulation stops; otherwise, proceed to $k+1$.
+Crucially, our method avoids computing the full conditional distribution of $\tilde{Y}_{k+1}\mid\bar{L}_{k},do(\bar{A}_{k}),Y_{k}=0$, but rather requiring just a quantile corresponding to our simulated value; this is central to the efficiency of our approach.
 Equation (12) involves $k$ calculations of the distributions of each $L_{j}\,|\,\bar{L}_{j-1}$ among those who survived to time $k$, in which all the quantities, including quantiles and copulas, are available from the previous step $k-1$.
-In (21) we iteratively apply $k+1$ $h$-functions to obtain the marginal quantile of $F(Y_{k+1}\,|\,Y_{k}=0)$.
-Once this quantile is computed, we simulate $Y_{k}$ from the target marginal structural model via inversion, as in (22).
-In total we evaluate $42k+$ 4 functions (including the simulation of $L_{k}$ and $A_{k}$) for this step.
+In (13), we recursively apply $k+1$ inverse $h$-functions to obtain the marginal quantile of $F(\tilde{Y}_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0)$.
+Finally, we simulate $\tilde{Y}_{k+1}$ from the target marginal structural model via inversion.
+In total, this time step requires $42k+$ 4 functions, including the simulation of $L_{k}$ and $A_{k}$.
 We proceed to show that the proposed method works using mathematical induction.
 Proposition 5.1.
-Given the quantiles $w_{k,1},w_{k,2},w_{k,3}$ for each $k=0,\ldots,K-1$ from (19) and associated univariate distributions, together with copula parameters $\Theta_{k,j}$ for each $k=0,\ldots,K$ and $j=0,\ldots,k$, we can use the procedure outlined above to simulate a survival outcome according to the specified distributions.
+Given the quantiles $w_{k,1},w_{k,2},w_{k,3}$ for each $k=0,\ldots,K-1$ and associated univariate distributions, together with copula parameters $\Theta_{k,j}$ for each $k=0,\ldots,K$ and $j=0,\ldots,k$, we can use the procedure outlined above to simulate a survival outcome according to the specified distributions.
+The proof is given in Appendix B.
+Remark 5.2.
+To simulate a binary survival outcome directly from a discrete marginal structural model, such as that defined in (9), we adopt an approach analogous to Fan et al.
+(2017).
+Specifically, we assume that there exists a continuous latent random variable, $S_{k+1}$, whose distribution under the intervention $do(\bar{A}_{k})$ is $F_{S_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}$, such that the binary survival indicator is obtained by dichotomizing $S_{k+1}$ at a constant threshold $C$.
+This formulation allows dependencies between $Y_{k+1}$ and past covariates to be modelled via the latent variable $S_{k+1}$, and the simulation procedure follows the algorithm described above, replacing $\tilde{Y}_{k+1}$ by $S_{k+1}$.
+Given the marginal quantile $u_{k+1}$, inversion yields $s_{k+1}=F^{-1}_{S_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}(u_{k+1}\,|\,do(\bar{a}_{k}),y_{k}=0)$ Finally, the binary outcome is obtained as $y_{k+1}=I(s_{k+1}>C)$.
+5.2 Extensions and computational complexity
+In this basic setting, simulating $K$ time steps for each individual requires $\sum_{k=0}^{K-1}(2k+4)=K^{2}+3K$ calculations.
+This implies that the computational complexity grows quadratically with $K$.
+However, if an individual experiences failure at any time step, the simulation terminates early, meaning that this complexity represents the maximum possible computational cost per individual.
+A key extension involves simulating multiple outcomes, and the approach depends on whether the additional outcomes pertain to survival or non-survival events.
+For non-survival outcomes simulated from marginal structural models—as opposed to conditional distributions—the number of calculations to derive $F(\bm{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y^{\triangle}_{k}=0)$ increases proportionally with the number of outcomes, $m$.
+Here, $\bm{Y}_{k+1}$ is an $m$-dimensional vector, and $Y_{k}^{\triangle}$ denotes the primary survival outcome, with all other outcomes assumed to be non-survival.
+To incorporate dependence between outcomes given past variables, we must specify the joint distribution $\bm{Y}_{k+1}\,|\,\bar{L}_{k},\bar{A}_{k}$ and simulate from this using a pair-copula construction.
+The total number of calculations is $\sum_{k=0}^{K-1}\big{\{}(k+3)m+k+1\big{\}}=(K+5)Km/2+K(K+1)/2$, which scales linearly with $m$.
+For competing risks, such as death from other causes, additional calculations are needed to determine the distribution of $L$ given survival from all risks, as well as its distribution conditioned on the occurrence of each specific risk.
+Consequently, the computational cost scales proportionally with $m$.
+Furthermore, iterative evaluations of inverse $h$-functions are required to obtain marginal interventional quantiles for each risk.
+The total number of calculations is given by $\sum_{k=0}^{K-1}(2mk+3m+1)=(K^{2}+2K)m+K$ which grows linearly with $m$.
+The simulation algorithm for a setting with a single competing risk is described in Appendix C.
+Another important extension involves including multiple time-varying confounders, observed or unobserved, such that $\bm{L}$ becomes a vector of $p$ variables.
+The number of calculations required for the distributions of ¯$\bm{\bar{}}L_{k-1}$, and $F(\tilde{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0)$ among survivors at $k$ scales proportionally with $p$.
+The total number of computations becomes $\sum_{k=0}^{K-1}(2pk+3p+2)=K(K+2)p+2K$.
+Thus, the computational complexity grows linearly with $p$.
+In the example presented in Section 5.3, we simulate data with two time-varying confounders.
+All of these calculations require evaluations of $h$-functions and their inverses. For commonly used copulas with analytic $h$-functions, these evaluations are computationally efficient and fast.
+5.3 Illustrative example
+We use an example to demonstrate the proposed algorithm simulating longitudinal time-to-event data, following a setup similar to Seaman and Keogh (2024).
+We consider baseline variables $B$ and $C$, two time-varying covariates $L_{k}$ and $W_{k}$, and treatment $A_{k}$ at follow-up $k$.
+The time-to-event is $T$, with $Y_{k+1}=I(T<k+1)$ indicating failure status observed at the end of the $k$th interval.
+The auxiliary continuous variable $\tilde{Y}_{k+1}$ is defined as previously.
+The data-generating models for the baseline and time-varying covariates, and treatment status are detailed in Appendix D.
+We assume a Cox marginal structural model given by:
+$\lambda(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),c,y_{k}=0)=\lambda_{0}\exp\left(\beta_{C}\,c+\beta_{A}\,a_{k}+\beta_{CA}\,c\,a_{k}\right),$ (15)
+where $\lambda_{0}=\exp(-2)$ is the baseline hazard.
+We set $\beta_{C}=-0.5$ and vary the values of $(\beta_{A},\beta_{CA})$ to construct scenarios with different strengths of treatment: a strong effect scenario $(-0.5,0.3)$ and a weak effect scenario $(0.2,0.1)$.
+This parameterization assumes a constant hazard rate between follow-up visits, influenced by the treatment status at the most recent visit.
+This implies that given survival at follow-up time $k$, the incremental survival time $\tilde{Y}_{k+1}$ follows an exponential distribution with the rate equal to the hazard above:
+$\tilde{Y}_{k+1}\sim\operatorname{Exp}(\lambda_{0}\exp\left(\beta_{C}\,c+\beta_{A}\,a_{k}+\beta_{CA}\,c\,a_{k}\right)).$
+If $\tilde{Y}_{k+1}<1$ then $Y_{k+1}=1$, and 00 otherwise.
+We set the total number of follow-ups $K$ to 6.
+We set the copulas describing the association between $L_{j}$ ($j=0,\ldots,k$) and $Y_{k+1}$ to be Student t with parameter $\rho_{L}$ and 5 degrees of freedom, although each pairwise dependence could, in principle, be distinct.
+Similarly, we model the copulas capturing the dependence between $Y_{k+1}$ and each $W_{j}$ ($j=0,\ldots,k$), given $L_{k}$, as Gaussian with correlation $\rho_{W}$, though again we note that different copulas could be specified for each time point.
+We vary the values of $\rho_{L}$ and $\rho_{W}$ to construct scenarios with different strengths of confounding (see Table 6 in Appendix D).
+We simulate data for sample size $n=\{1000,2000,5000\}$.
+Remark 5.3.
+Table 7 (in Appendix D) reports the runtime required to simulate data under the setup for the large causal effect with medium confounding scenario.
+The simulations were carried out on a laptop using a single score without parallelization.
+Generating data for $n$ up to 10,000 units takes only a few seconds, while even for a very large sample size ($n=50,\!000$) and 11 time steps ($K=10$), the runtime remains under one minute.
+This computational efficiency is crucial for tasks such as method evaluation and sample size calculations in clinical trial design, where simulations must be repeated across various parameter values.
+We contrast two approaches for fitting the Cox marginal structural model specified in (15): stabilized inverse-probability-weighting and unweighted estimation.
+Details on weight calculation are provided in Appendix D.
+Table 4 summarizes the bias of estimates from 500 replications, compared to corresponding Monte Carlo standard errors as a measure of uncertainty.
+As the sample size increases, generally, the biases of estimates for $\beta_{C}$, $\beta_{A}$ and $\beta_{AC}$ obtained with inverse probability weights converge to zero, with their absolute values substantially smaller than the corresponding Monte Carlo standard errors.
+In contrast, the unweighted estimator exhibits considerable bias, particularly for the estimate of the treatment effect, $\beta_{A}$, which persists even at $n=5000$.
+The magnitude of this bias far exceeds the Monte Carlo standard errors.
+The unweighted estimates are most biased under the high level of confounding by $L$ and $W$, since this is not adjusted for.
+Level of confounding Effect size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ IPTW Unweighted IPTW Unweighted IPTW Unweighted High $1000$ $-0.04(2.27)$ $-1.05(1.54)$ $-0.14(2.31)$ $3.73(1.73)$ $0.27(3.09)$ $0.49(2.22)$ $\beta_{A}=-0.5$ $2000$ $-0.07(1.68)$ $-1.03(1.06)$ $-0.02(1.63)$ $3.81(1.21)$ $0.02(2.09)$ $0.34(1.51)$ $5000$ $0.02(1.25)$ $-0.96(0.68)$ $0.10(1.08)$ $3.89(0.76)$ $-0.09(1.58)$ $0.24(0.97)$ $1000$ $-0.03(2.35)$ $-0.93(1.59)$ $0.02(1.99)$ $3.64(1.46)$ $0.03(2.78)$ $0.44(1.94)$ $\beta_{A}=0.2$ $2000$ $-0.11(1.60)$ $-0.93(1.11)$ $0.02(1.49)$ $3.65(1.18)$ $0.17(2.02)$ $0.49(1.51)$ $5000$ $-0.03(1.14)$ $-0.95(0.71)$ $0.04(1.01)$ $3.65(0.71)$ $0.07(1.36)$ $0.51(0.89)$ Medium $1000$ $-0.07(2.04)$ $-0.43(1.51)$ $-0.13(2.19)$ $1.90(1.73)$ $0.23(2.91)$ $0.08(2.22)$ $\beta_{A}=-0.5$ $2000$ $-0.02(1.60)$ $-0.35(1.01)$ $0.00(1.65)$ $1.95(1.22)$ $0.09(2.27)$ $-0.01(1.55)$ $5000$ $-0.09(1.10)$ $-0.42(0.64)$ $-0.01(1.21)$ $1.97(0.80)$ $0.07(1.59)$ $-0.02(1.03)$ $1000$ $-0.12(2.06)$ $-0.36(1.43)$ $-0.04(2.02)$ $1.75(1.54)$ $0.14(2.70)$ $0.09(1.92)$ $\beta_{A}=0.2$ $2000$ $-0.10(1.62)$ $-0.42(1.08)$ $0.03(1.49)$ $1.68(1.11)$ $0.09(2.09)$ $0.17(1.44)$ $5000$ $-0.03(1.04)$ $-0.39(0.66)$ $0.03(0.95)$ $1.78(0.68)$ $0.03(1.32)$ $0.14(0.88)$ Low $1000$ $0.01(1.82)$ $0.46(1.29)$ $-0.23(2.36)$ $-1.95(1.83)$ $0.21(2.90)$ $-0.20(2.20)$ $\beta_{A}=-0.5$ $2000$ $-0.10(1.32)$ $0.44(0.89)$ $-0.04(1.69)$ $-1.83(1.25)$ $0.07(2.13)$ $-0.30(1.59)$ $5000$ $-0.02(1.04)$ $0.47(0.58)$ $0.01(1.07)$ $-1.77(0.84)$ $0.03(1.46)$ $-0.32(0.99)$ $1000$ $0.04(1.98)$ $0.53(1.30)$ $0.05(1.98)$ $-1.88(1.46)$ $-0.03(2.65)$ $-0.31(1.88)$ $\beta_{A}=0.2$ $2000$ $-0.02(1.45)$ $0.43(0.99)$ $0.04(1.43)$ $-1.85(1.10)$ $0.01(1.87)$ $-0.26(1.38)$ $5000$ $-0.01(0.92)$ $0.43(0.59)$ $0.07(0.95)$ $-1.89(0.68)$ $-0.06(1.25)$ $-0.25(0.84)$
+Table 4: Bias and Monte Carlo standard errors (×$\times 10$) of estimates produced by inverse probability of treatment weighted (‘IPTW’) and unweighted Cox regressions, under different strengths of confounding, effect sizes, and sample sizes.
+Table 5 presents coverage rates of the 95% confidence intervals.
+Confidence intervals for the weighted estimated, computed using both sandwich variance estimation and bootstrap resampling, achieve coverage close to the nominal level, with slight undercoverage.
+In contrast, the sandwich-based confidence intervals for the unweighted estimator perform poorly, especially for the treatment effect estimate, $\beta_{A}$.
+Level of confounding Effect Size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ W-S W-B UW W-S W-B UW W-S W-B UW High 1000 93.8% 95.2% 88.4% 93.4% 92.8% 42.2% 93.8% 93.8% 93.4% $\beta_{A}=-0.5$ 2000 94.0% 94.0% 84.4% 95.2% 95.6% 12.4% 95.6% 95.6% 94.6% 5000 93.2% 92.8% 70.2% 94.4% 93.4% 0.2% 94.2% 93.6% 93.8% 1000 92.8% 93.8% 91.8% 96.2% 95.8% 31.6% 93.4% 94.0% 95.6% $\beta_{A}=0.2$ 2000 95.2% 95.0% 87.6% 94.6% 95.0% 9.8% 94.8% 94.6% 92.2% 5000 93.6% 93.6% 72.4% 93.8% 93.0% 0.2% 94.0% 94.0% 90.0% Medium 1000 93.8% 94.2% 93.4% 94.8% 93.6% 79.6% 93.8% 94.0% 93.6% $\beta_{A}=-0.5$ 2000 93.0% 93.8% 92.4% 95.2% 94.6% 64.8% 92.2% 93.4% 96.2% 5000 95.4% 94.8% 90.0% 93.2% 92.6% 30.2% 93.0% 93.8% 93.6% 1000 94.2% 94.8% 95.0% 95.0% 93.8% 77.2% 94.2% 93.2% 95.8% $\beta_{A}=0.2$ 2000 94.6% 93.8% 91.6% 93.6% 93.2% 64.0% 92.8% 92.2% 94.2% 5000 95.4% 94.6% 91.6% 95.4% 95.2% 26.8% 94.8% 94.4% 95.4% Low 1000 95.4% 95.4% 93.0% 95.0% 95.2% 84.0% 95.2% 95.2% 94.6% $\beta_{A}=-0.5$ 2000 95.6% 96.4% 93.0% 95.8% 96.0% 71.6% 95.0% 95.6% 94.6% 5000 93.2% 93.2% 87.2% 94.2% 94.8% 42.6% 94.0% 94.4% 94.4% 1000 94.6% 95.6% 94.8% 96.4% 96.6% 77.4% 92.8% 94.8% 96.0% $\beta_{A}=0.2$ 2000 94.8% 95.6% 92.8% 94.4% 94.6% 60.6% 95.0% 95.2% 94.2% 5000 94.8% 94.2% 89.4% 95.0% 95.4% 20.0% 95.0% 94.2% 93.8%
+Table 5: Coverage rates for IPTW-Sandwich (W-S), IPTW-Bootstrap (W-B), and unweighted (UW) estimators of $\beta_{C}$, $\beta_{A}$, and $\beta_{AC}$ under different strengths of confounding, effect sizes, and sample sizes.
+In summary, the asymptotically unbiased nature of the inverse probability weighted estimates, compared to the biased unweighted estimates, further confirms that the simulated data were generated in accordance with the Cox model in (15), and that confounding by $L$ and $W$ was appropriately introduced.
+6 Discussion
+In this paper, we proposed and demonstrated a principled strategy for simulating longitudinal data exactly from marginal structural models, leveraging the frugal parameterization (Evans and Didelez, 2024).
+Our proposed method is highly flexible and can accommodate a variety of practical scenarios, including multiple treatments (for example, primary and drop-in), dynamic treatment regimes, dose-response analyses, multiple outcomes (such as secondary outcomes and competing risks), and censoring.
+Additionally, it imposes no restrictions on the data-generating process or marginal structural causal models.
+In Section 5, we extended this method to survival models and time-to-event outcomes, which are common in clinical trials.
+Our methods are both conceptually and computationally simple, which crucial for practical applications such as sample size calculation.
+[…]
+Appendix A Extension to Include Time-Varying Latent Variables
+We provide an example of how to incorporate a latent process that confounds the covariates and the final outcome following the algorithm outlined in Section 4.
+The steps to simulate longitudinal data from the model depicted in Figure 3 are as follows:
+[…]
+Figure 3: Extension of the model in Figure 1 which allows for an unobserved process $H_{k}$ affecting only $Z$, $L_{k}$ and $Y$.
+  1. 1.
+    Baseline simulation: The directed acyclic graph in Figure 3 indicates that the latent variable $H_{0}$ precedes $Z$ in topological order, potentially indicating an initial selection bias.
+Accordingly, we simulate baseline values $(z,h_{0},l_{0},a_{0})$ sequentially:
+    $h_{0}$ $=F_{H_{0}}^{-1}(w_{0,0})$
+    $z$ $=F_{Z\mid H_{0}}^{-1}(w_{Z}\,|\,h_{0})$
+    $l_{0}$ $=F_{L_{0}\mid ZH_{0}}^{-1}(w_{0,1}\,|\,z,h_{0})$
+    $a_{0}$ $=F_{A_{0}\mid ZH_{0}L_{0}}^{-1}(w_{0,2}\,|\,z,h_{0},l_{0}),$
+    where $w_{Z},w_{0,0},w_{0,1},w_{0,2}$ are standard uniform variables.
+  2. 2.
+    Time-step simulation: For each time step $k=1,\dots,K$, simulate:
+    $h_{k}$ $=F_{H_{k}\mid\bar{H}_{k-1}}^{-1}(w_{k,1}\,|\,\bar{h}_{k-1})$
+    $l_{k}$ $=F_{L_{k}\mid Z\bar{H}_{k}\bar{L}_{k-1}\bar{A}_{k-1}}^{-1}(w_{k,2},\,|\,z,\bar{h}_{k},\bar{l}_{k-1},\bar{a}_{k-1})$
+    $a_{k}$ $=F_{A_{k}\mid Z\bar{H}_{k}\bar{L}_{k}\bar{A}_{k-1}}^{-1}(w_{k,3},\,|\,z,\bar{h}_{k},\bar{l}_{k},\bar{a}_{k-1})$
+    where $w_{k,1},w_{k,2},w_{k,3}\sim U(0,1)$.
+  3. 3.
+    Outcome simulation: Once all time-varying covariates and treatment have been simulated, we apply the same derivation following the frugal parameterization as in Section 4.2 and simulate $Y$ in two steps:
+    1. (a)
+      Compute the marginal quantile:
+      $u^{*}_{Y}$ $=h^{-1}\big{[}\dots h^{-1}[h^{-1}\left\{h^{-1}\left(w_{y},F(l_{K}\,|\,\bar{l}_{K-1},\bar{h}_{K},z,\bar{a}_{K-1});\Theta_{K,1}\right),F(h_{K}\,|\,\bar{l}_{K-1},\bar{h}_{K-1},z,\bar{a}_{K-1});\Theta_{K,2}\right\}$
+      $F(l_{K-1}\,|\,\bar{l}_{K-2},\bar{h}_{K-1},z,\bar{a}_{K-2});\Theta_{K-1,1}]\dots;\Theta_{0,2}\big{]}$
+      where $w_{Y}\sim\operatorname{U}(0,1)$;
+    2. (b)
+      Obtain the outcome: $y=F_{Y\mid Z,\bar{A}_{K}}^{*{-1}}(u^{*}_{Y}\,|\,z,do(\bar{a}_{K}))$.
+Compared to the case where without latent variable, the number of applications of inverse $h$-functions in Step 3(a) doubles.
+These transformations systematically disentangle the dependencies between $Y$ and $H_{k}$, as well as between $Y$ and $L_{k}$.
+In addition to the information required in the case without latent confounders, the user also specifies:
+  • the development of the latent confounder $H_{k}$ as influenced by $Z$ and treatment history $\bar{A}_{k}$;
+  • $K+1$ pair-copula densities describing dependencies between the outcome and each member of the latent process: $c_{YH_{k}\mid Z,\bar{H}_{k-1},\bar{L}_{k-1},do(\bar{A}_{k-1})}$ for $k=0,1,\ldots,K$.
+Extending the framework to include additional time-varying covariates follows a similar approach.
+In essence, simulating a latent process is technically equivalent to simulating any other time-varying confounder.
+Whether a confounder is observed or unobserved does not alter the fundamental simulation methodology.
+With common choices of parametric bivariate copulas, evaluating the $h$-functions is fast, making the simulation efficient.
+As the dimensionality of time-varying confounders, either observed or unobserved, the computational complexity grows linearly.
+Appendix B Proof of Proposition 5.1
 Proof.
-Base case: For $k=0$, initialize the algorithm by specifying:
-$F(L_{0})=w_{0,1}$
-$F(A_{0})=w_{0,2}$
-$F(Y_{1}\,|\,do(A_{0}),L_{0})=w_{0,3}$
-$C_{Y_{1}L_{0},do(A_{0})}:\Theta_{0,0}.$
-Using these inputs, we calculate
-$F(Y_{1}\,|\,do(A_{0}))$ $=h^{-1}\big{\{}F(Y_{1}\,|\,L_{0}),F(L_{0});\Theta_{0,0}\big{\}}$ (23)
-$Y_{1}$ $=F^{-1}\big{\{}F(Y_{1}\,|\,do(A_{0})\big{\}}.$
-For the next time step ($k=1$), the algorithm requires:
-$F(L_{0}\,|\,do(\bar{A}_{1}),Y_{1}=0)$ $=h\big{\{}F(L_{0}),F(Y_{1}=0\,|\,do(A_{0}));\Theta_{0,0}\big{\}}$ (24)
-$F(Y_{2}\,|\,L_{0},do(\bar{A}_{1}),Y_{1}=0)$ $=h^{-1}\big{\{}F(Y_{2}\,|\,\bar{L}_{1},do(\bar{A}_{1}),Y_{1}=0),F(L_{1}\,|\,L_{0},do(\bar{A}_{1}),Y_{1}=0);\Theta_{1,1}\big{\}}$
-$F(Y_{2}\,|\,Y_{1}=0)$ $=h^{-1}\big{\{}F(Y_{2}\,|\,L_{0},do(\bar{A}_{1}),Y_{1}=0),F(L_{0}\,|\,do(\bar{A}_{1}),Y_{1}=0);\Theta_{1,0}\big{\}}.$
-In (24) we derive the distribution of $L_{0}$ among survivors at $k=1$, using $F(Y_{1}=0)$, which is already computed in (23) in step $k=0$.
+We prove that the algorithm correctly simulates data for all $k\geq 0$.
+Base case ($k=0$):
+Initialize the algorithm by specifying the cumulative distribution functions $F_{L_{0}}$, $F_{A_{0}\mid L_{0}}$, $F_{\tilde{Y}_{1}\mid do(A_{0})}$, and the copula $C_{\tilde{Y}_{1}L_{0}\mid do(A_{0})}$, parameterized by $\theta_{0,0}$.
+Draw $w_{0,1}$, $w_{0,2}$, $w_{0,3}\sim U(0,1)$ and compute:
+$l_{0}=F^{-1}_{L_{0}}(w_{0,1})$ $a_{0}=F^{-1}_{A_{0}\mid L_{0}}(w_{0,2}\,|\,l_{0}),$ (16)
+to obtain a sample $(l_{0},a_{0})$. Using the transformation:
+$F_{\tilde{Y}_{1}\,|\,do(A_{0})}(\tilde{y}_{1}\,|\,do(a_{0}))$ $=h^{-1}\big{\{}F_{\tilde{Y}_{1}\,|\,L_{0},do(A_{0})}(\tilde{y}_{1}\,|\,l_{0},do(a_{0})),F_{L_{0}}(l_{0});\theta_{0,0}\big{\}},$ (17)
+set $w_{0,3}=F_{\tilde{Y}_{1}\,|\,L_{0},do(A_{0})}(\tilde{y}_{1}\,|\,l_{0},do(a_{0}))$ and substitute $F_{L_{0}}(l_{0})$ with $w_{0,1}$, yielding:
+$u_{1}=h^{-1}\left(w_{0,3},w_{0,1};\theta_{0,0}\right).$ (18)
+By inversion, compute $\tilde{y}_{1}=F_{Y_{1}\mid do(A_{0})}^{-1}(u_{1})$, and the corresponding binary outcome $y_{1}=I(\tilde{y}_{1}<1)$.
+This completes the first iteration, yielding a sample $(l_{0},a_{0},\tilde{y}_{1},y_{1})$.
+If $y_{1}=1$, stop; otherwise, proceed to $k=1$.
+For $k=1$, define the necessary cumulative distribution functions $F_{L_{1}\,|\,L_{0},A_{0}}$, $F_{A_{1}\mid\bar{L}_{1},A_{0}}$, $F_{\tilde{Y}_{2}\mid do(\bar{A}_{1})}$, and copulas $C_{\tilde{Y}_{2}L_{1}\mid L_{0}do(\bar{A}_{1})}$ and $C_{\tilde{Y}_{2}L_{0}\mid do(A_{0})}$, parameterized by $\theta_{1,1}$ and $\theta_{1,0}$, respectively.
+These two copulas associate $\tilde{Y}_{2}$ with previous time-varying confounders.
+Draw $w_{1,1}$, $w_{1,2}$, $w_{1,3}$ from $U(0,1)$, and obtain $(l_{1},a_{1})$ as before.
+We now need to obtain a sample of $\tilde{y}_{2}$. Calculate:
+$F(l_{0}\,|\,do(a_{0}),y_{1}=0)$ $=h\big{\{}F(l_{0}),F(\tilde{y}_{1}\,|\,do(a_{0}));\theta_{0,0}\big{\}}$ (19)
+$F(\tilde{y}_{2}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0)$ $=h^{-1}\big{\{}F(\tilde{y}_{2}\,|\,\bar{l}_{1},do(\bar{a}_{1}),y_{1}=0),F(l_{1}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0);\theta_{1,1}\big{\}}$
+$F(\tilde{y}_{2}\,|\,do(a_{1}),y_{1}=0)$ $=h^{-1}\big{\{}F(\tilde{y}_{2}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0),F(l_{0}\,|\,do(\bar{a}_{1}),y_{1}=0);\theta_{1,0}\big{\}}.$
+Note that we omit some subscripts on quantiles for brevity.
+In (19) we derive the quantile of $L_{0}$ given $a_{0}$ among survivors the start of the $k=1$ interval, where $F(l_{0})=w_{0,0}$ and $F(\tilde{y}_{1}\,|\,do(a_{0})$ is already computed in (18) in step $k=0$.
 All other necessary quantities are specified at $k=1$.
 Therefore, the base case is satisfied since the algorithm simulates the first time step and provides the required inputs for the next.
-Inductive step: Assumes that we have reached the $k$th iteration of the algorithm.
-We are given $F(L_{k}\mid\overline{L}_{k-1},do(\bar{A}_{k-1}),Y_{k}=0)$ and $F(Y_{k+1}\mid\overline{L}_{k},do(\bar{A}_{k}),Y_{k}=0)$, as well as a collection of pair copulas between $Y_{k+1}$ and $L_{j}$ with parameters $\Theta_{k+1,j}$.
-We need to compute the updated distributions $F(L_{j}\mid\overline{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0)$ for each $j<k$, given (from the $(k-1)$th iteration) $F(L_{j}\mid\overline{L}_{j-1},do(\bar{A}_{k-1}),Y_{k-1}=0)$ and $F(Y_{k}\mid\overline{L}_{j},do(\bar{A}_{k-1}),Y_{k-1}=0)$.
-This can be achieved by noting that $F(L_{j}\mid\overline{L}_{j-1},do(\bar{A}_{k-1}),Y_{k-1}=0)=F(L_{j}\mid\overline{L}_{j-1},do(\bar{A}_{k}),Y_{k-1}=0)$, and then applying (20).
-We must then obtain $F(Y_{k+1}\mid\overline{L}_{j},do(\bar{A}_{k}),Y_{k}=0)$ for each $j$ to use in the next iteration.
-This can be done using (21), and hence, we can move to the $(k+1)$th iteration.
-By the principle of induction, the statement is true for all $k\geq 0$.
-∎
-In this basic setting, simulating $K$ time steps for each individual requires $\sum_{k=0}^{K-1}(2k+4)=K^{2}+3K$ calculations.
-This implies that the computational complexity grows quadratically with $K$.
-If an individual ceases to survive at any time step, the simulation is terminated for them.
-Building upon this basic case, we analyze various extensions:
-  • Multiple time-varying confounders: If the number of time-vary confounders increases, making $\bm{L}$ a vector of $p$ variables, the number of calculations for the distributions of ¯$\bm{\bar{}}L_{k-1}$, and $F(Y_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0)$ among the survivors at $k$ scales proportionally with $p$.
-The total number of computations becomes $\sum_{k=0}^{K-1}(2pk+3p+2)=pK^{2}+(2p+2)K$.
-Hence the computational complexity grows linearly with $p$.
-In the demonstration presented in the next subsection, we simulate data where there are two time-varying confounders.
-  • Multiple outcomes: The approach to simulating additional outcomes depends on whether they are are survival-related or non-survival-related.
-    • For non-survival outcomes, such as weight loss, simulated from marginal structural models – as opposed to from conditional distributions) – the number of calculations to derive $F(\bm{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y^{\triangle}_{k}=0)$ increases proportionally with the number of outcomes, $m$.
-Here, $\bm{Y}_{k}$ is a m-dimensional vector, and $Y_{k}^{\triangle}$ denotes the primary survival outcome.
-Specifically, the total number of calculations becomes $\sum_{k=0}^{K-1}(pmk+pk+pm+p+m+1)=\frac{mp+p}{2}K^{2}+(mp+m+p+1)K$.
-    • For survival outcomes, such as competing risks (e.g., death from other causes), a hierarchical simulation approach is adopted.
-This involves defining a hypothetical sequence of events, where any order can be chosen, though some may be more intuitive than others.
-For instance, the simulation first evaluates death from the primary cause; if an individual survives, it then assesses death from other causes.
-The distribution of ¯$\bm{\bar{}}L_{k-1}$ is calculated sequentially conditional on surviving from both risks according to this order.
-As a result, the number of calculations scales with $m$.
-Additionally, similar to the non-survival case, deriving $(F(\bm{Y}_{k+1}\mid do(\bar{A}_{k}),\bm{Y}_{k}=0)$ requires $m$–fold more computations.
-In total, $\sum_{k=0}^{K-1}(pmk+pmk+pm+p+m+1)=mpK^{2}+(mp+m+p+1)K$ calculations are needed.
-    In both scenarios, the computational complexity grows linearly with the number of outcomes, $m$.
-  • Censoring: It is customary to specify and simulate censoring $M_{k}$ using a full conditional distribution $F(M_{k+1}\,|\,\bar{L}_{k},do(\bar{A}_{k}),\bm{Y}_{k}=0)$, which represents the likelihood of an individual being censored given their past history.
-This requires adding one additional calculation at the end of each time step to simulate $M_{k+1}$.
-These calculations are all evaluations of $h$-functions and their inverses. For commonly used copulas with analytic $h$-functions, these evaluations are generally efficient and computationally fast.
-6.2 Demonstration
-We use an example to demonstrate the proposed algorithm simulating longitudinal time-to-event data. This setup broadly follows the simulation study in Seaman and Keogh (2024).
-We consider baseline variables $B$, $C$ and two time-varying covariates $L$ and $W$.
-$A_{k}$ indicates treatment at follow-up $k$.
-Let the time to event be $T$ and $Y_{k}=I(T>k)$ be the survival status observed at each visit.
-Specifically we choose that at baseline:
+Inductive step:
+Assume the algorithm correctly simulates up to the $k$th iteration.
+We have $F(l_{k}\mid\overline{l}_{k-1},do(\bar{a}_{k-1}),y_{k}=0)$ and $F(y_{k+1}\mid\overline{l}_{k},do(\bar{a}_{k}),y_{k}=0)$, as well as a collection of pair copulas between $Y_{k+1}$ and $L_{j}$ given $(\bar{l}_{j-1},\bar{a}_{j-1})$ with parameters $\theta_{k,j}$.
+We need to compute the updated quantiles $F(l_{j}\mid\overline{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$ for each $j<k$, given from the previous iteration:
+$F(l_{j}\mid\overline{l}_{j-1},do(\bar{a}_{k-1}),y_{k-1}=0),$ $F(y_{k}\mid\overline{l}_{j},do(\bar{a}_{k-1}),y_{k-1}=0).$
+This can be achieved by noting that
+$F(l_{j}\mid\overline{l}_{j-1},do(\bar{a}_{k-1}),y_{k-1}=0)=F(l_{j}\mid\overline{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0),$
+which allows us to apply (12) to compute the required quantiles.
+We must then obtain $F(y_{k+1}\mid\overline{l}_{j},do(\bar{a}_{k}),y_{k}=0)$ for each $j$ to use in the next iteration. This can be done using (13), and hence, we can move to the $(k+1)$th iteration.
+By the principle of induction, the statement is true for all $k\geq 0$. ∎
+Appendix C Competing risk
+At each time step $k$, after simulating time-varying covariate $L_{k}$ and treatment $A_{k}$, we generate the primary outcome $Y_{k}$ and a competing risk $D_{k}$ using their respective auxiliary variables $\tilde{Y}_{k}$ and $\tilde{D}_{k}$, following the approach outlined in Section 5.
+For brevity, baseline covariates $\bm{Z}$ continue to be omitted, but in settings where they are present, they can be simulated as the initial step and included in the conditioning set of all subsequent quantities.
+Under this simulation framework, we specify:
+  • A marginal structural model for the primary event: $F^{*}(\tilde{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0,D_{k}=0)$;
+  • A marginal structural model for the competing event: $F^{*}(\tilde{D}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0,D_{k}=0)$;
+  • A bivariate copula parameterized by $\Psi_{k+1}$ to model the dependence between the above interventional margins of $\tilde{Y}_{k+1}$ and $\tilde{D}_{k+1}$;
+  • Bivariate copulas that ‘weave’ the multivariate dependence between $Y_{k+1}$ and past covariates $\bar{L}_{k}$, given $do(\bar{A}_{k})$, $Y_{k}=0$ and $D_{k}=0$, parameterized by $\Theta_{k,j}$ for $j=0,\ldots,k$;
+  • Bivariate copulas that ‘weave’ the multivariate dependence between $\tilde{D}_{k+1}$ and $\bar{L}_{k}$ conditional on $\tilde{Y}_{k+1}$, $do(\bar{A}_{k})$, $Y_{k}=0$ and $D_{k}=0$, parameterized by $\Phi_{k,j}$ for $j=0,\ldots,k$.
+The joint outcome model $p(\tilde{y}_{k+1},\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0)$ factorizes as follows:
+Primary event model:
+$p(\tilde{y}_{k+1}$ $\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0)$
+$=p^{*}(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)\times$
+$\prod_{j=0}^{k}c^{*}_{\tilde{Y}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},Y_{k}=0,D_{k}=0}\big{\{}F(\tilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0),$
+$\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0);\Theta_{k,j}\}.$ (20)
+Competing event model:
+$p(\tilde{d}_{k+1}$ $\,|\,\bar{l}_{k},\bar{a}_{k},\tilde{y}_{k+1},D_{k}=0)$
+$=p(\tilde{d}_{k+1}\,|\,do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)\times$
+$\prod_{j=0}^{k}c^{*}_{\tilde{D}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},\tilde{Y}_{k+1}D_{k}=0}\big{\{}F(d_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0),$
+$\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0);\Phi{k,j}\}.$ (21)
+The first term in (21), the distribution of $\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},\tilde{y}_{k+1},D_{k}=0$ can be obtained through:
+$F(\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},\tilde{y}_{k+1},D_{k}=0)=h\big{\{}F*(\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0),F^{*}(\tilde{y}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0)\big{\}}.$
+The challenge lies in computing $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$ in (20), and $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$ in (21), for $j=0,\ldots,k$.
+The former can be calculated in the same iterative fashion as described in Section 5.
+The latter can be derived as:
+$F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$ $=h\big{\{}F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0),$
+$\qquad F(\tilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0);\Theta_{k,j}\big{\}}.$ (22)
+Notably, the two components in the $h$-function in (22) exactly correspond to the those in the primary event model in (20), except with the order of conditioning swapped.
+Since these values are computed during the simulation of $\tilde{y}_{k+1}$, requiring only $k+1$ additional evaluations of the $h$-functions in (22).
+The algorithm for simulating each time step $k$ for an individual can be summarized into:
+  1. 1.
+    Simulate the time-varying confounder $L_{k}$ and treatment $A_{k}$ from their respective distributions.
+  2. 2.
+    Compute $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$ for $j=0,\dots,k-1$.
+  3. 3.
+    Compute $F(y_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$ for $j=1,\dots k$, leading to $u_{k+1}=F(\tilde{y}_{k+1}\,|\,z,do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$.
+  4. 4.
+    Compute: $\tilde{y}_{k+1}=F^{*-1}_{\tilde{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0,D_{k}=0}(u_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$.
+  5. 5.
+    Compute $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$ for $j=0,\ldots,k$ by evaluating $k+1$ $h$-functions following (22).
+  6. 6.
+    Compute $F(\tilde{d}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1}=0,D_{k}=0)$ for $j=1,\ldots k$, leading to $v_{k+1}=F(\tilde{d}_{k+1}\,|\,do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$.
+  7. 7.
+    Compute: $\tilde{d}_{k+1}=F_{\tilde{D}_{k+1}\,|\,do(\bar{A}_{k}),\tilde{Y}_{k+1},D_{k}=0}^{*-1}(\tilde{d}_{k+1}\,|\,do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$.
+This approach requires at most $74k+$ 7 evaluations of analytic functions per individual per time step. An example implementation is available at https://github.com/XiLinStats/LongSiMSM.
+Appendix D Additional details of the illustration in Section 5.3
+Models for baseline, time-varying covariates and treatment
+A baseline:
 $B\sim\operatorname{Exp}(2)$ $L_{0}\sim\operatorname{Gamma}(k=1+0.5\,B+0.5\,C,\;\theta=1)$
-$C\sim\operatorname{Bernoulli}(0.7)$ $W_{0}\sim\operatorname{Bernoulli}(\operatorname{expit}\{-0.2+0.5\,B+0.5\,C\})$
+$C\sim\operatorname{Bernoulli}(0.7)$ $W_{0}\sim\operatorname{Bernoulli}(\operatorname{expit}\{-0.2+0.5\,B+0.5\,C\}).$
 We specify that throughout the treatment assignment is influenced by $B$, $C$, $L_{k}$ and $W_{k}$ as:
@@
 $W_{k}$ $\sim\operatorname{Bernoulli}(\operatorname{expit}\{-0.2+0.5\,B+0.5\,C+W_{k-1}-0.6\,A_{k-1}\}).$
-We assume a Cox marginal structural model given by:
-$\lambda(t\,|\,do(\bar{A}_{\left\lfloor t\right\rfloor}),C)=\lambda_{0}\exp\left(\beta_{C}\,C+\beta_{A}\,A_{\left\lfloor t\right\rfloor}+\beta_{CA}\,C\,A_{\left\lfloor t\right\rfloor}\right),$ (25)
-where $\lambda_{0}$ is the baseline hazard and is assumed to be constant over time, $A_{\left\lfloor t\right\rfloor}$ represent the treatment status at the most recent visit prior to time $t$ (with $\left\lfloor t\right\rfloor$ denoting the largest integer less than $t$).
-We set $\lambda_{0}=\exp(-2)$ and $\beta_{C}=-0.5$.
-We vary the values of $(\beta_{A},\beta_{CA})$ to construct scenarios with different strengths of treatment: a strong effect scenario $(-0.5,0.3)$ and a weak effect scenario $(0.2,0.1)$.
-This parameterization assumes a constant hazard rate between follow-up visits, influenced by the treatment status at the most recent visit.
-This implies that given survival at follow-up time $k$, the survival time $T$ follows an exponential distribution with the rate equal to the hazard above:
-$T-k\,|\,T>k\sim\operatorname{Exp}(\lambda(t\,|\,do(\bar{A}_{\left\lfloor t\right\rfloor}),C)).$
-If $T-k<1$ then $Y_{k+1}=1$, and 00 otherwise.
-We set the total number of follow-ups $K$ to 6.
-For simplicity we set the copulas that weave the dependence between $L$, $W$ and $Y$ so those that describe the dependency between $L_{j}$ and $Y_{k+1}$ for $j=0,1,2,\dots,k$ are Student t with parameter $\rho_{L}$ and 5 degrees of freedom, and those for $W_{j}$ and $Y_{k+1}$ (given $L_{k}$) for $j=0,1,2,\dots,k$ are Gaussian with correlation $\rho_{W}$.
-We vary the values of $\rho_{L}$ and $\rho_{W}$ to construct scenarios with different strengths of confounding.
-Table 4 summarizes the different parameter values.
-In total, we consider 6 scenarios corresponding to the combination of different effect size and confounding scenarios, and three sample sizes: $n=\{1000,2000,5000\}$.
+Parameter Scenarios
+Table6 summarizes the parameter values for treatment effect size and confounding scenarios:
 Effect size Large $(\beta_{A},\beta_{CA})=(-0.5,0.3)$ Small $(\beta_{A},\beta_{CA})=(0.2,0.1)$ Level of confounding High $(\rho_{L},\rho_{W})=(0.9,0.7)$ Medium $(\rho_{L},\rho_{W})=(-0.5,0.4)$ Low $(\rho_{L},\rho_{W})=(-0.1,-0.2)$
 Table 6: Parameter values of treatment effect size and confounding scenarios.
-Remark 6.2.
-The proposed algorithm, as previously discussed, exclusively relies on evaluating analytic functions, making it highly efficient.
+Runtime details
 Table 7 reports the runtime required to simulate data under the setup for the large causal effect with medium confounding scenario.
@@
 Table 7: Runtime (in seconds) to simulate a dataset of size $n$, averaged over 20 repeats.
-We contrast the results of two approaches for fitting the Cox marginal structural model specified in (25).
-The first employs stabilized inverse probability of treatment weighting, with weights for subject $i$ at time step $k$ calculated as:
+Weight calculation
+To fit the Cox marginal model, we employ stabilized inverse probability of treatment weighting, with weights for subject $i$ at time step $k$ calculated as:
 $\prod_{j=0}^{k}\frac{\hat{p}_{j}(A_{j}\mid C,\bar{A}_{j-1},Y_{j}=0)}{\hat{p}_{j}(A_{j}\mid B,C,\bar{L}_{j},\bar{W}_{j},\bar{A}_{j-1},Y_{j}=0)},$
 where $\hat{p}_{j}(A_{j}\mid C,\bar{A}_{j-1},Y_{j}=0)$ and $\hat{p}_{j}(A_{j}\mid B,C,\bar{L}_{j},\bar{W}_{j},\bar{A}_{j-1},Y_{j}=0)$ are estimated using logistic regression models fitted to the survivors at time step $j$.
-Remark 6.3.
 The denominator in the weight formula represents the correctly specified propensity score model under the assumed data-generating process, while the numerator stabilizes the weights.
@@
 In accordance with the model in (15), we only include $C$ in the numerator for stabilization.
-The other method is to fit model (25) nominally, without weights or any adjustment for confounding.
-Table 6 summarizes the bias of estimates produced by inverse probability of treatment weighted and unweighted Cox regressions, from 500 replications.
-The biases are compared with their corresponding Monte Carlo standard errors as a measure of uncertainty.
-As the sample size increases, generally, the biases of estimates for $\beta_{C}$, $\beta_{A}$ and $\beta_{AC}$ obtained with inverse probability weights converge to zero, with their absolute values substantially smaller than the corresponding Monte Carlo standard deviations.
-In contrast, the unweighted estimates exhibit considerable bias, particularly for the estimate of the treatment effect, $\beta_{A}$, where the bias does not diminish as $n$ increases to $5000$.
-The magnitude of this bias far exceeds the Monte Carlo standard errors.
-The unweighted estimates are most biased under the high level of confounding by $L$ and $W$, since this is not adjusted for.
-Level of confounding Effect size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ IPTW Unweighted IPTW Unweighted IPTW Unweighted High $1000$ $-0.04(2.27)$ $0.27(3.09)$ $3.73(1.73)$ $-1.05(1.54)$ $0.49(2.22)$ $2.31(9.34)$ $\beta_{A}=-0.5$ $2000$ $-0.07(1.68)$ $0.02(2.09)$ $3.81(1.21)$ $-1.03(1.06)$ $0.34(1.51)$ $1.63(9.52)$ $5000$ $0.02(1.25)$ $-0.09(1.58)$ $3.89(0.76)$ $-0.96(0.68)$ $0.24(0.97)$ $1.08(9.44)$ $1000$ $-0.03(2.35)$ $0.03(2.78)$ $3.64(1.46)$ $-0.93(1.59)$ $0.44(1.94)$ $1.99(9.62)$ $\beta_{A}=0.2$ $2000$ $-0.11(1.60)$ $0.17(2.02)$ $3.65(1.18)$ $-0.93(1.11)$ $0.49(1.51)$ $1.49(9.46)$ $5000$ $-0.03(1.14)$ $0.07(1.36)$ $3.65(0.71)$ $-0.95(0.71)$ $0.51(0.89)$ $1.01(9.38)$ Medium $1000$ $-0.07(2.04)$ $0.23(2.91)$ $1.90(1.73)$ $-0.43(1.51)$ $0.08(2.22)$ $2.19(9.48)$ $\beta_{A}=-0.5$ $2000$ $-0.02(1.60)$ $0.09(2.27)$ $1.95(1.22)$ $-0.35(1.01)$ $-0.01(1.55)$ $1.65(9.52)$ $5000$ $-0.09(1.10)$ $0.07(1.59)$ $1.97(0.80)$ $-0.42(0.64)$ $-0.02(1.03)$ $1.21(9.32)$ $1000$ $-0.12(2.06)$ $0.14(2.70)$ $1.75(1.54)$ $-0.36(1.43)$ $0.09(1.92)$ $2.02(9.50)$ $\beta_{A}=0.2$ $2000$ $-0.10(1.62)$ $0.09(2.09)$ $1.68(1.11)$ $-0.42(1.08)$ $0.17(1.44)$ $1.49(9.36)$ $5000$ $-0.03(1.04)$ $0.03(1.32)$ $1.78(0.68)$ $-0.39(0.66)$ $0.14(0.88)$ $0.95(9.54)$ Low $1000$ $0.01(1.82)$ $0.21(2.90)$ $-1.95(1.83)$ $0.46(1.29)$ $-0.20(2.20)$ $2.36(9.50)$ $\beta_{A}=-0.5$ $2000$ $-0.10(1.32)$ $0.07(2.13)$ $-1.83(1.25)$ $0.44(0.89)$ $-0.30(1.59)$ $1.69(9.58)$ $5000$ $-0.02(1.04)$ $0.03(1.46)$ $-1.77(0.84)$ $0.47(0.58)$ $-0.32(0.99)$ $1.07(9.42)$ $1000$ $0.04(1.98)$ $-0.03(2.65)$ $-1.88(1.46)$ $0.53(1.30)$ $-0.31(1.88)$ $1.98(9.64)$ $\beta_{A}=0.2$ $2000$ $-0.02(1.45)$ $0.01(1.87)$ $-1.85(1.10)$ $0.43(0.99)$ $-0.26(1.38)$ $1.43(9.44)$ $5000$ $-0.01(0.92)$ $-0.06(1.25)$ $-1.89(0.68)$ $0.43(0.59)$ $-0.25(0.84)$ $0.95(9.50)$
-Table 6: Bias and Monte Carlo standard errors (×$\times 10$) of estimates produced by inverse probability of treatment weighted (‘IPTW’) and unweighted Cox regressions, under different strengths of confounding, effect sizes, and sample sizes.
-Table 7 presents the coverage rates of the 95% confidence intervals for both the inverse probability weighted and unweighted estimators.
-Consistent with the illustrative example in Section 5, the confidence intervals for the inverse probability weighted estimator were computed using two methods:one based on sandwich variance estimation and the other on bootstrap resampling.
-The results indicate that the confidence intervals for the inverse probability weighted estimator, obtained using both methods, achieve coverage rates close to the nominal 95% level.
-However, slight undercoverage is observed with both approaches, and the extent of undercoverage is comparable between them.
-In contrast, the sandwich-based confidence intervals for the unweighted estimator perform poorly, particularly for the treatment effect estimate, $\beta_{A}$.
-Level of confounding Effect Size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ W-S W-B UW W-S W-B UW W-S W-B UW High 1000 93.8% 95.2% 88.4% 93.4% 92.8% 42.2% 93.8% 93.8% 93.4% $\beta_{A}=-0.5$ 2000 94.0% 94.0% 84.4% 95.2% 95.6% 12.4% 95.6% 95.6% 94.6% 5000 93.2% 92.8% 70.2% 94.4% 93.4% 0.2% 94.2% 93.6% 93.8% 1000 92.8% 93.8% 91.8% 96.2% 95.8% 31.6% 93.4% 94.0% 95.6% $\beta_{A}=0.2$ 2000 95.2% 95.0% 87.6% 94.6% 95.0% 9.8% 94.8% 94.6% 92.2% 5000 93.6% 93.6% 72.4% 93.8% 93.0% 0.2% 94.0% 94.0% 90.0% Medium 1000 93.8% 94.2% 93.4% 94.8% 93.6% 79.6% 93.8% 94.0% 93.6% $\beta_{A}=-0.5$ 2000 93.0% 93.8% 92.4% 95.2% 94.6% 64.8% 92.2% 93.4% 96.2% 5000 95.4% 94.8% 90.0% 93.2% 92.6% 30.2% 93.0% 93.8% 93.6% 1000 94.2% 94.8% 95.0% 95.0% 93.8% 77.2% 94.2% 93.2% 95.8% $\beta_{A}=0.2$ 2000 94.6% 93.8% 91.6% 93.6% 93.2% 64.0% 92.8% 92.2% 94.2% 5000 95.4% 94.6% 91.6% 95.4% 95.2% 26.8% 94.8% 94.4% 95.4% Low 1000 95.4% 95.4% 93.0% 95.0% 95.2% 84.0% 95.2% 95.2% 94.6% $\beta_{A}=-0.5$ 2000 95.6% 96.4% 93.0% 95.8% 96.0% 71.6% 95.0% 95.6% 94.6% 5000 93.2% 93.2% 87.2% 94.2% 94.8% 42.6% 94.0% 94.4% 94.4% 1000 94.6% 95.6% 94.8% 96.4% 96.6% 77.4% 92.8% 94.8% 96.0% $\beta_{A}=0.2$ 2000 94.8% 95.6% 92.8% 94.4% 94.6% 60.6% 95.0% 95.2% 94.2% 5000 94.8% 94.2% 89.4% 95.0% 95.4% 20.0% 95.0% 94.2% 93.8%
-Table 7: Coverage rates for IPTW-Sandwich (W-S), IPTW-Bootstrap (W-B), and unweighted (UW) estimators of $\beta_{C}$, $\beta_{A}$, and $\beta_{AC}$ under different strengths of confounding, effect sizes, and sample sizes.
-In summary, the asymptotically unbiased nature of the inverse probability weighted estimates, compared to the biased unweighted estimates, further confirms that the simulated data were generated in accordance with the Cox model in (25), and that confounding by $L$ and $W$ was appropriately introduced.
-R code for implementing this study is provided in Appendix A.
-7 Discussion
-In this paper, we proposed and demonstrated a principled strategy for parametrically simulating longitudinal data from marginal structural models, leveraging the frugal parameterization (Evans and Didelez, 2024).
-Our proposed method is highly flexible and can accommodate a variety of practical scenarios, including multiple treatments (for example, primary and drop-in), dynamic treatment regimes, dose-response analyses, multiple outcomes (such as secondary outcomes and competing risks), and censoring.
-Additionally, it imposes no restrictions on the data-generating process or marginal structural causal models.
-In Section 6, we extended this method to survival models and time-to-event outcomes, which are common in clinical trials.
-Unlike the non-survival case, survival data simulation requires conditioning each time step $k+1$ on surviving through $k$, meaning all parameterized quantities must account for survival.
-This posed a challenge: while we can specify the distribution of the time-varying variable, say $L_{j}$, conditional on all prior history up to $j-1$, its distribution conditioning on a future survival state at time $k$ ($k>j$) is generally intractable.
-A very similar issue occurs in the approach proposed by Seaman and Keogh (2024) where Monte Carlo approximations are used to address the problem.
-While effective in theory, their approach is computationally expensive and imposes restrictions on the types of data that can be generated.
-We resolved this challenge elegantly by leveraging pair-copulas, so that these conditional quantiles can be iteratively computed through h-functions.
-Thus, our solution relies solely on evaluating analytic functions, making it highly efficient.
-This efficiency is crucial for practical applications, such as sample size calculation, which require repeated simulations across a wide range of parameter combinations.
-[…]
-If enforcing such Markov conditions is absolutely necessary, an alternative pair-copula construction can be adopted.
-However, such an approach requires computing marginal quantiles of time-varying confounders through numerical integration, which quickly becomes high-dimensional—a well-known computational challenge.
-Appendix A Example R code for Section 6.2
-Here is an example code to simulate data following the setup in Section 6.2.
-1library(data.table)
-2library(copula)
-3
-4sim_func <- function(n = 500,
-5 rho1 = -0.5, rho2 = 2,
-6 K = 5, betaS = -0.5,
-7 betaB2 = -0.5, betaSB2 = 0.3){
-8 # copulas
-9 copL <- tCopula(param = rho1, df = 5, dim = 2, dispstr = "un")
-10 copW <- normalCopula(param = rho2, dim = 2, dispstr = "un")
-11 dat <- data.table()
-12 dat\$surv <- rep(1,n)
-13
-14 # baseline vars
-15 dat[, B1 := rexp(n, rate = 2)]
-16 dat[, B2 := rbinom(n, size = 1, prob = 0.7)]
-17 baseline_vars <- c("B1","B2")
-18
-19 # a data table to store calculated quantiles to be carried over to the next time step
-20 comp_dat <- copy(dat[,baseline_vars, with = F])
-21
-22 for (k in 0:K){
-23 n_surv <- sum(dat\$surv, na.rm = T)
-24 if (k>0){
-25 comp_dat <- comp_dat[Y == 0,]
-26 if (k > 1){
-27 comp_dat[,grep("_prev\$", colnames(comp_dat), value = TRUE):= NULL]
-28 }
-29 # reset names
-30 setnames(comp_dat,
-31 old = setdiff(names(comp_dat), baseline_vars),
-32 new = paste0(setdiff(names(comp_dat), baseline_vars), "_prev"))
-33 }
-34
-35 # random quantiles
-36 comp_dat[, paste0("F_L",k,"|surv"):= runif(n_surv)]
-37 comp_dat[, paste0("F_W",k,"|L",k,"_surv"):= runif(n_surv)]
-38 comp_dat[, paste0("F_S",k,"|L",k,"_surv"):= runif(n_surv)]
-39 comp_dat[, paste0("F_Y|LW",k,"_surv"):= runif(n_surv)]
-40
-41 if (k == 0){
-42 comp_dat[,L := qgamma(get(paste0("F_L",k,"|surv")), shape = 1 + 0.5*B1 + 0.5*B2, scale = 1)]
-43 comp_dat[,W := qbinom(get(paste0("F_W",k,"|L",k,"_surv")), size = 1, prob = expit(-0.2 +0.5 * B1 + 0.5*B2))]
-44 } else{
-45 comp_dat[,L := qgamma(get(paste0("F_L",k,"|surv")), shape = 1 + 0.5*B1 + 0.5*B2 + 0.1 * L_prev - 0.5 * S_prev, scale = 1)]
-46 comp_dat[,W := qbinom(get(paste0("F_W",k,"|L",k,"_surv")), size = 1, prob = expit(-0.2 + 0.5 * B1 + 0.5*B2 + W_prev - 0.6* S_prev))]
-47 }
-48
-49 comp_dat[,S := qbinom(get(paste0("F_S",k,"|L",k,"_surv")),size = 1, prob = expit(-1.5 + 0.5*B1 + 0.5*B2 + 0.5* L + 0.5 * W))]
-50
-51 if (k >0){
-52 #get the distributions of Ls and Ws in the survivors
-53 for (j in (1:k)){
-54 if (j<k){
-55 comp_dat[,paste0("F_W",k-j,"|L",k-j,"_surv"):= cCopula(as.matrix(cbind(get(paste0("F_Y|L",k-j,"_W", k-j-1,"_surv_prev")),
-56 get(paste0("F_W",k-j,"|L",k-j,"_surv_prev")))),copula = copW)[,2]]
-57 comp_dat[,paste0("F_L",k-j,"|surv"):= cCopula(as.matrix(cbind(get(paste0("F_Y|LW",k-j-1,"_surv_prev")),
-58 get(paste0("F_L",k-j,"|surv_prev")))),copula = copL)[,2]]
-59 } else{
-60 comp_dat[,paste0("F_W",k-j,"|L",k-j,"_surv"):= cCopula(as.matrix(cbind(get(paste0("F_Y|L",k-j,"_surv_prev")),
-61 get(paste0("F_W",k-j,"|L",k-j,"_surv_prev")))),copula = copW)[,2]]
-62 comp_dat[,paste0("F_L",k-j,"|surv"):= cCopula(as.matrix(cbind(F_Y_surv_prev,
-63 get(paste0("F_L",k-j,"|surv_prev")))),copula = copL)[,2]]
-64 }
-65 }
-66 }
-67
-68 for (j in (0:k)){
-69 if (j<k){
-70 comp_dat[,paste0("F_Y|L",k-j,"_W", k-j-1,"_surv"):= cCopula(as.matrix(cbind(get(paste0("F_W",k-j,"|L",k-j,"_surv")),
-71 get(paste0("F_Y|LW",k-j,"_surv")))),copula = copW,inverse = T)[,2]]
-72 comp_dat[,paste0("F_Y|LW",k-j-1,"_surv"):= cCopula(as.matrix(cbind(get(paste0("F_L",k-j,"|surv")),
-73 get(paste0("F_Y|L",k-j,"_W", k-j-1,"_surv")))),copula = copL,inverse = T)[,2]]
-74 }else{
-75 comp_dat[,paste0("F_Y|L",k-j,"_surv"):= cCopula(as.matrix(cbind(get(paste0("F_W",k-j,"|L",k-j,"_surv")),
-76 get(paste0("F_Y|LW",k-j,"_surv")))),copula = copW,inverse = T)[,2]]
-77 comp_dat[,F_Y_surv:= cCopula(as.matrix(cbind(get(paste0("F_L",k-j,"|surv")),
-78 get(paste0("F_Y|L",k-j,"_surv")))),copula = copL,inverse = T)[,2]]
-79 }
-80 }
-81
-82 # hazard
-83 comp_dat[, lambda := exp(-2) *exp(betaS * S + betaB2 * B2 + betaSB2*B2*S)]
-84 comp_dat[, T := qexp(F_Y_surv, rate = lambda)]
-85 comp_dat[, Y := 1*(T<1)]
-86
-87 dat[surv == 1, paste0("L","_",k) := comp_dat\$L]
-88 dat[surv == 1, paste0("W","_",k) := comp_dat\$W]
-89 dat[surv == 1, paste0("S","_",k) := comp_dat\$S]
-90 dat[surv == 1, paste0("Y","_",k) := comp_dat\$Y]
-91 dat[surv == 1, T := (comp_dat\$T + k)]
-92 dat[,surv := 1*(get(paste0("Y_",k)) == 0)]
-93 return(dat)
-94 }

### v3

@@
 Abstract
-Simulating longitudinal data from specified marginal structural models is a crucial but challenging task for evaluating causal inference methods and clinical trial design.
-While data generation typically proceeds in a fully conditional manner using structural equations according to a temporal ordering, it is difficult to ensure alignment between conditional distributions and the target marginal causal effects.
-This misalignment presents a fundamental challenge in simulating data that adheres to marginal structural model specifications.
+Simulating longitudinal data from specified marginal structural models is a crucial but challenging task for evaluating causal inference methods and informing study design.
+While data generation typically proceeds in a fully conditional manner using structural equations according to a temporal ordering, it is difficult to ensure alignment between conditional distributions and the target marginal causal effects, which presents a fundamental challenge.
 To address this, we propose a flexible and efficient algorithm for simulating longitudinal data that adheres exactly to a specified marginal structural model.
-Recognizing the importance of time-to-event outcomes in clinical research, we extend our approach to accommodate survival models.
-Compared to existing approaches, our method offers several advantages: it enables exact simulation from a known causal model rather than relying on approximations; avoids restrictive assumptions about the data-generating process; and remains computationally efficient by requiring only the evaluation of analytic functions.
-This last benefit contrasts with methods that use computationally intensive techniques such as Monte Carlo approximations or numerical integration.
+Our approach accommodates time-to-event outcomes and extends naturally to survival settings, which are prevalent in applied research.
+Compared to existing approaches, it offers several advantages: it enables exact simulation from a known causal model rather than relying on approximations; avoids restrictive assumptions about the data-generating process; and remains computationally efficient by requiring only the evaluation of analytical expressions, rather than Monte Carlo methods or numerical integration.
 Through simulation studies replicating realistic scenarios, we validate the method’s accuracy and utility.
@@
 1 Introduction
-Longitudinal data are fundamental for studies analyzing the causal effect of a treatment or exposure on an outcome, especially for research into chronic diseases such as diabetes or HIV.
+Longitudinal data are fundamental for studies analyzing the causal effect of a treatment or exposure, especially for research into chronic diseases such as diabetes or HIV.
 Unlike cross-sectional studies, longitudinal designs capture the dynamic nature of treatment regimens, which often vary over time due to clinical decisions, patient adherence, or external factors.
@@
 A central challenge in analyzing longitudinal data is addressing time-varying confounding, where a time-dependent covariate predicts both the outcome and future treatments, while itself being influenced by past treatments.
-For example, in the case study in Robins et al.
+For example, in the case study by Robins et al.
 (2000) on the effect of zidovudine (AZT) treatment on HIV infection, the biomarker CD4 lymphocyte count is both affected by prior AZT treatment and is predictive of future therapy initiation and HIV progression.
@@
 Model parameters can be consistently estimated using inverse probability of treatment weighting (Robins, 1999, 2000).
-Among alternative methods, including structural nested models estimated through g-computation (Robins, 1989; Robins and Tsiatis, 1991), marginal structural models have gained prominence due to their interpretability and ease of implementation.
+Among alternative methods, including structural nested models estimated through g-estimation (Robins, 1989, 1992; Robins and Tsiatis, 1991), marginal structural models have gained prominence due to their interpretability and ease of implementation.
 A systematic review by Clare et al.
@@
 Sample size calculations also benefit from simulations.
-Prior to conducting a trial or an observational study, it is crucial to determine the number of participants or number of observational units required to achieve adequate statistical power.
-While standard formulas can be used for time-fixed treatments, more complex scenarios with time-varying treatments make deriving analytical formulas impractical or impossible.
+Prior to conducting a trial or an observational study, it is crucial to determine the number of participants or observed units required to achieve adequate statistical power.
+Although standard formulas can sometimes be used for time-fixed treatments, more complex scenarios with time-varying treatments make deriving analytical formulas impractical or impossible.
 In such cases, simulation is invaluable, allowing researchers to estimate sample sizes based on realistic data-generating processes and causal structures.
@@
 While generating data sequentially from structural equations—creating covariates, treatment status, and then the outcome conditioned on past information—may seem intuitive, these models rely on conditional rather than marginal specifications.
-This makes it difficult to sequentially simulate with conditional distributions in a way that aligns with a pre-specified marginal model; in particular, it is extremely challenging to sample from a null hypothesis —the so-called g-null paradox (Robins and Wasserman, 1997).
+This makes it difficult to sequentially simulate from conditional distributions that align with a pre-specified marginal model; in particular, it is extremely challenging to sample from a null hypothesis—the so-called g-null paradox (Robins and Wasserman, 1997).
 The remainder of this paper is structured as follows: Section 2 reviews existing methods for simulating longitudinal data from causal models.
 Section 3 introduces key preliminaries, including the frugal parameterization (Evans and Didelez, 2024) and copulas.
-In Section 4, we introduce our proposed methodology and illustrate it in Section 4.4 through a simulation study replicating a real-world study in epidemiology (Bodnar et al., 2004).
-Section 5 extends the simulation approach to time-to-event data, enabling simulations from survival models such as Cox marginal structural models, including an illustrative example.
-The paper concludes with a discussion in Section 6.
-Example code for the simulations in Sections 4.4 and 5.3 is available at https://github.com/XiLinStats/LongSiMSM.
+In Section 4, we introduce our proposed methodology.
+Section 5 extends the approach to time-to-event data, enabling simulations from survival models such as Cox marginal structural models, and Section 5.2 illustrates the method through a simulation study replicating a real-world clinical scenario involving concomitant treatment in a trial.
+We conclude with a discussion in Section 6.
+Example code for the simulations in Sections 4.4 and 5.2 is available at https://github.com/XiLinStats/LongSiMSM.
 2 Related Literature
@@
   1. (i)
-    Simulating from conditional models and deriving or approximating the corresponding marginal models post-hoc.
+    Simulating from conditional models and deriving or approximating the corresponding marginal models post hoc.
   2. (ii)
@@
 Xiao et al.
-(2010) used sequential conditional models to approximate Cox marginal structural models under the rare disease assumption, i.e.
-very low failure rates, which mitigates the impact of noncollapsibility but assumes the absence of unmeasured confounding between time-varying confounders and outcomes.
+(2010) used sequential conditional models to approximate Cox marginal structural models under the rare disease assumption, i.e., very low failure rates, which mitigates the impact of non-collapsibility but assumes the absence of unmeasured confounding between time-varying confounders and outcomes.
 Young and Tchetgen Tchetgen (2014) employed the g-formula to bridge conditional and marginal models in discrete-time Cox marginal structural models.
@@
 Havercroft and Didelez (2012) simulated directly from a discrete Cox marginal structural model, introducing time-varying confounding indirectly via a shared latent variable.
-However, this approach assumes no direct effect of time-varying confounders on outcomes, which may not hold in practice, e.g.
-the effect of CD4 count on HIV survival (Robins et al., 2000).
+However, this approach assumes no direct effect of time-varying confounders on outcomes, which may not hold in practice, e.g., the effect of CD4 count on HIV survival (Robins et al., 2000).
 Evans and Didelez (2024) proposed a frugal parameterization that prioritizes the causal margin and parameterizes the surrounding variables accordingly, enabling direct simulation from the exact marginal model while explicitly modeling time-varying covariates.
@@
 Building on this, Seaman and Keogh (2024) applied the frugal parameterization to simulate survival outcomes, using a one-dimensional “risk score” to summarize multi-dimensional covariates.
-While simplifying dependence structures, this method relies on computationally expensive Monte Carlo approximations to estimate the cumulative distribution functions of the risk score, making it infeasible for continuous treatments or baseline variables.
-The aims for a new simulation approach therefore include the ability to: precisely specify the marginal structural model; include multiple time-varying covariates, treatments, outcomes, and censoring; have good control over their distributions; avoid restrictive assumptions about the data-generating process.
-In addition, we would like it to scale well in the number of time-points and variables, and be applicable to real-world problems.
+While simplifying dependence structures, this method relies on computationally expensive Monte Carlo approximations to estimate the cumulative distribution functions of the risk score.
+In settings with continuous treatments or baseline variables, these approximations must be performed separately for each individual at each time point, substantially increasing the computational burden.
+The aims for a new simulation approach therefore include the ability to: simulate exactly from a specified marginal structural model; include multiple time-varying covariates, treatments, outcomes, and censoring; have good control over their distributions; avoid restrictive assumptions about the data-generating process.
+In addition, we would like it to scale well in the number of time-points and variables, and applicable in real-world settings.
 3 Preliminaries
@@
 This structure can be represented by a copula, which is a multivariate distribution function $C:[0,1]^{d}\rightarrow[0,1]$ with margins uniformly distributed on $(0,1)$.
-According to Sklar’s theorem (Sklar, 1959, 1973), any continuous multivariate distribution $F$ with marginals $F_{1},\ldots,F_{d}$ can be expressed as
+According to Sklar’s theorem (Sklar, 1959, 1973), let $F$ be a multivariate distribution with margins $F_{1},\ldots,F_{d}$.
+Then there exists a copula $C$ such that for all $(x_{1},\ldots x_{d})$ in $\mathbb{R}^{d}$,
 $F(x_{1},\ldots,x_{d})=C\left\{F_{1}(x_{1}),\ldots,F_{d}(x_{d})\right\}=C(u_{1},\ldots,u_{d}),$
-where $\bm{X}=(X_{1},\ldots,X_{d})$ is a vector of random variables, and $u_{i}=F_{i}(x_{i})$, $i=1,\ldots,d$. The copula then takes the form
+where $u_{i}=F_{i}(x_{i})$, $i=1,\ldots,d$. If $F_{1}(x_{1}),\ldots,F_{d}(x_{d})$ are all continuous, then $C$ is unique and takes the form
 $C(u_{1},\ldots,u_{d})=F\{F_{1}^{-1}(u_{1}),\ldots,F_{d}^{-1}(u_{d})\},\qquad(u_{1},\ldots,u_{d})\in[0,1]^{d}.$
+Otherwise, $C$ is only uniquely determined on $\operatorname{Ran}F_{1}\times\cdots\times\operatorname{Ran}F_{d}$, where $\operatorname{Ran}F_{i}$ denotes the range of $F_{i}$.
 The choice of copula can be considered independently of the univariate margins.
@@
 This hierarchical structure incorporates additional variables at each level, using pair-copulas as the fundamental building blocks.
-Following the general formula given in Joe (1997), the conditional distribution of one variable $y$ given $d$ other variables can be decomposed to $d$ bivariate copula density functions, each of which can be specified flexibly:
+Following the general formula given in Joe (1997), and assuming that the joint distribution of $(Y,X_{1},\ldots,X_{d})$ is absolutely continuous with continuous margins, the conditional distribution of $Y$ given $d$ other variables can be decomposed into $d$ bivariate copula density functions, each of which can be specified flexibly:
 $f(y\,|\,\bm{x})=c_{YX_{j}|\bm{X}_{-j}}\left\{F(y\,|\,\bm{x}_{-j}),F(x_{j}\,|\,\bm{x}_{-j})\right\}\cdot f(y\,|\,\bm{x}_{-j}),$ (1)
@@
 3.2.2 Simulating from pair-copula constructions
-The pair-copula construction includes conditional distributions of the form $F(y\,|\,\bm{x})$. Joe (1996) demonstrated the following result:
-$F(y\,|\,\bm{x})=\frac{\partial C_{YX_{j}\mid\bm{X}_{-j}}\left\{F(y\,|\,\bm{x}_{-j}),F(x_{j}\,|\,\bm{x}_{-j})\right\}}{\partial F(x_{j}\,|\,\bm{x}_{-j})}.$ (3)
+The pair-copula construction includes conditional distributions of the form $F(y\,|\,\bm{x})$.
+Assuming that the joint distribution of $(Y,X_{1},\ldots,X_{d})$ is absolutely continuous with continuous margins, and the bivariate copula $C_{YX_{j}\mid\bm{X}_{-j}}$ is differentiable, Joe (1996) demonstrated the following result:
+$F(y\,|\,\bm{x})=\frac{\partial C_{YX_{j}\mid\bm{X}_{-j}}\left\{F(y\,|\,\bm{x}_{-j}),F(x_{j}\,|\,\bm{x}_{-j})\right\}}{\partial F(x_{j}\,|\,\bm{x}_{-j})},$ (3)
+where $j\in\{1,\ldots,d\}$.
 For notational convenience, we introduce the h-function $h(u_{1},u_{2};\Theta)$ from Aas et al.
 (2009).
-Let $u_{1}$ and $u_{2}$ be random variables with a uniform distribution on $(0,1)$.
-The $h$-function is defined as
+Let $u_{1}$ and $u_{2}$ be random variables with a uniform distribution on $(0,1)$, for a sufficiently regular copula $C_{12}(u_{1},u_{2})$, the $h$-function is defined as
 $h(u_{1},u_{2};\Theta):=F(u_{1}\,|\,u_{2})=\frac{\partial}{\partial u_{2}}C_{12}(u_{1},u_{2};\Theta)$
@@
 The $h$-function is effectively the conditional copula $C(u_{1}\,|\,u_{2};\Theta)$.
+We further define $h^{-1}(u_{1},u_{2};\Theta)$, be the inverse of the $h$-function in the first variable $u_{1}$.
+Equivalently, $h^{-1}(u_{1},u_{2};\Theta):=F^{-1}(u_{1}\,|\,u_{2})$.
 Remark 3.1.
-The derivation of the analytical form of $h(\cdot)$ for common bivariate copulas is straightforward by definition for parametric families.
-Examples include Gaussian, Student t, Clayton, Frank, and Gumbel copulas.
+Here we define a copula $C$ as sufficiently regular if it has non-zero density everywhere on its domain.
+Under this condition, the associated $h$-functions exist and are invertible.
+For standard parametric families—including the Gaussian, Student-t, Clayton and Frank copulas—the analytical form of $h(\cdot)$ and its inverse $h^{-1}(\cdot)$ can be derived directly from their definitions.
+See Appendix C of Aas et al.
+(2009).
 The $h$-functions are foundational for simulating from $F(y\,|\,\bm{x})$.
@@
 At each follow-up $k$, treatment status $A_{k}$, and a set of time-varying covariates $L_{k}$ are recorded.
-Additionally, $Z$ represents a set of static baseline covariatesatent variables, if present, are written as $\{H_{k}\}_{k\in\mathcal{K}}$.
+Additionally, $Z$ represents a set of static baseline covariates.
+Latent variables, if present, are written as $\{H_{k}\}_{k\in\mathcal{K}}$.
 To indicate the history of a variable up to time $k$, we use bar notation.
@@
 Following the simulation strategy from pair-copula constructions outlined in Section 3.2.2, the fully-conditional distribution of the outcome is alternatively parameterized as:
-$F(y\,|\,\bar{l}_{K},\bar{a}_{K})$ $=h\big{[}\dots h\{h(F^{*}(y\,|\,do(\bar{a}_{K})),F(l_{0});\Theta_{0}),F(l_{1}\,|\,l_{0},a_{0});\Theta_{1}\}\dots;\Theta_{K}\big{]},$
-which is centered around the causal margin $F^{*}(y\,|\,\bar{a}_{K})$.
-Through inversion the outcome can be simulated in two steps:
+$F(y\,|\,\bar{l}_{K},\bar{a}_{K})$ $=h\big{[}\dots h\big{\{}h(F_{Y\mid\bar{A}_{K}}^{*}(y\,|\,do(\bar{a}_{K})),\,F(l_{0});\Theta_{0}),\,F(l_{1}\,|\,l_{0},a_{0});\Theta_{1}\big{\}}\dots;\Theta_{K}\big{]},$ (8)
+which is centered around the causal margin $F^{*}_{Y\mid\bar{A}_{K}}$.
+We write $h_{k}(u_{1},u_{2})=h(u_{1},u_{2};\Theta_{k})$, so as to avoid including the parameters $\Theta_{k}$.
+By the inversion method, it is straightforward that given ($\bar{l}_{K},\bar{a}_{K}$) and $U\sim\operatorname{U}(0,1)$,
+$Y$ $=F^{*-1}_{Y\mid\bar{A}_{K}}(h_{0}^{-1}\big{[}\dots h_{K-1}^{-1}\big{\{}h_{K}^{-1}(U,F(l_{K}\,|\,\bar{l}_{K-1},\bar{a}_{K-1})),F(l_{K-1}\,|\,\bar{l}_{K-2},\bar{a}_{K-2})\big{\}},\ldots,F(l_{0})\big{]})$
+follows the distribution parameterized by (7).
+This result means that following the approach outline in Section 4.1, given the simulated baseline variables and time-varying variables, we can simulate the outcome in two steps:
   1. 1.
     Calculate the marginal quantile of the outcome, $u^{*}_{Y}$, as:
-    $u^{*}_{Y}$ $=h^{-1}\big{[}\dots h^{-1}\big{\{}h^{-1}(w_{Y},F(l_{K}\,|\,\bar{l}_{K-1},\bar{a}_{K-1});\Theta_{K}),$
-    $\qquad\qquad\qquad F(l_{K-1}\,|\,\bar{l}_{K-2},\bar{a}_{K-2});\Theta_{K-1}\big{\}}\dots,F(l_{0});\Theta_{0}\big{]}$ (8)
-    $=h^{-1}\big{[}\dots h^{-1}\big{\{}h^{-1}(w_{y},w_{K,1};\Theta_{K}),w_{K-1,1};\Theta_{K-1}\big{\}},\ldots,w_{0,1};\Theta_{0}\big{]},$
+    $u^{*}_{Y}$ $=h_{0}^{-1}\big{[}\dots h_{K-1}^{-1}\big{\{}h_{K}^{-1}(w_{Y},F(l_{K}\,|\,\bar{l}_{K-1},\bar{a}_{K-1})),F(l_{K-1}\,|\,\bar{l}_{K-2},\bar{a}_{K-2})\big{\}}\dots,F(l_{0})\big{]}$ (9)
+    $=h_{0}^{-1}\big{[}\dots h_{K-1}^{-1}\big{\{}h_{K}^{-1}(w_{Y},w_{K,1}),w_{K-1,1}\big{\}},\ldots,w_{0,1}\big{]},$
     where $w_{Y}\sim\operatorname{U}(0,1)$.
@@
     Obtain $y$ through the inverse cumulative distribution function of the marginal distribution: $y=F^{*-1}_{Y\mid\bar{A}_{K}}(u^{*}_{Y}\,|\,do(\bar{a}_{K}))$.
-An intuitive way to understand the calculation in Step 1 is that the inverse $h$-functions “disentangle” the dependencies between $Y$ and each of $L_{K},L_{K-1},\dots L_{0}$ sequentially, and remove the conditioning of $Y$ on $\bar{L}_{K}$ until the marginal quantile of $Y$ is obtained.
+An intuitive way to understand the calculation in Step 1 is that the inverse $h$-functions “unravel” the dependencies between $Y$ and each of $L_{K},L_{K-1},\dots L_{0}$ sequentially, and remove the conditioning of $Y$ on $\bar{L}_{K}$ until the marginal quantile of $Y$ is obtained.
 The quantiles in the inverse $h$-functions, such as $F(l_{k}\,|\,\bar{l}_{k-1},\bar{a}_{k-1})$, correspond exactly to the random standard uniform variables $w_{k,1}$ used to simulate $l_{k}$ at time step $k$.
@@
 end while
-Sample $w_{y}\sim U(0,1)$
+Sample $w_{y}\sim\operatorname{U}(0,1)$
 Let $\nu_{0}=w_{y}$
 for $j=0$ to $K$ do
-     $\nu_{j+1}\leftarrow h^{-1}(\nu_{j},w_{(K-j),1};\Theta_{K-j})$
+     $\nu_{j+1}\leftarrow h_{K-j}^{-1}(\nu_{j},w_{(K-j),1})$
 end for
@@
 return $(z,\bar{l}_{K},\bar{a}_{K},y)$
-Extending the simulation algorithm to include a latent process $\{H_{k}\}_{k\in\mathcal{K}}$, or multiple time-varying confounders is straightforward.
+Extending the simulation algorithm to incorporate multiple time-varying confounders—whether assumed to be observed or unobserved — is straightforward.
 The key modification lies in the outcome simulation: the marginal quantile of $Y$ is obtained by sequentially removing the conditioning on all the time-dependent confounders at each step, through inverse $h$-functions.
-We provide an illustrative example including a latent variable $H_{k}$ in Appendix A.
+We provide an illustrative example in Appendix A, outlining the simulation procedure for incorporating an additional latent process $\{H_{k}\}_{k\in\mathcal{K}}$.
 4.4 Illustrative example
@@
   • Discrete-time marginal structural model, where the binary survival outcome $Y_{k}$ is modeled directly:
-    $P(Y_{k+1}=1\,|\,do(\bar{A}_{k})=\bar{a}_{k},Z=z,Y_{k}=0)=f(\bar{a}_{k},z;\alpha),$ (9)
+    $P(Y_{k+1}=1\,|\,do(\bar{A}_{k}=\bar{a}_{k}),Z=z,Y_{k}=0)=f(\bar{a}_{k},z;\alpha),$ (10)
     where $f$ can be any real-valued functions bounded between 0 and 1, for example, an expit function for a logistic marginal structural model.
-  • Continuous-time hazard MSM, where we introduce an auxiliary continuous variable, $\tilde{Y}_{k+1}=T-k\,|\,T>k$, representing the incremental failure time given survival at the start of the $k$th interval.
-We model incremental survival time $\tilde{Y}_{k+1}$ and truncated it to obtain the observed binary survival outcome $Y_{k+1}$.
-Given survival at $k$, the time to failure is $T\,|\,T>k=k+\tilde{Y}_{k+1}$.
-Examples of marginal models for $\tilde{Y}_{k+1}$ include:
-    • Cox proportional hazard: $\lambda(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),z,y_{k}=0)=\lambda(\tilde{y}_{k+1}\,|\,do(\bar{0}_{k}),z,y_{k}=0)\exp\{g(\bar{a}_{k},z;\beta)\}$,
-    • Additive hazard: $\lambda(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),z,y_{k}=0)=\lambda(\tilde{y}_{k+1}\,|\,do(\bar{0}_{k}),z,y_{k}=0)+g(\bar{a}_{k},z;\gamma)$.
+  • Continuous-time hazard MSM, where we introduce an auxiliary continuous variable, $\widetilde{Y}_{k+1}=T-k\,|\,T>k$, representing the incremental failure time given survival at the start of the $k$th interval.
+We model incremental survival time $\widetilde{Y}_{k+1}$ and truncated it to obtain the observed binary survival outcome $Y_{k+1}$.
+Given survival at $k$, the time to failure is $T\,|\,T>k=k+\widetilde{Y}_{k+1}$.
+Examples of marginal models for $\widetilde{Y}_{k+1}$ include:
+    • Cox proportional hazard:
+      $\lambda(\widetilde{y}_{k+1}\,|\,do(\bar{a}_{k}),z,y_{k}=0)=\lambda(\widetilde{y}_{k+1}\,|\,do(\bar{0}_{k}),z,y_{k}=0)\exp\{g(\bar{a}_{k},z;\beta)\},$
+    • Additive hazard:
+      $\lambda(\widetilde{y}_{k+1}\,|\,do(\bar{a}_{k}),z,y_{k}=0)=\lambda(\widetilde{y}_{k+1}\,|\,do(\bar{0}_{k}),z,y_{k}=0)+g(\bar{a}_{k},z;\gamma).$
 We focus on describing the simulation algorithm for time-to-event outcomes, with a remark on adapting the approach for discrete survival outcomes, which follows a similar but simpler framework.
@@
 Unlike the derivation of outcome simulation for the non-survival case in Section 4.2, all quantities involved in the copulas must now be conditioned on survival up to this point.
-Specifically, when simulating the incremental failure time, $\tilde{y}_{k+1}$, given $Y_{k}=0$, (7) becomes:
-$p(\tilde{y}_{k+1}$ $\,|\,\bar{l}_{k},\bar{a}_{k},y_{k}=0)$
-$=\prod_{j=0}^{k}c^{*}_{\tilde{Y}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},Y_{k}=0}\big{\{}F(\tilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0),$
+Specifically, when simulating the incremental failure time, $\widetilde{y}_{k+1}$, given $Y_{k}=0$, (7) becomes:
+$p(\widetilde{y}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},y_{k}=0)=$ $p^{*}(\widetilde{y}_{k+1}\,|\,do(\bar{a}_{k}),y_{k}=0)\cdot$
+$\prod_{j=0}^{k}c^{*}_{\widetilde{Y}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},Y_{k}=0}\big{\{}F(\widetilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)),$
 $\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0);\Theta_{k,j}\}.$ (11)
@@
 To reiterate, our goal is to simulate the outcome at the end of the $k$th interval, which corresponds to the $k$th panel in Figure 2.
-Each copula in (10) associates $\tilde{y}_{k+1}$ with all the prior values of the time-varying confounder $L_{j}$, for $j=0,\ldots,k$.
+Each copula in (11) associates $\widetilde{y}_{k+1}$ with all the prior values of the time-varying confounder $L_{j}$, for $j=0,\ldots,k$.
 In the non-survival case discussed in Section 4.2, the quantiles $F(l_{j}\,|\,\bar{l}_{j-1},\bar{a}_{j-1})$, for $j=0,\ldots,k$, are directly available at each simulation at step $j$ and can be simply plugged into the iterative application of inverse $h$-functions without further computation.
@@
 This means that the conditional quantiles of $L_{j}$ that we need are not directly available and require additional computation.
-Fortunately, employing copulas to define the dependence structure offers an efficient method to derive these conditional distributions, eliminating the need for resource-intensive techniques such as Monte Carlo approximation or numerical integration.
-The proposed simulation algorithm follows an iterative structure and is formally presented and validated using an inductive approach.
+Fortunately, employing copulas to define the dependence structure offers a solution to deriving these conditional distributions.
+Lemma 5.1.
+In the setting described above, suppose $Y_{k}=0\Leftrightarrow\widetilde{Y}_{k}>1$, then the conditional distribution $F_{L_{j}\,|\,\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0}$ is given by
+$F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$ $=\frac{v_{k-1,j}-C_{\widetilde{Y}_{k}L_{j}\mid\bar{L}_{j},do(\bar{A}_{k}),Y_{k-1}=0}(v_{k-1,j},q_{k-1,j};\Theta_{k-1,j})}{1-q_{k-1,j}},$ (12)
+where $v_{k-1,j}=F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0)$ and $q_{k-1,j}=F(\widetilde{y}_{k}=1\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0)$.
+The proof is given in Appendix B.
+Here we emphasize that $F_{L_{j}\,|\,\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0}$ is the “renormalized” distribution in the population who have survived up to time-point $k$.
+This is distinct from simply conditioning on an individual’s previous incremental failure time, which would give $F_{L_{j}\,|\,\bar{L}_{j-1},do(\bar{A}_{k}),\widetilde{Y}_{k}}$.
+The former corresponds to the expectation of the latter with respect to $\widetilde{Y}_{k}$ given $\widetilde{Y}_{k}>1$.
+The proposed simulation algorithm combines an iterative and recursive structure and is formally presented and validated using an inductive approach.
 Proposed simulation method: At time step $k$, we specify the cumulative distribution functions:
-$F_{L_{k}\mid\bar{L}_{k-1},\bar{A}_{k-1},Y_{k}=0}$ $F_{A_{k}\mid\bar{L}_{k},\bar{A}_{k-1},Y_{k}=0}$ $F_{\tilde{Y}_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}.$
-Additionally, we define $k+1$ bivariate copulas $C_{\tilde{Y}_{k+1}L_{j}\mid\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0}$, parameterized by $\Theta_{k,j}$ for $j=0,\ldots,k$.
-These copulas captures the dependence between $\tilde{Y}_{k+1}$ and each preceding time-varying covariate.
+$F_{L_{k}\mid\bar{L}_{k-1},\bar{A}_{k-1},Y_{k}=0}$ $F_{A_{k}\mid\bar{L}_{k},\bar{A}_{k-1},Y_{k}=0}$ $F_{\widetilde{Y}_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}.$
+Additionally, we define $k+1$ bivariate copulas $C_{\widetilde{Y}_{k+1}L_{j}\mid\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0}$, parameterized by $\Theta_{k,j}$ for $j=0,\ldots,k$.
+These copulas captures the dependence between $\widetilde{Y}_{k+1}$ and each preceding time-varying covariate.
 Draw $w_{k,1}$, $w_{k,2}$, $w_{k,3}\sim\operatorname{U}(0,1)$ and compute:
 $l_{k}=F^{-1}_{L_{k}\mid\bar{L}_{k-1},\bar{A}_{k-1},Y_{k}=0}(w_{k,1}\,|\,\bar{l}_{k-1},\bar{a}_{k-1})$ $a_{k}=F^{-1}_{A_{k}\mid\bar{L}_{k},\bar{A}_{k-1},Y_{k}=0}(w_{k,2}\,|\,\bar{l}_{k},\bar{a}_{k-1}),$ (13)
-to obtain a sample $(l_{k},a_{k})$. We initialize by setting $F(\tilde{y}_{k+1}\,|\,\bar{l}_{k},do(\bar{a}_{k}),y_{k}=0)=w_{k,3}$. Then, for $j=k-1,\ldots,0$, we iteratively apply the transformations:
-$F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$ $=h\big{\{}F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0),$
-$\qquad\quad F(\tilde{y}_{k}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0);\Theta_{{k-1},j}\big{\}}$ (12)
-$F(\tilde{y}_{k+1}\,|\,\bar{l}_{j},do(\bar{a}_{k}),y_{k}=0)$ $=h^{-1}\big{\{}F(\tilde{y}_{k+1}\,|\,\bar{l}_{j+1},do(\bar{a}_{k}),y_{k}=0),$
+to obtain a sample $(l_{k},a_{k})$.
+We initialize by setting $F(\widetilde{y}_{k+1}\,|\,\bar{l}_{k},do(\bar{a}_{k}),y_{k}=0)=w_{k,3}$.
+Then, for $j=k-1,\ldots,0$, we iteratively apply the transformations:
+$F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$ $=\frac{v_{k-1,j}-C_{\widetilde{Y}_{k}L_{j}\mid\bar{L}_{j},do(\bar{A}_{k}),Y_{k-1}=0}(v_{k-1,j},q_{k-1,j};\Theta_{k-1,j})}{1-q_{k-1,j}},$ (14)
+$F(\widetilde{y}_{k+1}\,|\,\bar{l}_{j},do(\bar{a}_{k}),y_{k}=0)$ $=h^{-1}\big{\{}F(\widetilde{y}_{k+1}\,|\,\bar{l}_{j+1},do(\bar{a}_{k}),y_{k}=0),$
 $\qquad\qquad F(l_{j+1}\,|\,\bar{l}_{j},do(\bar{a}_{k}),y_{k}=0);\Theta_{k,j+1}\big{\}}.$ (15)
-By induction, the quantiles in the $h$-function in (12) are available from the previous time step $k-1$. We prove this claim as Proposition 5.1.
-When the recursion reaches $j=0$, we obtain the quantile $F(\tilde{y}_{k+1}\,|\,l_{0},do(\bar{a}_{k}),y_{k}=0)$. Applying the $h$-function one final time, we set:
-$u_{k+1}=F(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}))$ $=h^{-1}\big{\{}F(\tilde{y}_{k+1}\,|\,l_{0},do(\bar{a}_{k}),y_{k}=0),$
+Expression (14) follows Lemma 5.1, by induction, $v_{k-1,j}$ and $q_{k-1,j}$ are available from the previous time step $k-1$.
+When the recursion reaches $j=0$, we obtain the quantile $F(\widetilde{y}_{k+1}\,|\,l_{0},do(\bar{a}_{k}),y_{k}=0)$. Applying the $h$-function one final time, we set:
+$u^{*}_{k+1}=F(\widetilde{y}_{k+1}\,|\,do(\bar{a}_{k}))$ $=h^{-1}\big{\{}F(\widetilde{y}_{k+1}\,|\,l_{0},do(\bar{a}_{k}),y_{k}=0),$
 $\qquad\qquad F(l_{0}\,|\,do(\bar{a}_{k}),y_{k}=0);\Theta_{k,0}\big{\}}.$
-A sample of $\tilde{y}_{k+1}$ is then obtained via inversion:
-$\tilde{y}_{k+1}=F_{Y_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}^{-1}(u_{k+1}),$ (14)
-with the corresponding binary outcome $y_{k+1}=I(\tilde{y}_{k+1}<1)$.
-This completes the $k$th time step, yielding the sample $(l_{k},a_{k},\tilde{y}_{k+1},y_{k+1})$.
-If $y_{k+1}=1$, the simulation stops; otherwise, proceed to $k+1$.
-Crucially, our method avoids computing the full conditional distribution of $\tilde{Y}_{k+1}\mid\bar{L}_{k},do(\bar{A}_{k}),Y_{k}=0$, but rather requiring just a quantile corresponding to our simulated value; this is central to the efficiency of our approach.
-Equation (12) involves $k$ calculations of the distributions of each $L_{j}\,|\,\bar{L}_{j-1}$ among those who survived to time $k$, in which all the quantities, including quantiles and copulas, are available from the previous step $k-1$.
-In (13), we recursively apply $k+1$ inverse $h$-functions to obtain the marginal quantile of $F(\tilde{Y}_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0)$.
-Finally, we simulate $\tilde{Y}_{k+1}$ from the target marginal structural model via inversion.
-In total, this time step requires $42k+$ 4 functions, including the simulation of $L_{k}$ and $A_{k}$.
+A sample of $\widetilde{y}_{k+1}$ is then obtained via inversion:
+$\widetilde{y}_{k+1}=F_{Y_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}^{-1}(u^{*}_{k+1}),$
+with the corresponding binary outcome $y_{k+1}=I(\widetilde{y}_{k+1}<1)$.
+If $y_{k+1}=1$, the simulation stops; otherwise, before proceeding to $k+1$, we store and compute the following quantities for the next step:
+  • Store $v_{k,j}$ = $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$, for $j=0,\ldots,k$ , which have been computed in the current step via (14).
+  • Compute $q_{k,j}=F(\widetilde{y}_{k+1}=1\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$: initialize $q_{k,0}=F(\widetilde{y}_{k+1}=1\,|\,do(\bar{a}_{k}),y_{k}=0)$, then for $j=1,\ldots k$, recursively update:
+    $q_{k,j}$ $=F(\widetilde{y}_{k+1}=1\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$
+    $=h\big{\{}F(\widetilde{y}_{k+1}=1\,|\,\bar{l}_{j-2},do(\bar{a}_{k}),y_{k}=0),F(l_{j-1}\,|\,\bar{l}_{j-2},do(\bar{a}_{k}),y_{k}=0);\Theta_{k,j-1}\big{\}}$
+    $=h\big{\{}q_{k,j-1},v_{k,j-1};\Theta_{k,j-1}\big{\}}.$ (16)
+This completes the $k$th time step, yielding the sample $(l_{k},a_{k},\widetilde{y}_{k+1},y_{k+1})$, along with the quantities required for the next step. We formalize this claim as Proposition 5.2.
+Algorithm 2 Simulation algorithm for an individual $i$ up to the $K$th follow-up
+$k\leftarrow 0$
+while $k\leq K$ do
+     Sample $w_{k,1},w_{k,2},w_{k,3}$ independent standard uniform variables
+     $l_{k}\leftarrow F^{-1}_{L_{k}\mid Z,\bar{L}_{k-1},\bar{A}_{k-1},Y_{k}=0}(w_{k,1}\,|\,z,\bar{l}_{k-1},\bar{a}_{k-1})$
+     $a_{k}\leftarrow F^{-1}_{A_{k}\mid Z,\bar{L}_{k},\bar{A}_{k-1},Y_{k}=0}(w_{k,2}\,|\,z,\bar{l}_{k},\bar{a}_{k-1})$
+     if k = 0 then
+         $v_{k,j}=w_{0,1}$
+     else
+         for $j\leftarrow 1,\ldots,k$ do $\triangleright$ $L_{j}$ dist. in survivors per (14)
+              $v_{k,j}\leftarrow\frac{v_{k-1,j}-C(v_{k-1,j},q_{k-1,j};\Theta_{k,k-j})}{1-q_{k-1,j}}$
+         end for
+     end if
+     $\xi_{k+1}\leftarrow w_{k,3}$
+     for $j\leftarrow 0,\ldots,k$ do $\triangleright$ ‘Unweaving’ as per (15)
+         $\xi_{k-j}\leftarrow h^{-1}(\xi_{k+1-j},v_{k,k-j};\Theta_{K-j})$
+     end for
+     $\widetilde{y}_{k+1}\leftarrow F_{\widetilde{Y}_{k+1}\mid\bar{A}_{k},Y_{k-1}=0}^{-1}(\xi_{0}\,|\,do(\bar{a}_{k}))$
+     $y_{k+1}=I(\widetilde{y}_{k+1}>1)$
+     if $y_{k+1}=1$ then
+         Stop
+     else $\triangleright$ Prepare $q$s for the next time step per (16)
+         $q_{k,0}\leftarrow F_{\widetilde{Y}_{k+1}\mid\bar{A}_{k},Y_{k-1}=0}(\widetilde{Y}_{k+1}=1\,|\,do(\bar{a}_{k}))$
+         for $j\leftarrow 1,\ldots k$ do
+              $q_{k,j}\leftarrow h\big{\{}q_{k,j-1},v_{k,j-1};\Theta_{k,j-1}\big{\}}$
+         end for
+         $k\leftarrow k+1$
+     end if
+end while
+return $(\bar{l}_{k},\bar{a}_{k},\widetilde{y}_{k+1},\bar{y}_{k+1})$
+Crucially, our method avoids computing the full conditional distribution of $\widetilde{Y}_{k+1}\mid\bar{L}_{k},do(\bar{A}_{k}),Y_{k}=0$, but rather requiring just a quantile corresponding to our simulated value; this is central to the efficiency of our approach.
+(14) involves $k$ calculations of the distributions of each $L_{j}\,|\,\bar{L}_{j-1}$ among those who survived to time $k$, in which all the quantities, including quantiles and copulas, are available from the previous step $k-1$.
+In (15), we recursively apply $k+1$ inverse $h$-functions to obtain the marginal quantile of $F(\widetilde{Y}_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0)$.
+Finally, we simulate $\widetilde{Y}_{k+1}$ from the target marginal structural model via inversion.
+Before proceeding to the next time step $k+1$, we prepare the $q_{k,j}$ through $k+1$ recursive computations in (16).
+In total, this time step requires $53k+$ 5 functions, including the simulation of $L_{k}$ and $A_{k}$.
 We proceed to show that the proposed method works using mathematical induction.
@@
 Given the quantiles $w_{k,1},w_{k,2},w_{k,3}$ for each $k=0,\ldots,K-1$ and associated univariate distributions, together with copula parameters $\Theta_{k,j}$ for each $k=0,\ldots,K$ and $j=0,\ldots,k$, we can use the procedure outlined above to simulate a survival outcome according to the specified distributions.
-The proof is given in Appendix B.
+The proof is given in Appendix C.
+In the derivation and proof presented above, we define the latent variable $\widetilde{Y}$ as the incremental time to failure which naturally aligns with the Cox marginal structural model that we use as a working example.
+Importantly, a key strength of our approach lies in its generality: our proposed algorithm is fully agnostic to the choice of survival model and holds for any specification of the latent variable and its distribution.
+This is a fundamental result that enables the direct simulation of survival outcomes from any causal marginal structural model, whether continuous or discrete.
 Remark 5.3.
-To simulate a binary survival outcome directly from a discrete marginal structural model, such as that defined in (9), we adopt an approach analogous to Fan et al.
+To illustrate this point, consider the case of simulating a binary survival outcome from a discrete marginal structural model, such as that defined in (10).
+Similar to the framework in Fan et al.
+(2017), we assume that there exists a continuous latent random variable, $S_{k+1}$, whose distribution under the intervention $do(\bar{A}_{k})$ is $F_{S_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}$, such that the binary survival indicator is obtained by dichotomizing $S_{k+1}$ at a threshold $C$.
+This formulation allows dependencies between $Y_{k+1}$ and past covariates to be modelled via the latent variable $S_{k+1}$, and the simulation procedure follows the algorithm described above, replacing $\widetilde{Y}_{k+1}$ by $S_{k+1}$.
+Given the marginal quantile $u_{k+1}$, inversion yields $s_{k+1}=F^{-1}_{S_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}(u_{k+1}\,|\,do(\bar{a}_{k}),y_{k}=0)$ Finally, the binary outcome is obtained as $y_{k+1}=I(s_{k+1}>C)$.
+In this basic setting, simulating $K$ time steps for each individual requires $\sum_{k=0}^{K-1}(3k+5)=3K^{2}/2+7K/2$ calculations.This implies that the computational complexity scales quadratically with $K$.
+If additional time-varying confounders—either observed or unobserved—are incorporated, such that $\bm{L}$ becomes a $p$-dimensional vector, the number of iterations needed of (14), (15) and (16) scales proportionally with $p$.
+As a result, the total number of computations becomes $\sum_{k=0}^{K-1}(3pk+2p+3)=K(3K+1)p/2+3K$.
+Thus, computational complexity grows linearly with the number of confounders $p$.
+In the example presented in Appendix E, we replicate the simulation study by Seaman and Keogh (2024), simulating data featuring two time-varying confounders.
+However, it should be noted that if an individual experiences failure at any time step, the simulation terminates early.
+Hence, these stated complexities represent the maximum possible computational cost per individual.
+All of these calculations require evaluations of $h$-functions and their inverses and Copula distribution functions.
+For commonly used copulas, such evaluations are computationally efficient.
+5.2 Illustration: Concomitant Treatment
+We illustrate how the proposed simulation framework can support sample size calculation under complex, real-world scenarios.
+To do so, we construct a synthetic study inspired by the SOUL trial (Semaglutide cardiOvascular oUtcomes triaL) (McGuire et al., 2023).
+The SOUL trial is a randomized controlled study assessing the effects of oral semaglutide, a glucagon-like peptide-1 receptor agonist (GLP-RA), on cardiovascular risk in individuals with type 2 diabetes, and cardiovascular and/or kidney diseases.
+The primary endpoint is time to first occurrence of major adverse cardiovascular events (MACE), analyzed under the intention-to-treat principle.
+A design challenge in SOUL arises from the permitted use of concomitant medications such as SGLT2 inhibitors (SGLT2-i), and there is demonstrably more uptake among participants in the control arm.
+The uneven ‘drop-in’ treatment between the semaglutide and placebo arms may dilute the observed intention-to-treat effect, because both groups end up receiving some active treatment, blurring the distinction between them, and potentially undermining statistical power if unaccounted for during sample size calculation.
+Our goal is to show how the proposed simulation approach can capture such intercurrent events while preserving a predefined causal estimand.
+We simulate data over seven half-yearly follow-up periods.
+Let $B\in\{0,1\}$ indicate baseline cardiovascular disease, $L_{k}$ denote HbA1c at time $k$, $A\in\{0,1\}$ the randomized treatment (semaglutide), $S_{k}\in\{0,1\}$ SGLT2-i use at time $k$ and $Y_{k}$ the binary survival status at time $k$.
+We assume perfect adherence to $A$ and model event times using a Cox marginal structural model:
+$\lambda(\widetilde{y}_{k+1}\,|\,do(A,\bar{S}_{k}),Y_{k}=0)=\lambda_{0}\exp\left(\beta_{A}\,A+\beta_{S}\,S_{k}+\beta_{AS}\,AS_{k}\right),$ (17)
+where $\lambda_{0}$ is a baseline hazard assumed to be constant between follow-up visits and the survival indicator is calculated as $Y_{k+1}=I(\widetilde{Y}_{k+1}<1)$.
+Baseline characteristics and model parameters are loosely based on the design assumptions of the SOUL trial (McGuire et al., 2023), including an annual MACE rate of 3.5% and a hazard ratio of 0.83 for semaglutide (i.e.
+$\beta_{A}=\log(0.83)$).
+We construct 10 simulation scenarios by varying both the prevalence (0%, 5%, 10% and 15% annual drop-in) in the placebo group and efficacy (hazard ratios of 0.95,0.9,0.85) of SGLT2-i use.
+These hazard ratios reflect estimates reported for SGLT2 inhibitors in McGuire et al.
 (2021).
-Specifically, we assume that there exists a continuous latent random variable, $S_{k+1}$, whose distribution under the intervention $do(\bar{A}_{k})$ is $F_{S_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}$, such that the binary survival indicator is obtained by dichotomizing $S_{k+1}$ at a constant threshold $C$.
-This formulation allows dependencies between $Y_{k+1}$ and past covariates to be modelled via the latent variable $S_{k+1}$, and the simulation procedure follows the algorithm described above, replacing $\tilde{Y}_{k+1}$ by $S_{k+1}$.
-Given the marginal quantile $u_{k+1}$, inversion yields $s_{k+1}=F^{-1}_{S_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}(u_{k+1}\,|\,do(\bar{a}_{k}),y_{k}=0)$ Finally, the binary outcome is obtained as $y_{k+1}=I(s_{k+1}>C)$.
-5.2 Extensions and computational complexity
-In this basic setting, simulating $K$ time steps for each individual requires $\sum_{k=0}^{K-1}(2k+4)=K^{2}+3K$ calculations.
-This implies that the computational complexity grows quadratically with $K$.
-However, if an individual experiences failure at any time step, the simulation terminates early, meaning that this complexity represents the maximum possible computational cost per individual.
-A key extension involves simulating multiple outcomes, and the approach depends on whether the additional outcomes pertain to survival or non-survival events.
-For non-survival outcomes simulated from marginal structural models—as opposed to conditional distributions—the number of calculations to derive $F(\bm{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y^{\triangle}_{k}=0)$ increases proportionally with the number of outcomes, $m$.
-Here, $\bm{Y}_{k+1}$ is an $m$-dimensional vector, and $Y_{k}^{\triangle}$ denotes the primary survival outcome, with all other outcomes assumed to be non-survival.
-To incorporate dependence between outcomes given past variables, we must specify the joint distribution $\bm{Y}_{k+1}\,|\,\bar{L}_{k},\bar{A}_{k}$ and simulate from this using a pair-copula construction.
-The total number of calculations is $\sum_{k=0}^{K-1}\big{\{}(k+3)m+k+1\big{\}}=(K+5)Km/2+K(K+1)/2$, which scales linearly with $m$.
-For competing risks, such as death from other causes, additional calculations are needed to determine the distribution of $L$ given survival from all risks, as well as its distribution conditioned on the occurrence of each specific risk.
-Consequently, the computational cost scales proportionally with $m$.
-Furthermore, iterative evaluations of inverse $h$-functions are required to obtain marginal interventional quantiles for each risk.
-The total number of calculations is given by $\sum_{k=0}^{K-1}(2mk+3m+1)=(K^{2}+2K)m+K$ which grows linearly with $m$.
-The simulation algorithm for a setting with a single competing risk is described in Appendix C.
-Another important extension involves including multiple time-varying confounders, observed or unobserved, such that $\bm{L}$ becomes a vector of $p$ variables.
-The number of calculations required for the distributions of ¯$\bm{\bar{}}L_{k-1}$, and $F(\tilde{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0)$ among survivors at $k$ scales proportionally with $p$.
-The total number of computations becomes $\sum_{k=0}^{K-1}(2pk+3p+2)=K(K+2)p+2K$.
-Thus, the computational complexity grows linearly with $p$.
-In the example presented in Section 5.3, we simulate data with two time-varying confounders.
-All of these calculations require evaluations of $h$-functions and their inverses. For commonly used copulas with analytic $h$-functions, these evaluations are computationally efficient and fast.
-5.3 Illustrative example
-We use an example to demonstrate the proposed algorithm simulating longitudinal time-to-event data, following a setup similar to Seaman and Keogh (2024).
-We consider baseline variables $B$ and $C$, two time-varying covariates $L_{k}$ and $W_{k}$, and treatment $A_{k}$ at follow-up $k$.
-The time-to-event is $T$, with $Y_{k+1}=I(T<k+1)$ indicating failure status observed at the end of the $k$th interval.
-The auxiliary continuous variable $\tilde{Y}_{k+1}$ is defined as previously.
-The data-generating models for the baseline and time-varying covariates, and treatment status are detailed in Appendix D.
-We assume a Cox marginal structural model given by:
-$\lambda(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),c,y_{k}=0)=\lambda_{0}\exp\left(\beta_{C}\,c+\beta_{A}\,a_{k}+\beta_{CA}\,c\,a_{k}\right),$ (15)
-where $\lambda_{0}=\exp(-2)$ is the baseline hazard.
-We set $\beta_{C}=-0.5$ and vary the values of $(\beta_{A},\beta_{CA})$ to construct scenarios with different strengths of treatment: a strong effect scenario $(-0.5,0.3)$ and a weak effect scenario $(0.2,0.1)$.
-This parameterization assumes a constant hazard rate between follow-up visits, influenced by the treatment status at the most recent visit.
-This implies that given survival at follow-up time $k$, the incremental survival time $\tilde{Y}_{k+1}$ follows an exponential distribution with the rate equal to the hazard above:
-$\tilde{Y}_{k+1}\sim\operatorname{Exp}(\lambda_{0}\exp\left(\beta_{C}\,c+\beta_{A}\,a_{k}+\beta_{CA}\,c\,a_{k}\right)).$
-If $\tilde{Y}_{k+1}<1$ then $Y_{k+1}=1$, and 00 otherwise.
-We set the total number of follow-ups $K$ to 6.
-We set the copulas describing the association between $L_{j}$ ($j=0,\ldots,k$) and $Y_{k+1}$ to be Student t with parameter $\rho_{L}$ and 5 degrees of freedom, although each pairwise dependence could, in principle, be distinct.
-Similarly, we model the copulas capturing the dependence between $Y_{k+1}$ and each $W_{j}$ ($j=0,\ldots,k$), given $L_{k}$, as Gaussian with correlation $\rho_{W}$, though again we note that different copulas could be specified for each time point.
-We vary the values of $\rho_{L}$ and $\rho_{W}$ to construct scenarios with different strengths of confounding (see Table 6 in Appendix D).
-We simulate data for sample size $n=\{1000,2000,5000\}$.
-Remark 5.3.
-Table 7 (in Appendix D) reports the runtime required to simulate data under the setup for the large causal effect with medium confounding scenario.
-The simulations were carried out on a laptop using a single score without parallelization.
-Generating data for $n$ up to 10,000 units takes only a few seconds, while even for a very large sample size ($n=50,\!000$) and 11 time steps ($K=10$), the runtime remains under one minute.
-This computational efficiency is crucial for tasks such as method evaluation and sample size calculations in clinical trial design, where simulations must be repeated across various parameter values.
-We contrast two approaches for fitting the Cox marginal structural model specified in (15): stabilized inverse-probability-weighting and unweighted estimation.
-Details on weight calculation are provided in Appendix D.
-Table 4 summarizes the bias of estimates from 500 replications, compared to corresponding Monte Carlo standard errors as a measure of uncertainty.
-As the sample size increases, generally, the biases of estimates for $\beta_{C}$, $\beta_{A}$ and $\beta_{AC}$ obtained with inverse probability weights converge to zero, with their absolute values substantially smaller than the corresponding Monte Carlo standard errors.
-In contrast, the unweighted estimator exhibits considerable bias, particularly for the estimate of the treatment effect, $\beta_{A}$, which persists even at $n=5000$.
-The magnitude of this bias far exceeds the Monte Carlo standard errors.
-The unweighted estimates are most biased under the high level of confounding by $L$ and $W$, since this is not adjusted for.
-Level of confounding Effect size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ IPTW Unweighted IPTW Unweighted IPTW Unweighted High $1000$ $-0.04(2.27)$ $-1.05(1.54)$ $-0.14(2.31)$ $3.73(1.73)$ $0.27(3.09)$ $0.49(2.22)$ $\beta_{A}=-0.5$ $2000$ $-0.07(1.68)$ $-1.03(1.06)$ $-0.02(1.63)$ $3.81(1.21)$ $0.02(2.09)$ $0.34(1.51)$ $5000$ $0.02(1.25)$ $-0.96(0.68)$ $0.10(1.08)$ $3.89(0.76)$ $-0.09(1.58)$ $0.24(0.97)$ $1000$ $-0.03(2.35)$ $-0.93(1.59)$ $0.02(1.99)$ $3.64(1.46)$ $0.03(2.78)$ $0.44(1.94)$ $\beta_{A}=0.2$ $2000$ $-0.11(1.60)$ $-0.93(1.11)$ $0.02(1.49)$ $3.65(1.18)$ $0.17(2.02)$ $0.49(1.51)$ $5000$ $-0.03(1.14)$ $-0.95(0.71)$ $0.04(1.01)$ $3.65(0.71)$ $0.07(1.36)$ $0.51(0.89)$ Medium $1000$ $-0.07(2.04)$ $-0.43(1.51)$ $-0.13(2.19)$ $1.90(1.73)$ $0.23(2.91)$ $0.08(2.22)$ $\beta_{A}=-0.5$ $2000$ $-0.02(1.60)$ $-0.35(1.01)$ $0.00(1.65)$ $1.95(1.22)$ $0.09(2.27)$ $-0.01(1.55)$ $5000$ $-0.09(1.10)$ $-0.42(0.64)$ $-0.01(1.21)$ $1.97(0.80)$ $0.07(1.59)$ $-0.02(1.03)$ $1000$ $-0.12(2.06)$ $-0.36(1.43)$ $-0.04(2.02)$ $1.75(1.54)$ $0.14(2.70)$ $0.09(1.92)$ $\beta_{A}=0.2$ $2000$ $-0.10(1.62)$ $-0.42(1.08)$ $0.03(1.49)$ $1.68(1.11)$ $0.09(2.09)$ $0.17(1.44)$ $5000$ $-0.03(1.04)$ $-0.39(0.66)$ $0.03(0.95)$ $1.78(0.68)$ $0.03(1.32)$ $0.14(0.88)$ Low $1000$ $0.01(1.82)$ $0.46(1.29)$ $-0.23(2.36)$ $-1.95(1.83)$ $0.21(2.90)$ $-0.20(2.20)$ $\beta_{A}=-0.5$ $2000$ $-0.10(1.32)$ $0.44(0.89)$ $-0.04(1.69)$ $-1.83(1.25)$ $0.07(2.13)$ $-0.30(1.59)$ $5000$ $-0.02(1.04)$ $0.47(0.58)$ $0.01(1.07)$ $-1.77(0.84)$ $0.03(1.46)$ $-0.32(0.99)$ $1000$ $0.04(1.98)$ $0.53(1.30)$ $0.05(1.98)$ $-1.88(1.46)$ $-0.03(2.65)$ $-0.31(1.88)$ $\beta_{A}=0.2$ $2000$ $-0.02(1.45)$ $0.43(0.99)$ $0.04(1.43)$ $-1.85(1.10)$ $0.01(1.87)$ $-0.26(1.38)$ $5000$ $-0.01(0.92)$ $0.43(0.59)$ $0.07(0.95)$ $-1.89(0.68)$ $-0.06(1.25)$ $-0.25(0.84)$
-Table 4: Bias and Monte Carlo standard errors (×$\times 10$) of estimates produced by inverse probability of treatment weighted (‘IPTW’) and unweighted Cox regressions, under different strengths of confounding, effect sizes, and sample sizes.
-Table 5 presents coverage rates of the 95% confidence intervals.
-Confidence intervals for the weighted estimated, computed using both sandwich variance estimation and bootstrap resampling, achieve coverage close to the nominal level, with slight undercoverage.
-In contrast, the sandwich-based confidence intervals for the unweighted estimator perform poorly, especially for the treatment effect estimate, $\beta_{A}$.
-Level of confounding Effect Size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ W-S W-B UW W-S W-B UW W-S W-B UW High 1000 93.8% 95.2% 88.4% 93.4% 92.8% 42.2% 93.8% 93.8% 93.4% $\beta_{A}=-0.5$ 2000 94.0% 94.0% 84.4% 95.2% 95.6% 12.4% 95.6% 95.6% 94.6% 5000 93.2% 92.8% 70.2% 94.4% 93.4% 0.2% 94.2% 93.6% 93.8% 1000 92.8% 93.8% 91.8% 96.2% 95.8% 31.6% 93.4% 94.0% 95.6% $\beta_{A}=0.2$ 2000 95.2% 95.0% 87.6% 94.6% 95.0% 9.8% 94.8% 94.6% 92.2% 5000 93.6% 93.6% 72.4% 93.8% 93.0% 0.2% 94.0% 94.0% 90.0% Medium 1000 93.8% 94.2% 93.4% 94.8% 93.6% 79.6% 93.8% 94.0% 93.6% $\beta_{A}=-0.5$ 2000 93.0% 93.8% 92.4% 95.2% 94.6% 64.8% 92.2% 93.4% 96.2% 5000 95.4% 94.8% 90.0% 93.2% 92.6% 30.2% 93.0% 93.8% 93.6% 1000 94.2% 94.8% 95.0% 95.0% 93.8% 77.2% 94.2% 93.2% 95.8% $\beta_{A}=0.2$ 2000 94.6% 93.8% 91.6% 93.6% 93.2% 64.0% 92.8% 92.2% 94.2% 5000 95.4% 94.6% 91.6% 95.4% 95.2% 26.8% 94.8% 94.4% 95.4% Low 1000 95.4% 95.4% 93.0% 95.0% 95.2% 84.0% 95.2% 95.2% 94.6% $\beta_{A}=-0.5$ 2000 95.6% 96.4% 93.0% 95.8% 96.0% 71.6% 95.0% 95.6% 94.6% 5000 93.2% 93.2% 87.2% 94.2% 94.8% 42.6% 94.0% 94.4% 94.4% 1000 94.6% 95.6% 94.8% 96.4% 96.6% 77.4% 92.8% 94.8% 96.0% $\beta_{A}=0.2$ 2000 94.8% 95.6% 92.8% 94.4% 94.6% 60.6% 95.0% 95.2% 94.2% 5000 94.8% 94.2% 89.4% 95.0% 95.4% 20.0% 95.0% 94.2% 93.8%
-Table 5: Coverage rates for IPTW-Sandwich (W-S), IPTW-Bootstrap (W-B), and unweighted (UW) estimators of $\beta_{C}$, $\beta_{A}$, and $\beta_{AC}$ under different strengths of confounding, effect sizes, and sample sizes.
-In summary, the asymptotically unbiased nature of the inverse probability weighted estimates, compared to the biased unweighted estimates, further confirms that the simulated data were generated in accordance with the Cox model in (15), and that confounding by $L$ and $W$ was appropriately introduced.
+Full specifications of the data-generating model and scenario parameters are provided in Appendix D.
+For each scenario and trial sizes of $n=\{5,\!000,10,\!000,20,\!000\}$, we simulate 400 datasets and fit a Cox MSM using inverse probability of treatment weights with stabilized weights.
+Naïve unweighted estimators are also computed for comparison.
+For reference, the total randomized population in SOUL is 9,650.
+Figure 3: Boxplots of the bias for $\beta_{A}$, $\beta_{S}$ and $\beta_{AS}$ by inverse probability weighting (IPTW) across different drop-in scenarios.
+Figure 3 and Figure 4 display boxplots of the estimation biases for the coefficients $\beta_{A}$, $\beta_{S}$ and $\beta_{AS}$ by inverse-probability-weighted and unweighted Cox regression, respectively.
+Across all scenarios, the inverse-probability-weighted estimates exhibit biases that are centered at zero.
+As the sample size increases, the distribution of these biases becomes narrower and more tightly concentrated around zero, indicating consistency.
+In contrast, unweighted estimates show substantial and persistent bias across all scenarios, with no evidence of convergence toward zero as the sample size increases.
+Table 4 presents the empirical coverage: the inverse-probability-weighted method achieves coverage close to the nominal 95% level, whereas the unweighted approach yields empirical coverage well below the nominal level, reflecting the impact of the time-varying confounding introduced to the data.
+Figure 4: Boxplots of the bias for $\beta_{A}$, $\beta_{S}$ and $\beta_{AS}$ by unweighted Cox regression across different drop-in scenarios.
+Effect of SGLT2-i Prevalence in placebo $n$ $\beta_{A}$ $\beta_{S}$ $\beta_{AS}$ IPTW Unweighted IPTW Unweighted IPTW Unweighted Strong effect $5000$ $96.5\%$ $65.0\%$ $94.0\%$ $0.0\%$ $93.0\%$ $74.0\%$ 5% drop-in $10000$ $94.8\%$ $47.3\%$ $95.0\%$ $0.0\%$ $94.3\%$ $49.0\%$ $20000$ $95.0\%$ $18.5\%$ $95.0\%$ $0.0\%$ $95.0\%$ $25.8\%$ $5000$ $94.0\%$ $28.0\%$ $95.0\%$ $0.0\%$ $92.3\%$ $64.5\%$ 10% drop-in $10000$ $95.8\%$ $4.5\%$ $96.5\%$ $0.0\%$ $95.3\%$ $42.5\%$ $20000$ $96.5\%$ $0.0\%$ $95.8\%$ $0.0\%$ $96.3\%$ $16.3\%$ $5000$ $92.8\%$ $4.8\%$ $94.0\%$ $0.0\%$ $94.0\%$ $60.8\%$ 15% drop-in $10000$ $94.5\%$ $0.0\%$ $96.0\%$ $0.0\%$ $94.3\%$ $34.0\%$ $20000$ $93.0\%$ $0.0\%$ $95.0\%$ $0.0\%$ $94.0\%$ $7.5\%$ Moderate effect $5000$ $95.3\%$ $68.0\%$ $96.5\%$ $0.0\%$ $95.5\%$ $69.3\%$ 5% drop-in $10000$ $94.5\%$ $43.3\%$ $93.8\%$ $0.0\%$ $94.3\%$ $50.8\%$ $20000$ $95.3\%$ $17.5\%$ $96.0\%$ $0.0\%$ $93.8\%$ $25.0\%$ $5000$ $96.0\%$ $25.5\%$ $97.3\%$ $0.0\%$ $95.5\%$ $63.8\%$ 10% drop-in $10000$ $95.0\%$ $4.5\%$ $96.3\%$ $0.0\%$ $93.5\%$ $40.8\%$ $20000$ $95.8\%$ $0.3\%$ $96.3\%$ $0.0\%$ $95.3\%$ $14.8\%$ $5000$ $94.3\%$ $4.8\%$ $95.5\%$ $0.0\%$ $95.8\%$ $58.0\%$ 15% drop-in $10000$ $95.3\%$ $0.0\%$ $94.5\%$ $0.0\%$ $94.0\%$ $37.0\%$ $20000$ $92.0\%$ $0.0\%$ $95.3\%$ $0.0\%$ $94.0\%$ $10.0\%$ Weak effect $5000$ $96.3\%$ $65.5\%$ $95.3\%$ $0.0\%$ $93.5\%$ $63.5\%$ 5% drop-in $10000$ $96.0\%$ $42.8\%$ $97.0\%$ $0.0\%$ $95.3\%$ $45.3\%$ $20000$ $95.0\%$ $13.3\%$ $95.0\%$ $0.0\%$ $95.0\%$ $18.5\%$ $5000$ $96.0\%$ $26.8\%$ $96.0\%$ $0.0\%$ $94.5\%$ $59.8\%$ 10% drop-in $10000$ $96.3\%$ $5.0\%$ $96.5\%$ $0.0\%$ $95.0\%$ $37.0\%$ $20000$ $95.5\%$ $0.0\%$ $96.8\%$ $0.0\%$ $96.3\%$ $11.5\%$ $5000$ $93.3\%$ $5.3\%$ $95.8\%$ $0.0\%$ $93.8\%$ $57.5\%$ 15% drop-in $10000$ $93.3\%$ $0.0\%$ $93.5\%$ $0.0\%$ $92.5\%$ $34.8\%$ $20000$ $95.5\%$ $0.0\%$ $95.5\%$ $0.0\%$ $95.0\%$ $7.3\%$ $5000$ $96.0\%$ $96.0\%$- -- -- -- No drop-in $10000$ $94.5\%$ $94.5\%$- -- -- -- $20000$ $93.5\%$ $93.5\%$- -- -- --
+Table 4: Empirical coverage rates of 95% confidence intervals of inverse-probability-weighted (IPTW) and unweighted Cox regression methods.
+These results confirm that the simulated data adhere to the marginal structural model specified in (17), and that time-varying confounding is induced.
+We now demonstrate how the proposed simulation framework can inform trial design, specifically for sample size determination under realistic scenarios involving intercurrent events.
+The SOUL trial was designed to evaluate the effect of semaglutide on CV outcomes under the intention-to-treat principle, contrasting the outcomes of individuals randomized to semaglutide ($A=1$) versus placebo ($A=0$).
+Accordingly, we fit the following Cox proportional hazards model:
+$\lambda(t\,|\,A)=\lambda_{0}\exp(\beta_{ITT}A),$
+where $t$ is time to the first occurrence of MACE and $\exp(\beta_{ITT})$ represents the hazard ratio associated with being assigned semaglutide, irrespective of SGLT2-i use.
+Table 5 reports the estimated intention-to-treat effects and empirical power based on a one-sided type I error rate of $0.025$.
+In the hypothetical scenario with no SGLT2-i use in either arm, the estimated intention-to-treat effect closely approximates the true hazard ratio for semaglutide (0.83).
+Under this scenario, a trial with 10,000 participants achieves approximately 85% power.
+However, in all other scenarios, the intention-to-treat estimates are attenuated due to drop-in and consequently reduce power.
+To maintain at least 80% power, our simulations suggest that a sample size of 20,000 is sufficient but may be unnecessarily large and resource-intensive.
+Conversely, a trial with 5,000 participants is consistently underpowered, even in the absence of drop-in.
+A sample size of 10,000 appears to strike a reasonable balance: it achieves adequate power in scenarios with weak SGLT2-i effects, but falls short when the effect of SGLT2-i is strong or if the prevalence of drop-in in the placebo arm exceeds 5%.
+For the moderate scenarios, where SGLT2 inhibitors have a moderate effect, 10,000 participants is borderline sufficient, but if drop-in rates in the placebo arm exceed 5%, power may fall below acceptable thresholds.
+These findings suggest that 10,000 participants may be an appropriate design choice, but the sample size should be adjusted based on the plausibility of drop-in and effect size assumptions.
+The 10 scenarios considered here are illustrative.
+The proposed simulation framework is highly flexible.
+For example, users can vary the marginal effect of semaglutide by adjusting $\beta_{A}$, alter the baseline incidence rate, or introduce more time-varying covariates.
+Because the algorithm relies on analytic function evaluations, it is computationally efficient: simulating 10,000 individuals over seven time steps requires approximately 40 seconds on a single core.
+This efficiency makes it practical and feasible to conduct large-scale simulation studies using parallel computation to support power analyses and inform robust trial design.
+Effect of SGLT2-i Prevalence in placebo n Estimate Hazard Ratio Power Strong effect $5000$ $-0.164$ $0.849\%$ $46.5\%$ 5% drop-in $10000$ $-0.165$ $0.848\%$ $74.3\%$ $20000$ $-0.169$ $0.845\%$ $97.5\%$ $5000$ $-0.153$ $0.858\%$ $43.0\%$ 10% drop-in $10000$ $-0.148$ $0.862\%$ $65.5\%$ $20000$ $-0.151$ $0.860\%$ $93.0\%$ $5000$ $-0.143$ $0.867\%$ $37.5\%$ 15% drop-in $10000$ $-0.143$ $0.867\%$ $63.8\%$ $20000$ $-0.135$ $0.874\%$ $87.8\%$ Moderate effect $5000$ $-0.176$ $0.839\%$ $52.0\%$ 5% drop-in $10000$ $-0.173$ $0.841\%$ $81.0\%$ $20000$ $-0.179$ $0.836\%$ $98.8\%$ $5000$ $-0.160$ $0.852\%$ $45.0\%$ 10% drop-in $10000$ $-0.164$ $0.848\%$ $75.8\%$ $20000$ $-0.161$ $0.851\%$ $97.8\%$ $5000$ $-0.161$ $0.851\%$ $46.0\%$ 15% drop-in $10000$ $-0.157$ $0.855\%$ $76.3\%$ $20000$ $-0.154$ $0.857\%$ $94.5\%$ Weak effect $5000$ $-0.174$ $0.841\%$ $52.0\%$ 5% drop-in $10000$ $-0.181$ $0.834\%$ $84.5\%$ $20000$ $-0.184$ $0.832\%$ $98.3\%$ $5000$ $-0.176$ $0.838\%$ $53.3\%$ 10% drop-in $10000$ $-0.180$ $0.835\%$ $82.8\%$ $20000$ $-0.176$ $0.839\%$ $98.3\%$ $5000$ $-0.182$ $0.834\%$ $55.0\%$ 15% drop-in $10000$ $-0.176$ $0.838\%$ $82.3\%$ $20000$ $-0.173$ $0.841\%$ $98.3\%$ $5000$ $-0.187$ $0.830\%$ $56.8\%$ No drop-in $10000$ $-0.183$ $0.833\%$ $85.0\%$ $20000$ $-0.187$ $0.830\%$ $99.5\%$
+Table 5: Under the intention-to-treat principal, estimates of $\beta_{ITT}$, hazard ratios ($\exp{\beta_{ITT}}$), and power using a one-sided type I error rate of $0.025$ for various scenarios and sample sizes ($n$).
+5.3 Competing risks
+An important consideration in research concerning time-to-event outcomes is that of competing risks; these are events whose occurrence prevents the observation of the outcome of interest—for example, non-cardiovascular death in studies of MACE.
+Classical literature defines various statistical estimands for competing risk settings, including the marginal, cause-specific and subdistribution hazards (see, for example, Pintilie, 2007; Geskus, 2016; Young et al., 2020).
+These estimands differ in interpretation and causal relevance.
+Our proposed simulation framework offers the flexibility to efficiently simulate data from marginal structural models that target the controlled direct effect:
+$F(\widetilde{Y}_{k+1}\,|\,do(\bar{A}_{k}=\bar{a}_{k},\bar{D}_{k}=\bar{0}_{k}),Z=z,Y_{k}=0)=f(\bar{a}_{k},z;\alpha),$ (18)
+where $\{D_{k}\}_{k\in\mathcal{K}_{+}}$ denote the sequence of competing events and $do(\bar{D}_{k}=\bar{0}_{k})$ represents a hypothetical intervention eliminating these events.
+This aligns with the definition of marginal hazard in Young et al.
+(2020).
+Under this framework, competing risks are essentially treated as equivalent to censoring.
+To generate the competing event process, two approaches can be considered.
+The first specifies a fully conditional distribution: $F(\widetilde{D}_{k+1}\,|\,\bar{L}_{k},\bar{A}_{k},Z,Y_{k}=0)$.
+Under this approach, the outcome process $Y_{1:K}$ is simulated according to Algorithm 2, targeting the causal margin defined in (18).
+Once simulation of the failure process is complete—suppose the trajectory terminates at time step $K^{\prime}$ when $Y_{K^{\prime}+1}=1$—the competing event process is then simulated for each $k\in\{0,\ldots,K^{\prime}\}$.
+At each of these time steps, $\widetilde{D}_{k+1}$ is drawn from the specified conditional distribution $F(\widetilde{D}_{k+1}\,|\,\bar{L}_{k},\bar{A}_{k},Z,Y_{k}=0)$.
+If a competing event occurs before time $K^{\prime}$, the individual’s trajectory is truncated at the corresponding time step.
+Censoring events, such as loss to follow-up, can also be simulated using this approach.
+Alternatively, a marginal structural model may be specified for the competing event hazard, mirroring the approach used for the primary outcome:
+$F(\widetilde{D}_{k+1}\,|\,do(\bar{A}_{k}=\bar{a}_{k},\bar{Y}_{k}=\bar{0}_{k}),Z=z,D_{k}=0)=g(\bar{a}_{k},z;\beta).$ (19)
+Additionally, we impose the following conditional independence:
+[…]
+which states that, under intervention, and conditional on baseline and time-varying covariates, $\widetilde{D}_{k+1}$ and $\widetilde{Y}_{k+1}$ are independent.
+Graphically, this corresponds to the absence of a direct edge between $Y_{k+1}$ and $D_{k+1}$ in the underlying causal graph—an edge that would not be identifiable from observed data in any case.
+If both the outcome and the competing event are simulated using MSMs, the simulation proceeds as follows:
+  1. 1.
+    Simulate baseline covariates $Z$. Set $Y_{0}=D_{0}=0$.
+  2. 2.
+    For time point $k=0,\dots,K$:
+    1. (a)
+      If $Y_{k}=0$ or $D_{k}=0$, simulate $L_{k},A_{k}$; otherwise, stop;
+    2. (b)
+      If $Y_{k}=0$, simulate $Y_{k}$ following Algorithm 2; otherwise set $Y_{k}$ to null;
+    3. (c)
+      If $D_{k}=0$, simulate $D_{k}$ following Algorithm 2, substituting $Y$ with $D$, along with the corresponding distributions and copulas; otherwise, set $D_{k}$ to null.
+  3. 3.
+    Truncate this individual’s trajectory at the first occurrence of either the outcome of interest or the competing event.
+Notably, in step (b), the decision to simulate $Y_{k+1}$ depends solely on $Y_{k}$ irrespective of $D_{k}$ and vice versa for step (c).
+This reflects the fact that we target a controlled direct effect, as specified in (18) and (19), where the other event is conceptually eliminated through intervention.
+This approach is particularly helpful in settings like the estrogen therapy example in Young et al.
+(2020).
+By explicitly specifying the direction and magnitude of the direct effects of treatment on both prostate cancer mortality and other-cause mortality, one can simulate data to understand how these effects jointly determine the total effect, i.e.
+the cause-specific hazard.
+This framework can inform trial design: clinicians can input plausible direct effect assumptions and compute the sample size required to achieve adequate power for detecting a treatment effect on the cause-specific hazard.
 6 Discussion
@@
 In Section 5, we extended this method to survival models and time-to-event outcomes, which are common in clinical trials.
-Our methods are both conceptually and computationally simple, which crucial for practical applications such as sample size calculation.
-[…]
-Appendix A Extension to Include Time-Varying Latent Variables
-We provide an example of how to incorporate a latent process that confounds the covariates and the final outcome following the algorithm outlined in Section 4.
-The steps to simulate longitudinal data from the model depicted in Figure 3 are as follows:
+The method’s computational efficiency, combined with its analytic precision, makes it well-suited to practical tasks such as simulation-based method evaluation and sample size determination.
+In Section 5.3, we discussed how our proposed simulation framework can flexibly generate data with competing risks from marginal structural models targeting the marginal hazard.
+However, extending the framework to simulate data under alternative estimands—specifically, the cause-specific and subdistribution hazards, which represent total effects—remains nontrivial.
+Future work may explore how to extend the proposed method to support simulation from marginal structural models that target these alternative causal quantities.
+A limitation of the method is difficulty in enforcing Markov-type conditional independences between the outcome and the history of time-varying confounders.
+For instance, the pair-copula construction we employed in Section 4.2 to model the association between $Y$ and $L_{0},\dots,L_{k}$ makes it difficult to impose restrictions that might be considered natural; for example, that the distribution of the outcome only depends upon the most recent few values of a time-varying covariate.
+Parameterizing such a model is infeasible unless all the bivariate copulas in the pair-copula construction are Gaussian; in this case, conditional independence can be enforced by appropriately choosing the partial correlation parameters.
+If enforcing such Markov conditions is absolutely necessary, we would suggest using either this Gaussian approximation or numerical integration.
+However, numerical integration in high-dimensions is well known to be infeasible.
+Appendix A Extension to include multiple time-varying variables
+We outline how the proposed algorithm in Section 4 can be extended to include multiple time-varying covariates.
+From a simulation perspective, there is no distinction between whether such covariate is observed or unobserved; for illustrative purposes, we assume that the additional time-varying variable $H$ follows a latent process$\{H_{k}\}_{k\in\mathcal{K}}$.
+The steps for simulate data from the model depicted in Figure 5 are as follows:
 […]
@@
 As the dimensionality of time-varying confounders, either observed or unobserved, the computational complexity grows linearly.
-Appendix B Proof of Proposition 5.1
+Appendix B Proof of Lemma 5.1
+Proof.
+By definition:
+$F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$
+[…]
+[…]
+where $q_{k-1,j}=F(\widetilde{y}_{k}=1\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0)$.
+Recall the result from Joe (1996) in (3), let $v_{k-1,j}=F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0)$ and $u=F(\widetilde{y}_{k}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0)$, then the numerator in (20) can be written as:
+[…]
+$=C(v_{k-1j},1)-C(v_{k-1j},q_{k-1,j})$
+$=v_{k-1j}-C(v_{k-1j},q_{k-1,j})$
+where the copula $C$ is $C_{L_{j}\widetilde{Y}_{k}\mid\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k-1}=0}(\cdot;\Theta_{k-1,j})$, specifically. ∎
+Appendix C Proof of Proposition 5.2
 Proof.
@@
 Base case ($k=0$):
-Initialize the algorithm by specifying the cumulative distribution functions $F_{L_{0}}$, $F_{A_{0}\mid L_{0}}$, $F_{\tilde{Y}_{1}\mid do(A_{0})}$, and the copula $C_{\tilde{Y}_{1}L_{0}\mid do(A_{0})}$, parameterized by $\theta_{0,0}$.
+Initialize the algorithm by specifying the cumulative distribution functions $F_{L_{0}}$, $F_{A_{0}\mid L_{0}}$, $F_{\widetilde{Y}_{1}\mid do(A_{0})}$, and the copula $C_{\widetilde{Y}_{1}L_{0}\mid do(A_{0})}$, parameterized by $\Theta_{0,0}$.
 Draw $w_{0,1}$, $w_{0,2}$, $w_{0,3}\sim U(0,1)$ and compute:
-$l_{0}=F^{-1}_{L_{0}}(w_{0,1})$ $a_{0}=F^{-1}_{A_{0}\mid L_{0}}(w_{0,2}\,|\,l_{0}),$ (16)
+$l_{0}=F^{-1}_{L_{0}}(w_{0,1})$ $a_{0}=F^{-1}_{A_{0}\mid L_{0}}(w_{0,2}\,|\,l_{0}),$
 to obtain a sample $(l_{0},a_{0})$. Using the transformation:
-$F_{\tilde{Y}_{1}\,|\,do(A_{0})}(\tilde{y}_{1}\,|\,do(a_{0}))$ $=h^{-1}\big{\{}F_{\tilde{Y}_{1}\,|\,L_{0},do(A_{0})}(\tilde{y}_{1}\,|\,l_{0},do(a_{0})),F_{L_{0}}(l_{0});\theta_{0,0}\big{\}},$ (17)
-set $w_{0,3}=F_{\tilde{Y}_{1}\,|\,L_{0},do(A_{0})}(\tilde{y}_{1}\,|\,l_{0},do(a_{0}))$ and substitute $F_{L_{0}}(l_{0})$ with $w_{0,1}$, yielding:
-$u_{1}=h^{-1}\left(w_{0,3},w_{0,1};\theta_{0,0}\right).$ (18)
-By inversion, compute $\tilde{y}_{1}=F_{Y_{1}\mid do(A_{0})}^{-1}(u_{1})$, and the corresponding binary outcome $y_{1}=I(\tilde{y}_{1}<1)$.
-This completes the first iteration, yielding a sample $(l_{0},a_{0},\tilde{y}_{1},y_{1})$.
-If $y_{1}=1$, stop; otherwise, proceed to $k=1$.
-For $k=1$, define the necessary cumulative distribution functions $F_{L_{1}\,|\,L_{0},A_{0}}$, $F_{A_{1}\mid\bar{L}_{1},A_{0}}$, $F_{\tilde{Y}_{2}\mid do(\bar{A}_{1})}$, and copulas $C_{\tilde{Y}_{2}L_{1}\mid L_{0}do(\bar{A}_{1})}$ and $C_{\tilde{Y}_{2}L_{0}\mid do(A_{0})}$, parameterized by $\theta_{1,1}$ and $\theta_{1,0}$, respectively.
-These two copulas associate $\tilde{Y}_{2}$ with previous time-varying confounders.
+$F_{\widetilde{Y}_{1}\,|\,do(A_{0})}(\widetilde{y}_{1}\,|\,do(a_{0}))$ $=h^{-1}\big{\{}F_{\widetilde{Y}_{1}\,|\,L_{0},do(A_{0})}(\widetilde{y}_{1}\,|\,l_{0},do(a_{0})),F_{L_{0}}(l_{0});\Theta_{0,0}\big{\}},$
+set $w_{0,3}=F_{\widetilde{Y}_{1}\,|\,L_{0},do(A_{0})}(\widetilde{y}_{1}\,|\,l_{0},do(a_{0}))$ and substitute $F_{L_{0}}(l_{0})$ with $w_{0,1}$, yielding:
+$u_{1}=h^{-1}\left(w_{0,3},w_{0,1};\Theta_{0,0}\right).$
+By inversion, compute $\widetilde{y}_{1}=F_{Y_{1}\mid do(A_{0})}^{-1}(u_{1})$, and the corresponding binary outcome $y_{1}=I(\widetilde{y}_{1}<1)$.
+This completes the first iteration, yielding a sample $(l_{0},a_{0},\widetilde{y}_{1},y_{1})$.
+If $y_{1}=1$, stop; otherwise, before proceeding to $k=1$, we store the following values:
+$u_{0,0}=F_{L_{0}}(l_{0})$ $q_{0,0}=F_{\widetilde{Y}_{1}\mid do(A_{0})}(\widetilde{y}_{1}=1\,|\,do(a_{0})).$
+For $k=1$, define the necessary cumulative distribution functions $F_{L_{1}\,|\,L_{0},A_{0}}$, $F_{A_{1}\mid\bar{L}_{1},A_{0}}$, $F_{\widetilde{Y}_{2}\mid do(\bar{A}_{1})}$, and copulas $C_{\widetilde{Y}_{2}L_{1}\mid L_{0}do(\bar{A}_{1})}$ and $C_{\widetilde{Y}_{2}L_{0}\mid do(A_{0})}$, parameterized by $\Theta_{1,1}$ and $\Theta_{1,0}$, respectively.
+These two copulas associate $\widetilde{Y}_{2}$ with previous time-varying confounders.
 Draw $w_{1,1}$, $w_{1,2}$, $w_{1,3}$ from $U(0,1)$, and obtain $(l_{1},a_{1})$ as before.
-We now need to obtain a sample of $\tilde{y}_{2}$. Calculate:
-$F(l_{0}\,|\,do(a_{0}),y_{1}=0)$ $=h\big{\{}F(l_{0}),F(\tilde{y}_{1}\,|\,do(a_{0}));\theta_{0,0}\big{\}}$ (19)
-$F(\tilde{y}_{2}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0)$ $=h^{-1}\big{\{}F(\tilde{y}_{2}\,|\,\bar{l}_{1},do(\bar{a}_{1}),y_{1}=0),F(l_{1}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0);\theta_{1,1}\big{\}}$
-$F(\tilde{y}_{2}\,|\,do(a_{1}),y_{1}=0)$ $=h^{-1}\big{\{}F(\tilde{y}_{2}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0),F(l_{0}\,|\,do(\bar{a}_{1}),y_{1}=0);\theta_{1,0}\big{\}}.$
+We now need to obtain a sample of $\widetilde{y}_{2}$. Calculate:
+$F(l_{0}\,|\,do(a_{0}),y_{1}=0)$ $=\frac{u_{0,0}-C^{*}_{\widetilde{Y}_{1}L_{0}\mid A_{0}}(u_{0,0},q_{0,0};\Theta_{0,0})}{1-q_{0,0}}$ (21)
+$F(\widetilde{y}_{2}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0)$ $=h^{-1}\big{\{}F(\widetilde{y}_{2}\,|\,\bar{l}_{1},do(\bar{a}_{1}),y_{1}=0),F(l_{1}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0);\Theta_{1,1}\big{\}}$
+$F(\widetilde{y}_{2}\,|\,do(a_{1}),y_{1}=0)$ $=h^{-1}\big{\{}F(\widetilde{y}_{2}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0),F(l_{0}\,|\,do(\bar{a}_{1}),y_{1}=0);\Theta_{1,0}\big{\}}.$
 Note that we omit some subscripts on quantiles for brevity.
-In (19) we derive the quantile of $L_{0}$ given $a_{0}$ among survivors the start of the $k=1$ interval, where $F(l_{0})=w_{0,0}$ and $F(\tilde{y}_{1}\,|\,do(a_{0})$ is already computed in (18) in step $k=0$.
+In (21) we derive the quantile of $L_{0}$ given $a_{0}$ among survivors the start of the $k=1$ interval, where $u_{0,0}$ and $q_{0,0}$ are already computed in step $k=0$.
 All other necessary quantities are specified at $k=1$.
+These computations yields a sample $(l_{1},a_{1},\widetilde{y}_{2},y_{2})$.
+If $y_{2}=1$, the simulation stops; otherwise, we store and compute the following values:
+$u_{1,0}$ $=F_{L_{0}\mid Y_{1}}(l_{0}\,|\,y_{1}=0)$
+$q_{1,0}$ $=F_{\widetilde{Y}_{2}\mid do(\bar{A}_{1})Y_{1}}(\widetilde{y}_{2}=1\,|\,do(\bar{a}_{1}),y_{1}=0)$
+$u_{1,1}$ $=F_{L_{1}\mid L_{0}Y_{1}}(l_{0}\,|\,l_{0},y_{1}=0)$
+$q_{1,1}$ $=F_{\widetilde{Y}_{2}\mid L_{0}do(\bar{A}_{1})Y_{1}}(\widetilde{y}_{2}=1\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0)$
+$=h\big{\{}q_{1,0};u_{1,0};\Theta_{1,0}\big{\}},$
+where we simply store the first three quantities, which are directly available from the current step, while only the last one requires computation.
 Therefore, the base case is satisfied since the algorithm simulates the first time step and provides the required inputs for the next.
@@
 We need to compute the updated quantiles $F(l_{j}\mid\overline{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$ for each $j<k$, given from the previous iteration:
-$F(l_{j}\mid\overline{l}_{j-1},do(\bar{a}_{k-1}),y_{k-1}=0),$ $F(y_{k}\mid\overline{l}_{j},do(\bar{a}_{k-1}),y_{k-1}=0).$
+$u_{k-1,j}=F(l_{j}\mid\overline{l}_{j-1},do(\bar{a}_{k-1}),y_{k-1}=0),$ $q_{k-1,j}=F(\widetilde{y}_{k}\mid\overline{l}_{j-1},do(\bar{a}_{k-1}),y_{k-1}=0).$
 This can be achieved by noting that
 $F(l_{j}\mid\overline{l}_{j-1},do(\bar{a}_{k-1}),y_{k-1}=0)=F(l_{j}\mid\overline{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0),$
+and
+$F(\widetilde{y}_{k}\mid\overline{l}_{j-1},do(\bar{a}_{k-1}),y_{k-1}=0)=F(\widetilde{y}_{k}\mid\overline{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0),$
 which allows us to apply (14) to compute the required quantiles.
@@
 By the principle of induction, the statement is true for all $k\geq 0$. ∎
-Appendix C Competing risk
-At each time step $k$, after simulating time-varying covariate $L_{k}$ and treatment $A_{k}$, we generate the primary outcome $Y_{k}$ and a competing risk $D_{k}$ using their respective auxiliary variables $\tilde{Y}_{k}$ and $\tilde{D}_{k}$, following the approach outlined in Section 5.
-For brevity, baseline covariates $\bm{Z}$ continue to be omitted, but in settings where they are present, they can be simulated as the initial step and included in the conditioning set of all subsequent quantities.
-Under this simulation framework, we specify:
-  • A marginal structural model for the primary event: $F^{*}(\tilde{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0,D_{k}=0)$;
-  • A marginal structural model for the competing event: $F^{*}(\tilde{D}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0,D_{k}=0)$;
-  • A bivariate copula parameterized by $\Psi_{k+1}$ to model the dependence between the above interventional margins of $\tilde{Y}_{k+1}$ and $\tilde{D}_{k+1}$;
-  • Bivariate copulas that ‘weave’ the multivariate dependence between $Y_{k+1}$ and past covariates $\bar{L}_{k}$, given $do(\bar{A}_{k})$, $Y_{k}=0$ and $D_{k}=0$, parameterized by $\Theta_{k,j}$ for $j=0,\ldots,k$;
-  • Bivariate copulas that ‘weave’ the multivariate dependence between $\tilde{D}_{k+1}$ and $\bar{L}_{k}$ conditional on $\tilde{Y}_{k+1}$, $do(\bar{A}_{k})$, $Y_{k}=0$ and $D_{k}=0$, parameterized by $\Phi_{k,j}$ for $j=0,\ldots,k$.
-The joint outcome model $p(\tilde{y}_{k+1},\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0)$ factorizes as follows:
-Primary event model:
-$p(\tilde{y}_{k+1}$ $\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0)$
-$=p^{*}(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)\times$
-$\prod_{j=0}^{k}c^{*}_{\tilde{Y}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},Y_{k}=0,D_{k}=0}\big{\{}F(\tilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0),$
-$\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0);\Theta_{k,j}\}.$ (20)
-Competing event model:
-$p(\tilde{d}_{k+1}$ $\,|\,\bar{l}_{k},\bar{a}_{k},\tilde{y}_{k+1},D_{k}=0)$
-$=p(\tilde{d}_{k+1}\,|\,do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)\times$
-$\prod_{j=0}^{k}c^{*}_{\tilde{D}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},\tilde{Y}_{k+1}D_{k}=0}\big{\{}F(d_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0),$
-$\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0);\Phi{k,j}\}.$ (21)
-The first term in (21), the distribution of $\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},\tilde{y}_{k+1},D_{k}=0$ can be obtained through:
-$F(\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},\tilde{y}_{k+1},D_{k}=0)=h\big{\{}F*(\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0),F^{*}(\tilde{y}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0)\big{\}}.$
-The challenge lies in computing $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$ in (20), and $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$ in (21), for $j=0,\ldots,k$.
-The former can be calculated in the same iterative fashion as described in Section 5.
-The latter can be derived as:
-$F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$ $=h\big{\{}F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0),$
-$\qquad F(\tilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0);\Theta_{k,j}\big{\}}.$ (22)
-Notably, the two components in the $h$-function in (22) exactly correspond to the those in the primary event model in (20), except with the order of conditioning swapped.
-Since these values are computed during the simulation of $\tilde{y}_{k+1}$, requiring only $k+1$ additional evaluations of the $h$-functions in (22).
-The algorithm for simulating each time step $k$ for an individual can be summarized into:
-  1. 1.
-    Simulate the time-varying confounder $L_{k}$ and treatment $A_{k}$ from their respective distributions.
-  2. 2.
-    Compute $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$ for $j=0,\dots,k-1$.
-  3. 3.
-    Compute $F(y_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$ for $j=1,\dots k$, leading to $u_{k+1}=F(\tilde{y}_{k+1}\,|\,z,do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$.
-  4. 4.
-    Compute: $\tilde{y}_{k+1}=F^{*-1}_{\tilde{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0,D_{k}=0}(u_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$.
-  5. 5.
-    Compute $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$ for $j=0,\ldots,k$ by evaluating $k+1$ $h$-functions following (22).
-  6. 6.
-    Compute $F(\tilde{d}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1}=0,D_{k}=0)$ for $j=1,\ldots k$, leading to $v_{k+1}=F(\tilde{d}_{k+1}\,|\,do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$.
-  7. 7.
-    Compute: $\tilde{d}_{k+1}=F_{\tilde{D}_{k+1}\,|\,do(\bar{A}_{k}),\tilde{Y}_{k+1},D_{k}=0}^{*-1}(\tilde{d}_{k+1}\,|\,do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$.
-This approach requires at most $74k+$ 7 evaluations of analytic functions per individual per time step. An example implementation is available at https://github.com/XiLinStats/LongSiMSM.
 Appendix D Additional details of the illustration in Section 5.2
-Models for baseline, time-varying covariates and treatment
-A baseline:
+[…]
+$B\sim\operatorname{Bernoulli}(0.7)$ $L_{0}\sim\operatorname{Gamma}\!\left(k=(25+0.8B),\;\theta=2.5\right)$
+$A\sim\operatorname{Bernoulli}(0.5)$ $S_{0}\sim\operatorname{Bernoulli}(0.1).$
+The choice of these parameter values are loosely based on the baseline characteristics of SOUL reported in McGuire et al. (2023). For subsequent follow-up visits $k>0$:
+$S_{k}$ $\sim\operatorname{Bernoulli}(\operatorname{expit}(\alpha_{0}+\alpha_{1}\,B+\alpha_{2}\,L_{k}+\alpha_{3}\,A))$
+$L_{k}$ $\sim\operatorname{Gamma}(k=400(L_{k-1}+0.2\,B-0.5\,S_{k-1}-A_{k-1}),\;\theta=\frac{1}{400}).$
+We target a Cox marginal structural model given by:
+$\lambda(\widetilde{y}_{k+1}\,|\,do(A,\bar{S}_{k}),Y_{k}=0)=\lambda_{0}\exp\left(\beta_{A}\,A+\beta_{S}\,S_{k}+\beta_{AS}\,AS_{k}\right)=g_{k}(A,S_{k}),$
+This parameterization assumes a constant hazard rate between follow-up visits, influenced by the uses of semaglutide and SGLT2-i at the most recent visit.
+This implies that given survival at follow-up time $k$, the incremental survival time $\widetilde{Y}_{k+1}$ follows an exponential distribution with the rate equal to the hazard above:
+$\widetilde{Y}_{k+1}\sim\operatorname{Exp}\big{(}g_{k}(A,S_{k})\big{)}.$
+If $\widetilde{Y}_{k+1}<1$ then $Y_{k+1}=1$, and 00 otherwise.
+To understand the impact of differential use of SGLT2-i on the estimation of treatment effect of the primary treatment, semaglutide, we create several scenarios with different prevalence of SGLT2-i in the placebo arm.
+We achieve this by varying the coefficients ($\alpha_{0}$,$\alpha_{1}$,$\alpha_{2}$ and $\alpha_{3}$) of the model for $S_{k}$.
+We also explored varying the efficacy of SGLT2-i.
+Table 6 below describes 4 levels of prevalence and 3 levels of efficacy of SGLT2-i.
+In total, we simulate 10 scenarios.
+SGLT2-i Prevalence Scenarios Baseline $S_{k}\sim\text{Bernoulli}(\text{expit}(-9.5+0.5B+0.1L_{k}-1.5A))$.
+Annual take-up: 10% in placebo, 2.5% in semaglutide group.
+High drop-in $S_{k}\sim\text{Bernoulli}(\text{expit}(-9+0.5B+0.1L_{k}-2A))$.
+Annual take-up: 15% in placebo, 2.5% in semaglutide group.
+Low drop-in $S_{k}\sim\text{Bernoulli}(\text{expit}(-10+0.5B+0.1L_{k}-A))$.
+Annual take-up: 5% in placebo, 2.5% in semaglutide group.
+No drop-in $S_{k}=0$ Effect Scenarios Moderate effect $\beta_{S}=\log(0.9)$, $\beta_{AS}=0.05$.
+Strong effect $\beta_{S}=\log(0.85)$, $\beta_{AS}=0.1$.
+Weak effect $\beta_{S}=\log(0.95)$, $\beta_{AS}=0.01$.
+Table 6: Descriptions of SGLT2-i prevalence and efficacy scenarios.
+The prevalence scenarios vary the annual uptake of SGLT2-i in the placebo and semaglutide groups, while efficacy scenarios are defined by combinations of $\beta_{S}$ and $\beta_{AS}$, representing the isolated and interaction effects, respectively.
+The values of $\beta_{S}$ are based on the findings of McGuire et al.
+(2021), who reported that SGLT2 inhibitors were associated with a reduced risk of MACE, with a hazard ratio of $0.90$ ($95\%$ confidence interval: 0.85–0.95) from a meta-analysis of six clinical trials.
+The interaction term represents the expectation that the combined effect of using both treatments is less than the total of their individual effects.
+For each simulated dataset, we fit the above Cox marginal structural model $\lambda(\widetilde{y}_{k+1}\,|\,do(A,\bar{S}_{k}),Y_{k}=0)=\lambda_{0}\exp\left(\beta_{A}\,A+\beta_{S}\,S_{k}+\beta_{AS}\,AS_{k}\right)$ using stabilized inverse-probability-of-treatment weights (Hernán et al., 2000), calculated as
+$W_{k}=\prod_{j=0}^{k}\frac{\hat{p}(S_{j}\mid A,\bar{S}_{j-1},Y_{j}=0)}{\hat{p}(S_{j}\mid B,A,\bar{L}_{j},\bar{S}_{j-1},Y_{j}=0)}.$ (22)
+Remark D.1.
+The denominator in the weight formula represents the correctly specified propensity score model under the assumed data-generating process, while the numerator stabilizes the weights.
+To avoid unintended confounding, the baseline covariates included in the numerator must also be adjusted for in the marginal structural model (Cole and Hernán, 2008).
+In accordance with the model in (23), we only include $A$ in the numerator for stabilization.
+Appendix E Seaman and Keogh (2024) simulations
+We present an example that broadly replicates the setting described in Seaman and Keogh (2024).
+We consider baseline variables $B$ and $C$, two time-varying covariates $L_{k}$ and $W_{k}$, and treatment $A_{k}$ at follow-up $k$.
+The time-to-event is $T$, with $Y_{k+1}=I(T<k+1)$ indicating failure status observed at the end of the $k$th interval.
+The auxiliary continuous variable $\widetilde{Y}_{k+1}$ is defined as previously.
+Specifically we choose that at baseline:
 $B\sim\operatorname{Exp}(2)$ $L_{0}\sim\operatorname{Gamma}(k=1+0.5\,B+0.5\,C,\;\theta=1)$
@@
 $W_{k}$ $\sim\operatorname{Bernoulli}(\operatorname{expit}\{-0.2+0.5\,B+0.5\,C+W_{k-1}-0.6\,A_{k-1}\}).$
-Parameter Scenarios
-Table6 summarizes the parameter values for treatment effect size and confounding scenarios:
+We target a Cox marginal structural model given by:
+$\lambda(\widetilde{y}_{k+1}\,|\,do(\bar{a}_{k}),c,y_{k}=0)=\lambda_{0}\exp\left(\beta_{C}\,c+\beta_{A}\,a_{k}+\beta_{CA}\,c\,a_{k}\right),$ (23)
+where $\lambda_{0}=\exp(-2)$ is the baseline hazard.
+We set $\beta_{C}=-0.5$ and vary the values of $(\beta_{A},\beta_{CA})$ to construct scenarios with different strengths of treatment: a strong effect scenario $(-0.5,0.3)$ and a weak effect scenario $(0.2,0.1)$.
+This parameterization assumes a constant hazard rate between follow-up visits, influenced by the treatment status at the most recent visit.
+This implies that given survival at follow-up time $k$, the incremental survival time $\widetilde{Y}_{k+1}$ follows an exponential distribution with the rate equal to the hazard above:
+$\widetilde{Y}_{k+1}\sim\operatorname{Exp}(\lambda_{0}\exp\left(\beta_{C}\,c+\beta_{A}\,a_{k}+\beta_{CA}\,c\,a_{k}\right)).$
+If $\widetilde{Y}_{k+1}<1$ then $Y_{k+1}=1$, and 00 otherwise.
+We set the total number of follow-ups $K$ to 6.
+We set the copulas describing the association between $L_{j}$ ($j=0,\ldots,k$) and $Y_{k+1}$ to be Student t with parameter $\rho_{L}$ and 5 degrees of freedom, although each pairwise dependence could, in principle, be distinct.
+Similarly, we model the copulas capturing the dependence between $Y_{k+1}$ and each $W_{j}$ ($j=0,\ldots,k$), given $L_{k}$, as Gaussian with correlation $\rho_{W}$, though again we note that different copulas could be specified for each time point.
+We vary the values of $\rho_{L}$ and $\rho_{W}$ to construct scenarios with different strengths of confounding.
+Table 7 summarizes the combinations of treatment effect size and confounding levels considered.
 Effect size Large $(\beta_{A},\beta_{CA})=(-0.5,0.3)$ Small $(\beta_{A},\beta_{CA})=(0.2,0.1)$ Level of confounding High $(\rho_{L},\rho_{W})=(0.9,0.7)$ Medium $(\rho_{L},\rho_{W})=(-0.5,0.4)$ Low $(\rho_{L},\rho_{W})=(-0.1,-0.2)$
 Table 7: Parameter values of treatment effect size and confounding scenarios.
-Runtime details
-Table 7 reports the runtime required to simulate data under the setup for the large causal effect with medium confounding scenario.
-The simulations were carried out on a local laptop using a single score without parallelization.
-Generating data for $n$ up to 10,000 units takes only a few seconds, while even for a very large sample size ($n=50,\!000$) and 11 time steps ($K=10$), the runtime remains under one minute.
-This computational efficiency is crucial for tasks such as method evaluation and sample size calculations in clinical trial design, where simulations must be repeated across various parameter values.
-For example, determining the sample size might require testing 10 different parameter combinations (such as effect sizes, prevalence of drop-in treatments, or adherence levels) across four sample sizes, each bootstrapped 200 times to estimate statistical power.
-This requires $10\times 4\times 200=8,\!000$ simulations.
-Using our method, suppose $n=10,000$ and $K=10$, the total runtime is approximately 13 hours on a single core, which can be reduced to about 2 hours using an 8-core laptop or under 10 minutes on a 50-core CPU through parallelization.
-In contrast, simulation methods requiring computationally intensive operations, such as Monte Carlo approximations (see, for example, Seaman and Keogh, 2024) or numerical integration, are often too slow to support such large-scale simulations.
-This highlights a key advantage of our proposed method: its ability to handle extensive simulation demands efficiently, making it well-suited for practical applications.
-$n$ $K=5$ $K=10$ 1,000 0.27 0.79 2,000 0.45 1.68 5,000 1.69 2.49 10,000 2.70 6.09 50,000 16.50 45.49
-Table 7: Runtime (in seconds) to simulate a dataset of size $n$, averaged over 20 repeats.
-Weight calculation
-To fit the Cox marginal model, we employ stabilized inverse probability of treatment weighting, with weights for subject $i$ at time step $k$ calculated as:
+We simulate data for sample size $n=\{1000,2000,5000\}$.
+We contrast two approaches for fitting the Cox marginal structural model specified in (23): stabilized inverse-probability-weighting and unweighted estimation.
+The first employs stabilized inverse probability of treatment weighting, with weights for subject $i$ at time step $k$ calculated as:
 $\prod_{j=0}^{k}\frac{\hat{p}_{j}(A_{j}\mid C,\bar{A}_{j-1},Y_{j}=0)}{\hat{p}_{j}(A_{j}\mid B,C,\bar{L}_{j},\bar{W}_{j},\bar{A}_{j-1},Y_{j}=0)},$
 where $\hat{p}_{j}(A_{j}\mid C,\bar{A}_{j-1},Y_{j}=0)$ and $\hat{p}_{j}(A_{j}\mid B,C,\bar{L}_{j},\bar{W}_{j},\bar{A}_{j-1},Y_{j}=0)$ are estimated using logistic regression models fitted to the survivors at time step $j$.
-The denominator in the weight formula represents the correctly specified propensity score model under the assumed data-generating process, while the numerator stabilizes the weights.
-To avoid unintended confounding, the baseline covariates included in the numerator must also be adjusted for in the marginal structural model (Cole and Hernán, 2008).
-In accordance with the model in (15), we only include $C$ in the numerator for stabilization.
+Table 8 summarizes the bias of estimates from 500 replications, compared to corresponding Monte Carlo standard errors as a measure of uncertainty.
+As the sample size increases, generally, the biases of estimates for $\beta_{C}$, $\beta_{A}$ and $\beta_{AC}$ obtained with inverse probability weights converge to zero, with their absolute values substantially smaller than the corresponding Monte Carlo standard errors.
+In contrast, the unweighted estimator exhibits considerable bias, particularly for the estimate of the treatment effect, $\beta_{A}$, which persists even at $n=5000$.
+The magnitude of this bias far exceeds the Monte Carlo standard errors.
+The unweighted estimates are most biased under the high level of confounding by $L$ and $W$, since this is not adjusted for.
+Level of confounding Effect size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ IPTW Unweighted IPTW Unweighted IPTW Unweighted High $1000$ $0.00(2.18)$ $-0.74(1.46)$ $-0.16(2.26)$ $2.75(1.88)$ $0.21(2.99)$ $0.22(2.20)$ $\beta_{A}=-0.5$ $2000$ $-0.07(1.50)$ $-0.74(1.00)$ $-0.13(1.56)$ $2.71(1.25)$ $0.24(2.06)$ $0.24(1.51)$ $5000$ $-0.04(1.13)$ $-0.78(0.67)$ $-0.02(1.03)$ $2.89(0.80)$ $0.04(1.46)$ $0.11(1.02)$ $1000$ $0.09(2.26)$ $-0.51(1.60)$ $0.27(1.92)$ $3.00(1.56)$ $-0.16(2.76)$ $0.07(2.02)$ $\beta_{A}=0.2$ $2000$ $0.04(1.64)$ $-0.63(1.15)$ $0.07(1.54)$ $2.84(1.15)$ $-0.02(2.04)$ $0.21(1.47)$ $5000$ $-0.04(1.07)$ $-0.73(0.70)$ $0.03(0.91)$ $2.81(0.68)$ $0.03(1.26)$ $0.31(0.88)$ Medium $1000$ $-0.13(2.12)$ $-0.77(1.48)$ $0.08(2.40)$ $2.66(1.78)$ $0.00(3.12)$ $0.20(2.23)$ $\beta_{A}=-0.5$ $2000$ $-0.01(1.60)$ $-0.79(0.97)$ $0.03(1.72)$ $2.53(1.25)$ $0.10(2.29)$ $0.39(1.52)$ $5000$ $0.00(1.04)$ $-0.74(0.60)$ $-0.04(1.09)$ $2.58(0.79)$ $0.10(1.52)$ $0.30(0.96)$ $1000$ $0.01(2.34)$ $-0.70(1.56)$ $0.28(1.97)$ $2.63(1.54)$ $-0.11(2.90)$ $0.30(2.05)$ $\beta_{A}=0.2$ $2000$ $0.01(1.73)$ $-0.67(1.04)$ $0.06(1.38)$ $2.50(1.08)$ $0.01(2.02)$ $0.35(1.37)$ $5000$ $-0.02(1.12)$ $-0.73(0.68)$ $0.05(1.03)$ $2.46(0.66)$ $0.05(1.41)$ $0.41(0.86)$ Low $1000$ $-0.09(2.04)$ $-0.11(1.41)$ $-0.05(2.40)$ $0.64(1.75)$ $0.12(3.12)$ $-0.12(2.25)$ $\beta_{A}=-0.5$ $2000$ $0.02(1.57)$ $-0.12(0.97)$ $0.02(1.70)$ $0.63(1.22)$ $0.09(2.26)$ $0.05(1.50)$ $5000$ $-0.06(1.07)$ $-0.12(0.65)$ $-0.01(1.11)$ $0.70(0.77)$ $0.08(1.47)$ $-0.07(0.94)$ $1000$ $0.14(2.21)$ $0.04(1.44)$ $0.27(2.09)$ $0.72(1.54)$ $-0.16(2.85)$ $-0.14(1.92)$ $\beta_{A}=0.2$ $2000$ $-0.02(1.75)$ $-0.08(1.05)$ $0.01(1.62)$ $0.58(1.11)$ $0.12(2.23)$ $-0.01(1.41)$ $5000$ $0.11(1.03)$ $-0.07(0.60)$ $0.10(0.89)$ $0.58(0.64)$ $-0.11(1.32)$ $-0.02(0.81)$
+Table 8: Bias and Monte Carlo standard errors (×$\times 10$) of estimates produced by inverse probability of treatment weighted (‘IPTW’) and unweighted Cox regressions, under different strengths of confounding, effect sizes, and sample sizes.
+Table 9 presents coverage rates of the 95% confidence intervals.
+Confidence intervals for the weighted estimator, computed using both sandwich variance estimation and bootstrap resampling, achieve coverage close to the nominal level, with slight undercoverage.
+In contrast, the sandwich-based confidence intervals for the unweighted estimator perform poorly, especially for the treatment effect estimate, $\beta_{A}$.
+Level of confounding Effect Size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ W-S W-B UW W-S W-B UW W-S W-B UW High 1000 93.5% 94.3% 92.5% 93.8% 93.5% 60.3% 94.3% 93.8% 95.0 % $\beta_{A}=-0.5$ 2000 94.8% 95.3% 89.8% 95.3% 93.3% 39.0% 95.3% 93.8% 95.5 % 5000 92.3% 92.0% 75.5% 93.8% 94.5% 5.8% 94.0% 92.8% 93.5 % 1000 93.5% 94.3% 93.0% 95.5% 95.5% 49.3% 94.5% 94.3% 93.0 % $\beta_{A}=0.2$ 2000 92.5% 93.0% 90.0% 93.8% 92.5% 26.5% 94.0% 93.5% 92.8 % 5000 93.5% 94.0% 82.0% 94.8% 93.5% 1.0% 94.8% 94.3% 94.0 % Medium 1000 94.8% 94.8% 91.3% 94.3% 95.0% 64.0% 93.5% 93.3% 95.8 % $\beta_{A}=-0.5$ 2000 94.3% 94.3% 88.8% 93.8% 93.8% 44.3% 95.0% 94.0% 95.8 % 5000 97.5% 96.8% 83.0% 95.0% 94.8% 8.8% 94.3% 93.5% 92.8 % 1000 93.3% 93.3% 91.8% 93.8% 93.5% 58.3% 92.0% 93.8% 94.8 % $\beta_{A}=0.2$ 2000 93.5% 94.8% 91.0% 96.5% 96.8% 36.3% 94.8% 95.8% 93.5 % 5000 96.0% 94.5% 79.8% 94.5% 93.8% 6.5% 96.0% 95.0% 91.8 % Low 1000 93.5% 93.8% 94.0% 93.5% 93.3% 94.0% 93.5% 94.0% 94.5 % $\beta_{A}=-0.5$ 2000 94.8% 95.3% 94.0% 95.8% 95.3% 92.5% 94.0% 94.5% 95.8 % 5000 93.8% 94.3% 92.3% 95.3% 95.0% 86.0% 95.3% 93.8% 95.0 % 1000 93.0% 94.8% 94.8% 93.3% 94.5% 92.8% 95.3% 95.8% 94.5 % $\beta_{A}=0.2$ 2000 92.0% 92.3% 94.0% 92.5% 93.8% 90.8% 92.3% 94.0% 94.0 % 5000 96.5% 96.2% 96.0% 95.2% 95.2% 88.5% 95.0% 94.5% 96.5 %
+Table 9: Coverage rates for IPTW-Sandwich (W-S), IPTW-Bootstrap (W-B), and unweighted (UW) estimators of $\beta_{C}$, $\beta_{A}$, and $\beta_{AC}$ under different strengths of confounding, effect sizes, and sample sizes.
+In summary, the asymptotically unbiased nature of the inverse probability weighted estimates, compared to the biased unweighted estimates, further confirms that the simulated data were generated in accordance with the Cox model in (23), and that confounding by $L$ and $W$ was appropriately introduced.